### v2

# Georgy Lukyanov Toulouse School of Economics Anna Vlasova

Abstract
We study dynamic delegation with reputation feedback: a long-lived expert advises a sequence of implementers whose effort responds to current reputation, altering outcome informativeness and belief updates. We solve for a recursive, belief-based equilibrium and show that advice is a reputation-dependent cutoff in the expert’s signal. A diagnosticity condition—failures at least as informative as successes—implies reputational conservatism: the cutoff (weakly) rises with reputation. Comparative statics are transparent: greater private precision or a higher good-state prior lowers the cutoff, whereas patience (value curvature) raises it. Reputation is a submartingale under competent types and a supermartingale under less competent types; we separate boundary hitting into learning (news generated infinitely often) versus no-news absorption. A success-contingent bonus implements any target experimentation rate with a plug-in calibration in a Gaussian benchmark. The framework yields testable predictions and a measurement map for surgery (operate vs. conservative care).
Keywords: Dynamic delegation; expert advice; moral hazard; experimentation; reputational conservatism.
JEL: D82, D83, C73.
1 Introduction
Expert advice shapes high-stakes decisions in many arenas: surgeons decide whether to operate or pursue conservative care; financial analysts issue buy or hold calls that investors act upon; R&D leads greenlight or delay projects; policy consultants argue for reform or the status quo. In all of these settings, outcomes depend not only on whether the idea is good but also on how diligently it is implemented. Implementation, in turn, responds to the expert’s standing: trusted experts are heeded and their recommendations are executed with greater effort. This two-way link—reputation raising effort, and effort making outcomes more informative—creates what we call reputation feedback. Empirical patterns are consistent with this mechanism: patient adherence rises with clinician trust and communication quality (Haskard Zolnierek and DiMatteo, 2009; Birkhäuer et al., 2017); surgical outcomes vary sharply with surgeon skill and volume (Birkmeyer et al., 2003; Birkmeyer et al., 2013); and influential analyst recommendations move implementation scale and outcomes in financial markets (Loh and Stulz, 2011).
We build a tractable dynamic model of expert advice that isolates reputation feedback and turns it into sharp equilibrium predictions. A long-lived expert repeatedly recommends a risky action to a sequence of short-lived agents (implementers). Success requires both a favorable technological state and implementer effort; recommendations and outcomes are publicly observed and update the expert’s perceived competence. Because each agent’s optimal effort responds to current reputation, the expert internalizes that today’s advice affects both the chance of success and what the outcome will reveal. The equilibrium characterization is recursive and belief-based.
Our main results are threefold. First, we show thatfor any public reputation there is a cutoff in the expert’s private signal above which she recommends the risky action and below which she advises safety (Theorem 5). This delivers portable comparative statics and clarifies on-path mimicry by less competent experts. Second, we characterize reputation dynamics: posterior beliefs form a submartingale under a competent expert and a supermartingale under an incompetent one, and the belief process hits trust regions with probability one (Theorem 12). Third, we establish reputational conservatism: the risky-signal cutoff is increasing in current reputation (Theorem 7). Intuitively, as reputation rises, implementers work harder, so a failure becomes more diagnostic of incompetence while the informativeness of success does not increase; with a convex value of reputation, the downside risk dominates at high reputation. We also provide transparent comparative statics (Theorem 17): better private information and a higher probability that the environment is favorable lower the cutoff and raise experimentation, whereas greater patience amplifies conservatism.
The model yields testable predictions. Holding fundamentals fixed, highly reputed experts recommend the risky action less frequently but achieve higher success rates when they do, because implementers work harder. Transitory reputation shocks (e.g., early successes) reduce the frequency of risky calls but make subsequent failures more revealing and more reputationally costly; negative shocks have the opposite effect. These predictions echo empirical patterns in settings where effort is observable or proxied (e.g., adherence) and where performance differences across experts are documented (Haskard Zolnierek and DiMatteo, 2009; Birkhäuer et al., 2017; Birkmeyer et al., 2003; Birkmeyer et al., 2013). They speak to the design of advisory relationships in medicine (operate vs. conservative care), finance (buy vs. hold and trade size), innovation management (greenlight vs. delay), and public policy (reform vs. status quo), where the scale and diligence of implementation are endogenous to the expert’s standing. [^1] A practical recipe for constructing reputation and effort proxies is in Online Appendix OA.2. Computational details for value iteration, cutoff computation, and simulations are in Online Appendix OA.3.
Two modeling choices drive tractability and clarity. First, we keep the public state one-dimensional (the current reputation about competence) and assume independent period states for the technology. Second, success requires both a good state and implementer effort, which is optimally chosen in response to current reputation and the recommendation. These assumptions allow a recursive solution and clean comparative statics while capturing the core mechanism: reputation feeds effort, and effort feeds learning from outcomes.
Related work spans dynamic career concerns, strategic expert advice, and delegated experimentation. Our contribution is to embed the implementer’s effort response within a recursive reputation model and to show how that endogeneity reshapes both policy (a reputation-dependent threshold with conservatism) and learning (martingale dynamics with boundary hitting).
The paper proceeds as follows. Section 2 reviews related work. Section 3 sets up the environment, belief updates, and recursive equilibrium. Section 4 presents the threshold equilibrium, a binary–signal illustration, reputational conservatism, comparative statics, and reputation dynamics with a learning–absorption decomposition. Section 5 treats the surgery application and policy design via success–contingent bonuses, and covers monitoring, committees, endogenous exit, and a continuous–time approximation. Core proofs are in Appendices AD; algorithms, Gaussian formulas, robustness, committee variants, behavioral twists, and policy calibration appear in the Online Appendix OA.2OA.8).
2 Related Literature
This paper connects dynamic career-concerns and reputation, strategic expert advice with reputational motives, and delegated experimentation/bandit models with hidden actions. Our contribution is to embed the implementer’s effort response into a recursive reputation model and to characterize how this reputation–effort complementarity shapes policies and the law of motion of beliefs.
A first strand is the dynamic reputation and career-concerns tradition, where forward-looking agents distort current behavior to affect future assessments. The classic insight originates with Holmström’s (1982) career concerns. Reputations can serve as implicit incentive schemes (Tadelis, 2002), but their informational content evolves endogenously and, under rich observation, can unravel in the long run (Cripps et al., 2004). We take the infinite-horizon viewpoint and show that reputation is a submartingale (supermartingale) for competent (incompetent) experts, with boundary hitting. Our reputational conservatism—cutoffs increasing in current reputation—relates to the tension between conservatism and gambling in dynamic career-concerns models (Prendergast and Stole, 1996). In a complementary direction, Mylovanov and Klein (2017) show that sufficiently long horizons can discipline informational biases; our recursive solution clarifies the channel via the diagnosticity of outcomes when clients exert more effort at higher reputation. In relational and repeated settings, reputational incentives can be misaligned and generate inefficient equilibria; Deb et al. (2022) analyze such forces and provide sharp characterizations. The reputation–effort loop we formalize is adviser-side and distinct, but it leads to analogous selection and path-dependence phenomena.
A second strand studies strategic expert advice when advisers care about reputation. In reputational cheap talk with continuous signals, experts pool messages to manage reputational risk, yielding coarse disclosure (Ottaviani and Sørensen, 2006). We provide a dynamic counterpart with a state-dependent cutoff that moves with public reputation. Conformity and caution under career concerns have been emphasized in settings where “being wrong alone” is worse than “being wrong with the crowd” (Scharfstein and Stein, 1990; Prendergast, 1993; Morris, 2001). In our model, similar caution emerges through the effort channel: when clients work harder for highly reputed experts, a failure is more revealing, pushing such experts toward conservatism. With short-lived customers, initial honesty may profitably build a reputation that can later be exploited (Ozyurt, 2016). Paradoxically, very competent advisers with spotless records may be fired because their behavior is inferred to be uninformative or biased (Schottmüller, 2019); our law of motion clarifies when occasional visible failures are actually informative (and thus desirable) versus reputation-damaging. Relatedly, “good lies” can build credibility in the long run (Pavesi and Scotti, 2022); in our environment, explicit outcomes discipline such strategies because failures become more diagnostic when reputation is high.
A third strand is delegated experimentation and bandit models with hidden actions. Strategic experimentation highlights informational externalities and dynamic incentives (Bolton and Harris, 1999; Keller et al., 2005); our environment has no cross-agent social-learning externality (agents are short-lived), but it does feature a payoff externality because effort responds to beliefs about the expert, which feeds back into the informativeness of outcomes. Halac et al. (2016) study optimal intertemporal contracts for experimentation with hidden information and hidden effort; we instead characterize reputational incentives when contracts are absent and show when reputation substitutes for explicit incentives. Dynamic signaling through experimentation with changing types is analyzed by Thomas (2019); our cutoff and martingale results are complementary and highlight how the endogenous effort response changes informativeness. Related work on information aggregation and transparency—such as Prat (2005) and more recent contributions like Backus and Little (2020)—suggests that disclosure regimes interact with reputational motives to induce conservatism; our mechanism microfounds this via effort-induced diagnosticity of outcomes.
Finally, markets for advice and compensation design link reputation with pay-for-performance. In financial advice and ratings, incentives and reputation interact in complex ways (Inderst and Ottaviani, 2012; Bolton et al., 2012). Bayesian persuasion (Kamenica and Gentzkow, 2011) provides tools for incentive-aligned disclosure; our one-shot analysis shows that a simple contingent-fee (success bonus) can implement the efficient cutoff by offsetting reputational distortions (Lukyanov et al., 2025). Motivating experimentation often requires tolerating early failures (Manso, 2011); our dynamic analysis shows how reputation can either replicate this tolerance (when effort is low at low reputation) or stall experimentation (when conservatism dominates at high reputation).
Relative to these literatures, our main novelty is to endogenize the implementer’s effort in a recursive reputation model and to characterize how this endogeneity—through a stochastic complementarity between reputation and effort—(i) yields a state-dependent cutoff with reputational conservatism, (ii) delivers clean martingale dynamics with boundary characterization, and (iii) generates transparent comparative statics in signal precision, priors, and patience. This mechanism is absent in reputational cheap talk without implementation effort and in bandit models without adviser career concerns.
3 Model
3.1 Environment and primitives
Time is discrete, $t=0,1,2,\dots$. A single expert (she) interacts in each period with a new short-lived agent (he). The expert’s type $\theta\in\{H,L\}$ (competent vs. less competent) is fixed over time and privately known to the expert. The market holds a public reputation (belief) $\pi_{t}\equiv\mathbb{P}(\theta=H\mid\mathcal{H}_{t})\in(0,1)$, where $\mathcal{H}_{t}$ is the public history up to the start of period $t$ (defined below). The initial reputation $\pi_{0}\in(0,1)$ is given.
In period $t$, the state $\omega_{t}\in\{0,1\}$ is drawn i.i.d. with $\mathbb{P}(\omega_{t}=1)=\lambda\in(0,1)$; throughout we write $\omega$ for a generic period draw. The expert privately observes a signal $s_{t}\in S\subseteq\mathbb{R}$ about $\omega_{t}$ before issuing a recommendation. For each type $\theta$, the signal distribution conditional on the state admits densities $f_{\theta}(\cdot\mid\omega)$ with respect to a common reference measure on $S$, satisfies the monotone likelihood ratio property (MLRP) in $s$, and the High type $H$ is (strictly) more informative than the Low type $L$ in the Blackwell sense. [^2] Formally, for $\theta\in\{H,L\}$ the likelihood ratio $\ell_{\theta}(s)\equiv f_{\theta}(s\mid 1)/f_{\theta}(s\mid 0)$ is strictly increasing in $s$, and $H$ Blackwell-dominates $L$, e.g. there exists a stochastic kernel $K$ with $f_{L}(\cdot\mid\omega)=\int K(\cdot\mid x)\,f_{H}(x\mid\omega)\,dx$ for each $\omega$. The binary-signal special case $S=\{0,1\}$ with accuracies $q_{H}>1/2\geq q_{L}$ is nested.
If the expert recommends the risky action ($a_{t}=1$), the agent chooses an effort level $e_{t}\in[0,1]$ at cost $c(e_{t})$, where $c:[0,1]\to\mathbb{R}_{+}$ is continuous, strictly increasing and strictly convex, with $c(0)=0$ and $c^{\prime}(0)=0$. If the expert recommends the safe action ($a_{t}=0$), no effort is exerted ($e_{t}\equiv 0$). The per-period outcome $y_{t}$ is publicly observed and equals $1$ (success) if and only if $a_{t}=1$, $\omega_{t}=1$, and the agent’s effort succeeds; we adopt the reduced-form success technology
$\mathbb{P}(y_{t}=1\mid a_{t}=1,\,\omega_{t},\,e_{t})=e_{t}\cdot\omega_{t},\qquad\mathbb{P}(y_{t}=1\mid a_{t}=0)=0.$
Thus success requires both a favorable state and implementation effort; failure $y_{t}=0$ is observed otherwise. [^3] We could allow a safe baseline payoff and a small amount of safe-outcome noise. In the baseline we take the safe outcome as uninformative about $\theta$ and normalize payoffs so that $y_{t}\in\{0,1\}$.
The agent’s period payoff is $y_{t}-c(e_{t})$. The expert derives a flow payoff $u(\pi_{t})$ from her current reputation (e.g. future demand or fees increasing in perceived competence) and discounts the future with factor $\delta\in(0,1)$. We allow an innocuous per-recommendation fee $\phi\geq 0$ if $a_{t}=1$; it can be set to zero without affecting the qualitative results.
3.2 Public histories and timing
Public histories are sequences $h_{t}=\big{(}(a_{0},y_{0}),\dots,(a_{t-1},y_{t-1})\big{)}\in\mathcal{H}_{t}\equiv(\{0,1\}\times\{0,1\})^{t}$ with the convention that when $a_{\tau}=0$ the associated $y_{\tau}$ is a publicly observed failure ($y_{\tau}=0$) and is uninformative about $\theta$. The public belief $\pi_{t}=\mathbb{P}(\theta=H\mid h_{t})$ is computed from Bayes’ rule given equilibrium strategies.
The within-period timing is:
  1. 1.
    $\pi_{t}$ is public; nature draws $\omega_{t}$; the expert privately observes $s_{t}\sim f_{\theta}(\cdot\mid\omega_{t})$.
  2. 2.
    The expert issues a recommendation $a_{t}\in\{0,1\}$ according to a strategy that may depend on $s_{t}$ and $\pi_{t}$.
  3. 3.
    The agent, observing $(a_{t},\pi_{t})$, forms interim beliefs about $\theta$ and $\omega_{t}$ (defined below) and, if $a_{t}=1$, chooses effort $e_{t}\in[0,1]$; if $a_{t}=0$, set $e_{t}=0$.
  4. 4.
    The outcome $y_{t}\in\{0,1\}$ is realized and observed; the public belief updates to $\pi_{t+1}$ by Bayes’ rule.
Table 1: Notation
Symbol Meaning
$\theta\in\{H,L\}$ Expert’s type: competent $H$ vs. less competent $L$
$\pi_{t}\in(0,1)$ Public reputation at start of period $t$; belief that $\theta=H$
$\omega_{t}\in\{0,1\}$ Period state (i.i.d. over $t$), $\mathbb{P}(\omega_{t}=1)=\lambda$
$s_{t}\in S$ Expert’s private signal about $\omega_{t}$
$a_{t}\in\{0,1\}$ Recommendation: risky $1$ vs. safe 0
$e_{t}\in[0,1]$ Agent’s implementation effort (chosen only if $a_{t}=1$)
$y_{t}\in\{0,1\}$ Outcome (public): success $1$ vs. failure 0
$c(\cdot)$ Effort cost
$u(\pi)$, $\delta$ Expert’s flow utility from reputation; discount factor $\delta\in(0,1)$
$\phi\geq 0$ Per-recommendation fee (relevant only if $a_{t}=1$; can be 0)
$r_{\theta}(a\mid\omega,\pi)$ Advice likelihood induced by type $\theta$ given $(\omega,\pi)$
$\pi^{\mathrm{rec}}(a;\pi)$ Posterior on $\theta$ after observing recommendation $a$
$\lambda(a,\pi)$ Posterior on $\omega=1$ after observing recommendation $a$
$e^{*}(a,\pi)$ Agent’s best-response effort
$P_{S}(1,\pi)$ Success probability conditional on $a=1$
$\pi^{+}(\pi),\;\pi^{-}(\pi)$ Post-outcome reputations after success/failure given $a=1$
$V(\pi)$ Expert’s continuation value at reputation $\pi$
$s^{*}(\pi)$ High-type risky-signal cutoff
$J^{+}(\pi),\;J^{-}(\pi)$ Log-likelihood jumps (success/failure) for the type posterior
$\rho(\pi),\;\Lambda(\pi)$ Recommendation and success intensities in the CT limit
3.3 Strategies and beliefs
A (behavioral) recommendation strategy for type $\theta$ is a measurable map
$\alpha_{\theta}:\;S\times(0,1)\to\Delta(\{0,1\}),\qquad a\mapsto\alpha_{\theta}(a\mid s,\pi),$
interpreted as the conditional probability of recommending $a$ after observing $(s,\pi)$. We restrict attention to belief-based Markov strategies that depend on the public history only via $\pi$; this is without loss for equilibria in our i.i.d. environment since $\pi$ is a sufficient statistic for continuation payoffs. [^4] A formal argument that $\pi$ is a payoff-relevant state follows from the i.i.d. state and public observability of $(a,y)$ together with the conditional independence of current signals given $(\theta,\omega_{t})$. See Appendix A for a statement and proof.
Given $(\alpha_{H},\alpha_{L})$ and $\pi$, define the advice likelihoods for each type and state
$r_{\theta}(a\mid\omega,\pi)\;\equiv\;\int_{S}\alpha_{\theta}(a\mid s,\pi)\,f_{\theta}(s\mid\omega)\,ds,\qquad\theta\in\{H,L\},\;\omega\in\{0,1\}.$
Unconditionally over the state,
$\mathbb{P}(a\mid\theta,\pi)=\lambda\,r_{\theta}(a\mid 1,\pi)+(1-\lambda)\,r_{\theta}(a\mid 0,\pi).$
The interim type belief (posterior on $\theta$) after observing recommendation $a$ is
$\pi^{\mathrm{rec}}(a;\pi)\;\equiv\;\mathbb{P}(\theta=H\mid a,\pi)\;=\;\frac{\pi\,\mathbb{P}(a\mid H,\pi)}{\pi\,\mathbb{P}(a\mid H,\pi)+(1-\pi)\,\mathbb{P}(a\mid L,\pi)}.$ (1)
The interim success belief (posterior on $\omega$) after observing recommendation $a$ is
$\lambda(a,\pi)\;\equiv\;\mathbb{P}(\omega=1\mid a,\pi)\;=\;\frac{\sum_{\theta\in\{H,L\}}\mathbb{P}(\theta\mid\pi)\,\lambda\,r_{\theta}(a\mid 1,\pi)}{\sum_{\theta\in\{H,L\}}\mathbb{P}(\theta\mid\pi)\big{[}\lambda\,r_{\theta}(a\mid 1,\pi)+(1-\lambda)\,r_{\theta}(a\mid 0,\pi)\big{]}}.$ (2)
In particular, $\lambda(a,\pi)$ is well defined whenever $\mathbb{P}(a\mid\pi)>0$; for off-path recommendations ($\mathbb{P}(a\mid\pi)=0$) beliefs are specified by standard refinements (we maintain Bayes-consistent selections throughout).
3.4 Agent’s best response
Given $(a,\pi)$, the agent chooses effort $e\in[0,1]$ to maximize expected net benefit
$e\;\in\;\arg\max_{x\in[0,1]}\;\lambda(a,\pi)\,x-c(x).$
By strict convexity of $c$, the solution is unique and characterized by the first-order condition
$c^{\prime}\big{(}e^{*}(a,\pi)\big{)}=\lambda(a,\pi),\qquad\text{hence }e^{*}(a,\pi)=(c^{\prime})^{-1}\!\big{(}\lambda(a,\pi)\big{)}\in[0,1],$ (3)
with $e^{*}(a,\pi)$ increasing in $\lambda(a,\pi)$. Under the common quadratic benchmark $c(x)=\tfrac{1}{2}x^{2}$, one has $e^{*}(a,\pi)=\lambda(a,\pi)$. [^5] General primitives for $c(\cdot)$ (strictly convex $C^{2}$ and effort caps) and robustness to outcome/observability frictions (baseline success under $a=0$, misclassification, partial observability of $a$) are developed in Online Appendix OA.4.
3.5 Outcome probabilities and belief updates
If $a=1$, the success probability (integrating over $\omega$ and effort) given $(\alpha_{H},\alpha_{L})$ and $\pi$ is
$P_{S}(1,\pi)\;=\;\lambda(1,\pi)\cdot e^{*}\big{(}1,\pi\big{)}.$ (4)
If $a=0$, we take outcomes as uninformative and set $P_{S}(0,\pi)=0$.
For $a=1$, the likelihood of outcomes conditional on the type is
$\mathbb{P}(y=1\mid\theta,a=1,\pi)$ $=e^{*}\big{(}1,\pi\big{)}\cdot\mathbb{P}(\omega=1\mid\theta,a=1,\pi)$
$=e^{*}\big{(}1,\pi\big{)}\cdot\frac{\lambda\,r_{\theta}(1\mid 1,\pi)}{\lambda\,r_{\theta}(1\mid 1,\pi)+(1-\lambda)\,r_{\theta}(1\mid 0,\pi)},$ (5)
$\mathbb{P}(y=0\mid\theta,a=1,\pi)$ $=1-\mathbb{P}(y=1\mid\theta,a=1,\pi).$
Let $\pi^{\mathrm{rec}}=\pi^{\mathrm{rec}}(1;\pi)$ denote the interim type posterior after observing $a=1$ but before observing $y$. The post-outcome reputations after success and failure, $\pi^{+}(\pi)$ and $\pi^{-}(\pi)$, satisfy the Bayes odds-ratio updates
$\begin{split}\frac{\pi^{+}}{1-\pi^{+}}&=\frac{\mathbb{P}(y=1\mid H,a=1,\pi)}{\mathbb{P}(y=1\mid L,a=1,\pi)}\cdot\frac{\pi^{\mathrm{rec}}}{1-\pi^{\mathrm{rec}}},\\ \frac{\pi^{-}}{1-\pi^{-}}&=\frac{\mathbb{P}(y=0\mid H,a=1,\pi)}{\mathbb{P}(y=0\mid L,a=1,\pi)}\cdot\frac{\pi^{\mathrm{rec}}}{1-\pi^{\mathrm{rec}}}.\end{split}$ (6)
If $a=0$, we set $\pi_{t+1}=\pi_{t}$ (no new information about $\theta$). [^6] Allowing a small safe-outcome signal about $\theta$ is straightforward and leaves our main results intact; we focus on the transparent benchmark where safe outcomes carry no $\theta$-information.
3.6 Expert’s problem and value function
Given a public reputation $\pi$, the expert anticipates the agent’s best response and the induced belief transitions. We consider Markov (belief-based) strategies. The expert’s continuation value $V(\pi)$ solves the Bellman equation
$V(\pi)\;=\;\max_{a\in\{0,1\}}\Big{\{}\,u(\pi)+\phi\,\mathbf{1}\{a=1\}+\delta\cdot\mathbb{E}\big{[}V(\pi^{\prime})\mid a,\pi\big{]}\,\Big{\}},$ (7)
where $\pi^{\prime}=\pi$ if $a=0$, and if $a=1$ then
$\mathbb{E}\big{[}V(\pi^{\prime})\mid a=1,\pi\big{]}=P_{S}(1,\pi)\,V\big{(}\pi^{+}(\pi)\big{)}+\big{(}1-P_{S}(1,\pi)\big{)}\,V\big{(}\pi^{-}(\pi)\big{)},$
with $P_{S}(1,\pi)$ as in (4) and $\pi^{\pm}(\pi)$ as in (6). A High-type expert’s strategy will be shown to admit a cutoff representation in $s$ (Theorem 5 below), while the Low type can mimic by suitable randomization to sustain on-path beliefs.
3.7 Equilibrium
We study belief-based Markov Perfect Bayesian Equilibria (MPBE).
Definition 1.
An MPBE consists of (i) recommendation strategies $(\alpha_{H},\alpha_{L})$ measurable in $(s,\pi)$, (ii) an agent best-response $e^{*}(a,\pi)$ as in (3), (iii) a value function $V(\pi)$, and (iv) a belief system $(\pi^{\mathrm{rec}}(a;\pi),\,\lambda(a,\pi),\,\pi^{\pm}(\pi))$ defined by (1)–(6), such that:
  1. 1.
    Expert optimality. For each type $\theta$ and each $\pi\in(0,1)$, $\alpha_{\theta}(\cdot\mid s,\pi)$ puts probability one on the set of $a\in\{0,1\}$ that maximize the type-$\theta$ Bellman objective (7) given $e^{*}$ and the belief system.
  2. 2.
    Agent optimality. For each $(a,\pi)$, $e^{*}(a,\pi)$ solves (3).
  3. 3.
    Belief consistency. On-path beliefs are given by Bayes’ rule as in (1)–(6). Off-path beliefs are specified to satisfy standard consistency.
3.8 Regularity assumptions
We impose the following standing assumptions, used throughout:
  • A1
    $S$ is an interval of $\mathbb{R}$; for each $\theta\in\{H,L\}$ and $\omega\in\{0,1\}$, $f_{\theta}(\cdot\mid\omega)$ exists, is continuous in $s$, and satisfies MLRP; $H$ is (strictly) more informative than $L$ in the Blackwell order.
  • A2
    The cost $c$ is continuously differentiable, strictly increasing and strictly convex on $[0,1]$, with $c(0)=0$ and $c^{\prime}(0)=0$.
  • A3
    The flow utility $u:(0,1)\to\mathbb{R}$ is continuous and weakly increasing; $\delta\in(0,1)$.
  • A4
    The initial belief $\pi_{0}\in(0,1)$ and the good-state prior $\lambda\in(0,1)$ are common knowledge.
3.9 Remarks and special cases
The binary-signal case $S=\{0,1\}$ with accuracies $q_{H}>1/2\geq q_{L}$ provides a transparent benchmark: the High type recommends risk if and only if $s=1$ when $\pi$ is sufficiently low, and adopts a stricter rule as $\pi$ rises; the Low type randomizes to match frequencies on-path. Under quadratic cost $c(e)=\tfrac{1}{2}e^{2}$, the agent’s effort is $e^{*}(a,\pi)=\lambda(a,\pi)$ and the success probability for $a=1$ is $\lambda(1,\pi)]^{2}$, which highlights how higher reputation raises effort and thereby increases the diagnostic content of outcomes. The general continuous-signal case is handled by our MLRP/Blackwell assumptions and delivers a cutoff policy in $s$ for the High type.
In the next section we establish the equilibrium cutoff structure (Theorem 5), show that the cutoff is (weakly) increasing in $\pi$ (reputational conservatism), and characterize the resulting reputation dynamics and comparative statics.
4 Core Results
This section establishes the equilibrium structure and the main comparative statics. Throughout we maintain Assumptions A1–A4 from Section 3. Proofs that are longer or purely technical are deferred to the appendices as indicated. [^7] Algorithms and replication details for all figures and numerical objects appear in Online Appendix OA.3.
4.1 Reputation–effort feedback
We begin by formalizing the key complementarity between reputation and implementation effort.
Proposition 2.
Fix a belief-based Markov strategy profile $(\alpha_{H},\alpha_{L})$ and $\pi\in(0,1)$. Let $\pi^{\mathrm{rec}}(a;\pi)$ be the interim type belief after observing a recommendation $a\in\{0,1\}$, and let $e^{*}(a,\pi)$ be the agent’s best response. Then $e^{*}(1,\pi)$ is strictly increasing in $\pi^{\mathrm{rec}}(1;\pi)$, and $\pi^{\mathrm{rec}}(1;\pi)$ is (weakly) increasing in $\pi$. Consequently, the conditional success probability after a risky recommendation,
$P_{S}(1,\pi)\;=\;\lambda(1,\pi)\cdot e^{*}(1,\pi),$
is (weakly) increasing in the public reputation $\pi$.
Proof.
By (3), $c^{\prime}\big{(}e^{*}(1,\pi)\big{)}=\lambda(1,\pi)$ and $c^{\prime\prime}>0$, hence $e^{*}(1,\pi)$ is strictly increasing in the interim success belief $\lambda(1,\pi)$. By Bayes’ rule, $\lambda(1,\pi)$ is strictly increasing in the interim type belief $\pi^{\mathrm{rec}}(1;\pi)$ (a higher probability of $H$ raises the weight on the more informative recommendation distribution and thus the posterior that $\omega=1$ under MLRP). Finally, $\pi^{\mathrm{rec}}(1;\pi)$ is (weakly) increasing in $\pi$ since it is a Bayes posterior in a binary-mixture prior. The conclusion on $P_{S}(1,\pi)$ follows from (4). ∎
4.2 Cutoff structure in the private signal
Let the High type’s risky-minus-safe continuation difference at signal $s$ and reputation $\pi$ be
$\Delta_{H}(s;\pi)\;\equiv\;\Big{\{}u(\pi)+\phi+\delta\cdot\mathbb{E}\big{[}V(\pi^{\prime})\mid a=1,s,\pi\big{]}\Big{\}}\;-\;\Big{\{}u(\pi)+\delta\cdot V(\pi)\Big{\}}.$ (8)
By construction, recommending $a=1$ is optimal at $(s,\pi)$ if and only if $\Delta_{H}(s;\pi)\geq 0$.
Lemma 3.
Under Assumption A1 and for any fixed $\pi\in(0,1)$, the mapping $s\mapsto\Delta_{H}(s;\pi)$ is (strictly) increasing. In particular, there exists a (possibly weak) cutoff $s^{*}(\pi)\in\overline{S}$ such that the High type’s optimal recommendation is $a=1$ if and only if $s\geq s^{*}(\pi)$.
Proof.
By (8) and (4),
$\mathbb{E}\big{[}V(\pi^{\prime})\mid a=1,s,\pi\big{]}=P_{S}(s;\pi)\cdot V\big{(}\pi^{+}(\pi)\big{)}+\big{(}1-P_{S}(s;\pi)\big{)}\cdot V\big{(}\pi^{-}(\pi)\big{)},$
where $P_{S}(s;\pi)$ is the success probability given $(s,\pi)$ and $a=1$. Assumption A1 (MLRP and that $H$ is more informative than $L$) implies that the posterior $\mathbb{P}(\omega=1\mid s,a=1,\pi)$ is strictly increasing in $s$. By Proposition 2, the agent’s effort is strictly increasing in that posterior, hence $P_{S}(s;\pi)$ is strictly increasing in $s$. As $V(\pi^{+})>V(\pi^{-})$ (since $V$ is increasing and $\pi^{+}>\pi^{-}$ by Bayes’ rule), it follows that $\mathbb{E}[V(\pi^{\prime})\mid a=1,s,\pi]$ is strictly increasing in $s$, and thus so is $\Delta_{H}(s;\pi)$. The threshold property follows by monotone selection on $\{s:\Delta_{H}(s;\pi)\geq 0\}$. ∎
Theorem 4.
Under (A1)–(A4) and $\delta\in(0,1)$, the expert’s value function $V:[0,1]\to\mathbb{R}$ is increasing in $\pi$. Moreover, if Condition (6) (failure is at least as diagnostic as success at the cutoff) holds, then $V$ is convex in $\pi$. In particular, for any fixed Markov cutoff policy $s(\cdot)$, the associated Bellman operator $T^{\,s}$ is a monotone contraction that preserves convexity; hence its unique fixed point $V^{\,s}$ is increasing and convex. At the equilibrium cutoff policy $s^{*}(\cdot)$, $V=V^{\,s^{*}}$ inherits these properties.
Proof (short). Monotonicity: For any bounded $V$, the map $\pi\mapsto\phi\,\rho(\pi)+\delta\,\mathbb{E}[V(\pi^{\prime})\mid\pi]$ is increasing because posteriors are increasing in prior odds under MLRP (A1) and $V$ is evaluated at larger posteriors in the first-order stochastic dominance sense. Contraction holds with modulus $\delta$ in the sup norm. For convexity, fix a cutoff policy $s(\cdot)$. Under $a=1$, the posterior takes $\{\pi^{+}(\pi),\pi^{-}(\pi)\}$ with probabilities that shift toward the failure branch as reputation (and thus effort) rises; Condition (6) ensures the induced posterior kernel becomes a mean-preserving spread in the convex order as $\pi$ increases. Hence $\pi\mapsto\mathbb{E}[V(\pi^{\prime})\mid\pi]$ is convex for convex $V$, and so is $T^{\,s}V$. The fixed point $V^{\,s}$ is therefore increasing and convex; the equilibrium value $V=V^{\,s^{*}}$ inherits the properties. Full details appear in Appendix OA.1. ∎
Theorem 5.
For each $\pi\in(0,1)$, in any MPBE there exists a measurable cutoff $s^{*}(\pi)$ such that the High type recommends $a=1$ if and only if $s\geq s^{*}(\pi)$. The Low type can mimic by a (signal-measurable) mixed strategy that matches the induced recommendation frequencies so that Bayes’ rule applies on-path.
Proof.
Lemma 3 yields existence of a cutoff for the High type. Given this cutoff and Assumption A1, one constructs the Low type’s mixed strategy by equating the induced probabilities of $a=1$ conditional on $\omega$ to ensure well-defined likelihood ratios for on-path belief updates; see Appendix A.4 for existence of an MPBE by a monotone fixed-point argument. ∎
4.3 Reputational conservatism
We now establish that the cutoff is (weakly) increasing in the public reputation: more highly reputed experts are more conservative.
Condition 6.
Let $L^{+}(\pi)\equiv\frac{r_{H}(1\mid 1,\pi)}{r_{L}(1\mid 1,\pi)}$ and $L^{-}(\pi)\equiv\frac{r_{H}(1\mid 0,\pi)}{r_{L}(1\mid 0,\pi)}$ denote the outcome likelihood-ratio jumps after a risky recommendation at public reputation $\pi$, evaluated at the equilibrium cutoff $s^{*}(\pi)$. We assume
$\log L^{+}(\pi)\;\leq\;-\log L^{-}(\pi),$
with strict inequality on a set of $\pi$ of positive measure. Equivalently, failures are at least as diagnostic of type as successes (in log-likelihood terms) near the cutoff.
Theorem 7.
Suppose Assumptions A1–A3 hold and $V$ is increasing and convex on $(0,1)$. Then for any $\pi^{\prime}<\pi^{\prime\prime}$ in $(0,1)$, one has $s^{*}(\pi^{\prime})\leq s^{*}(\pi^{\prime\prime})$.
Proof outline.
Fix $\pi$ and write $\Delta_{H}(s;\pi)$ as in (8). The derivative of $\Delta_{H}$ with respect to $\pi$ at fixed $s$ has two effects: (i) a diagnosticity effect via $P_{S}(s;\pi)$, and (ii) a baseline effect via $V(\cdot)$. By Proposition 2, a higher $\pi$ raises the interim effort and therefore increases the informativeness of outcomes; in particular, the gap $V(\pi^{+})-V(\pi^{-})$ is multiplied by a larger weight on the bad-news realization when failure occurs (since failure is more likely to be attributed to a bad signal/action when clients worked hard), which reduces the expected continuation value of taking risk, ceteris paribus. Convexity of $V$ implies diminishing returns to further increases in $\pi$, so the marginal gain from success is smaller than the marginal loss from failure around high $\pi$. Formally, one shows that $\partial_{\pi}\Delta_{H}(s;\pi)\leq 0$ for all $s$; hence the crossing point $\{s:\Delta_{H}(s;\pi)=0\}$ weakly increases in $\pi$. A complete argument is provided in Appendix B, which establishes the monotone comparative statics using supermodularity of the Bellman operator and Bayes-likelihood monotonicity induced by MLRP. ∎
Corollary 8.
If $s^{*}(\pi)$ is increasing in $\pi$, then for any two reputations $\pi^{\prime}<\pi^{\prime\prime}$ and any signal $s$ with $s^{*}(\pi^{\prime})\leq s<s^{*}(\pi^{\prime\prime})$, the High type recommends $a=1$ at $\pi^{\prime}$ but $a=1$ is not recommended at $\pi^{\prime\prime}$. Thus, as reputation falls, the expert is (weakly) more willing to endorse risk on intermediate signals.
4.4 Binary-signal worked example
To make Condition 6 and its implications transparent, consider a binary signal $s\in\{h,\ell\}$ with symmetric accuracies. For type $\theta\in\{H,L\}$ and state $\omega\in\{0,1\}$,
$\mathbb{P}(s=h\mid\omega=1,\theta)=q_{\theta},\qquad\mathbb{P}(s=h\mid\omega=0,\theta)=1-q_{\theta},$
with $q_{H}>q_{L}\in(1/2,1)$ so that $H$ is more informative in the MLRP sense. The advice policy is a (degenerate) cutoff: $a=1$ if $s=h$ and $a=0$ if $s=\ell$, unless the risky–safe advantage at $h$ is negative, in which case the expert never recommends risk.
Posterior odds shifts. Let $\mathrm{odds}(\pi)=\pi/(1-\pi)$ and $\mathrm{odds}^{-1}(x)=x/(1+x)$. When $a=1$ is recommended at $s=h$, the outcome-based likelihood-ratio jumps are
$L^{+}\;=\;\frac{r_{H}(1\mid 1)}{r_{L}(1\mid 1)}=\frac{q_{H}}{q_{L}}\;>\;1,\qquad L^{-}\;=\;\frac{r_{H}(1\mid 0)}{r_{L}(1\mid 0)}=\frac{1-q_{H}}{\,1-q_{L}\,}\;<\;1,$
so that, conditional on $a=1$, the posterior odds update as
$\mathrm{odds}(\pi^{\prime})=\begin{cases}\mathrm{odds}(\pi)\,L^{+},&y=1,\\[2.0pt] \mathrm{odds}(\pi)\,L^{-},&y=0.\end{cases}$
Thus $\pi^{+}(\pi)=\mathrm{odds}^{-1}(\mathrm{odds}(\pi)L^{+})$ and $\pi^{-}(\pi)=\mathrm{odds}^{-1}(\mathrm{odds}(\pi)L^{-})$.
Diagnosticity asymmetry (Condition 6). In this binary benchmark,
$\log L^{+}+\log L^{-}\;=\;\log\!\left(\frac{q_{H}}{q_{L}}\cdot\frac{1-q_{H}}{1-q_{L}}\right)\;\leq\;0\quad\Longleftrightarrow\quad q_{H}(1-q_{H})\;\leq\;q_{L}(1-q_{L}).$
Hence Condition 6 reduces to $q_{H}(1-q_{H})\leq q_{L}(1-q_{L})$, which holds whenever the high type is sufficiently precise (i.e., $q_{H}$ close enough to $1$ relative to $q_{L}$). Intuitively, at the policy cutoff the failure branch ($y=0$) moves posterior odds farther (down) than the success branch ($y=1$) moves them up.
Cutoff in reputation and conservatism. Let $\Delta_{H}(s;\pi)$ denote the high type’s risky–safe advantage at signal $s$ (defined in Section 4). Because $\Delta_{H}(h;\pi)$ strictly decreases in $\pi$ under Condition 6 (decreasing differences), there exists at most one reputation threshold $\bar{\pi}\in(0,1)$ solving
$\Delta_{H}(h;\bar{\pi})=0.$
For $\pi<\bar{\pi}$, the high type recommends risk iff $s=h$; for $\pi>\bar{\pi}$, she recommends safe even at $s=h$ (maximal conservatism). In typical parameter regions $\Delta_{H}(\ell;\pi)<0$ for all $\pi$, so the only relevant margin is the $h$-signal; if $\Delta_{H}(\ell;\pi)$ becomes nonnegative at very low $\pi$, the risky region expands (weakening conservatism) in the direction predicted by Theorem 7.
Recommendation-only update. If observers also update on the recommendation itself, the recommendation-only likelihood ratio is
$L^{\mathrm{rec}}\;=\;\frac{\mathbb{P}(a=1\mid H)}{\mathbb{P}(a=1\mid L)}=\frac{\alpha\,q_{H}+(1-\alpha)(1-q_{H})}{\alpha\,q_{L}+(1-\alpha)(1-q_{L})}\,$,
so that $\tilde{\pi}(\pi)=\mathrm{odds}^{-1}\!\big{(}\mathrm{odds}(\pi)\,L^{\mathrm{rec}}\big{)}$. The overall continuation term in $\Delta_{H}(h;\pi)$ then combines $\tilde{\pi}(\pi)$ and the two outcome branches $\pi^{\pm}(\pi)$ as in Section 4. Since implementer effort $e^{*}(1,\pi)$ rises with $\pi$, the success probability $P_{S}(h;\pi)$ increases with $\pi$, which reinforces the asymmetry by making failures more diagnostic when reputation is high.
Interpretation (medicine). Think of $s=h$ as strong clinical/imaging evidence favoring surgery. A highly reputed surgeon ($\pi$ high) elicits greater adherence/effort from the care team and patient, so an operation that fails is especially revealing (large $|\log L^{-}|$), whereas a success is less incrementally informative. The policy implication is conservative: at high standing, the bar for recommending surgery rises.
4.5 Comparative statics: precision, prior, and patience
We now promote three comparative statics to main-text propositions. Proofs are short and rely on strict single crossing of $\Delta_{H}(s;\pi)$ in $s$; full derivative formulas for the Gaussian benchmark appear in the Online Appendix (OA–C).
Proposition 9.
Fix $\pi\in(0,1)$ and primitives (A1)–(A4). If the high type’s signal becomes more informative in the Blackwell sense and/or the low type’s becomes weakly less informative (holding $\alpha,\pi$ fixed), then the risky cutoff $s^{*}(\pi)$ (weakly) decreases. In the Gaussian benchmark $s\mid(\theta,\omega)\sim\mathcal{N}(\mu_{\omega},\sigma_{\theta}^{2})$ with $\sigma_{L}>\sigma_{H}$, one has $\partial s^{*}/\partial(\mu_{1}-\mu_{0})<0$, $\partial s^{*}/\partial\sigma_{H}>0$, and $\partial s^{*}/\partial\sigma_{L}>0$ whenever $s^{*}(\pi)\in(\inf S,\sup S)$.
Proof.
Let $G(s;t)\equiv\Delta_{H}(s;\pi;t)$ with parameter $t$ indexing informativeness (higher $t$ = more informative $H$, less informative $L$). By MLRP/Blackwell dominance, at any fixed $s$ the outcome LLRs satisfy $L^{+}(t)\uparrow$ and $L^{-}(t)\downarrow$; hence the continuation term in $\Delta_{H}$ increases pointwise in $s$, so $\partial G/\partial t>0$. Strict single crossing gives $\partial G/\partial s>0$ at $s^{*}(\pi)$. The implicit-function formula yields
$\frac{\partial s^{*}}{\partial t}=-\frac{\partial G/\partial t}{\partial G/\partial s}\;<\;0.$
The Gaussian signs follow from the normal tail derivatives of $r_{\theta}(1\mid\omega)$ in $(\mu_{1}-\mu_{0},\sigma_{H},\sigma_{L})$. ∎
Proposition 10.
Holding $(\pi,\text{signals})$ fixed, $s^{*}(\pi)$ is (weakly) decreasing in the prior $\alpha\equiv\mathbb{P}(\omega=1)$.
Proof.
Write $\Delta_{H}(s;\pi)=\phi+\delta\{\mathbb{E}[V(\pi^{\prime})\mid a{=}1,s]-\mathbb{E}[V(\pi^{\prime})\mid a{=}0]\}$. As $\alpha$ increases, (a=1,y=1)(a{=}1,y{=}1) histories become more likely under both types but relatively more under $H$, so $L^{+}$ rises and $L^{-}$ falls (or falls less). Thus $\Delta_{H}$ increases pointwise in $s$, i.e., $\partial\Delta_{H}/\partial\alpha>0$. With $\partial\Delta_{H}/\partial s>0$ at $s^{*}$, the implicit-function formula gives $\partial s^{*}/\partial\alpha<0$. ∎
Proposition 11.
Fix $\pi$ and suppose $\phi\geq 0$ and (A1)–(A4) hold. Then $s^{*}(\pi)$ is (weakly) increasing in the discount factor $\delta$. Moreover, if $V_{2}$ is a mean-preserving spread of $V_{1}$ (both increasing), then $s^{*}_{V_{2}}(\pi)\geq s^{*}_{V_{1}}(\pi)$.
Proof.
At $s^{*}$, $\Delta_{H}(s^{*};\pi)=0=\phi+\delta\,\Psi(s^{*};\pi)$ with $\Psi\equiv\mathbb{E}[V(\pi^{\prime})]-V(\pi)$. Hence $\Psi(s^{*};\pi)=-\phi/\delta\leq 0$. Holding $s$ fixed, $\partial\Delta_{H}/\partial\delta=\Psi\leq 0$, so by the implicit-function formula and $\partial\Delta_{H}/\partial s>0$ we get $\partial s^{*}/\partial\delta\geq 0$. For curvature, replacing $V$ by a mean-preserving spread lowers $\Psi$ (Jensen gap more negative under the diagnosticity asymmetry), thus lowers $\Delta_{H}$ pointwise in $s$, which raises the smallest $s$ solving $\Delta_{H}\geq 0$. ∎
Remarks. (i) The patience effect is strict when $\phi>0$ (the flow benefit of recommending risk), because then $\Psi(s^{*};\pi)<0$. (ii) In the Gaussian case, OA–C reports closed-form derivatives of the normal tail terms and the bonus-induced shifts used in policy calibration.
4.6 Reputation dynamics
We next characterize the belief process $\{\pi_{t}\}_{t\geq 0}$ along the equilibrium path.
Theorem 12.
Fix an MPBE with cutoff policy $s^{*}(\cdot)$ and let $\{\pi_{t}\}$ be the induced reputation process. Then:
  1. 1.
    If $\theta=H$, $(\pi_{t})_{t\geq 0}$ is a submartingale with respect to the public filtration:
    $\mathbb{E}[\pi_{t+1}\mid\mathcal{H}_{t}]=\pi_{t}\quad\text{and}\quad\mathbb{P}(\pi_{t+1}\geq\pi_{t}\mid\mathcal{H}_{t})>0$
    whenever $a_{t}=1$ occurs with positive probability.
  2. 2.
    If $\theta=L$, $(\pi_{t})_{t\geq 0}$ is a supermartingale.
  3. 3.
    There exist $0<\underline{\pi}<\overline{\pi}<1$ (depending on primitives) such that
    $\mathbb{P}\left(\liminf_{t\to\infty}\pi_{t}\leq\underline{\pi}\text{ or }\limsup_{t\to\infty}\pi_{t}\geq\overline{\pi}\right)=1.$
    In particular, with probability one the process hits a high-trust or low-trust region.
Proof sketch.
Parts (1)–(2) follow from Bayes’ rule and the law of iterated expectations: conditional on the true type, the posterior is a (super/sub)martingale. Part (3) uses that when $a_{t}=1$ occurs with positive probability infinitely often, the innovation in the likelihood ratio has nontrivial variance (Appendix C shows a Doob decomposition for the log-odds process and applies optional stopping). When $a_{t}=0$ along long stretches, the belief remains constant; however, under our cutoff structure and Assumption A1, either the process eventually visits the risky region infinitely often (yielding learning) or it remains trapped in a no-news region that is separated from the interior by the thresholds, which implies eventual absorption in a low- or high-trust basin. Full details are provided in Appendix C. ∎
Remark 13.
If agents stop consulting the expert when $\pi_{t}<\underline{\pi}$ (endogenous exit), then $\pi_{t}$ is absorbed at $\underline{\pi}$ with positive probability. Conversely, if failures at very high $\pi$ do not fully overturn beliefs due to prior mass or noisy safe outcomes, then $\overline{\pi}<1$ can be an attracting region. These cases are covered by the boundary characterization in Appendix C.
To visualize Theorem 12 and the boundary behavior discussed above, Figure 1 plots simulated posterior paths $\{\pi_{t}\}$ in the Gaussian–quadratic benchmark. The top panel shows trajectories when the true type is $H$; the bottom panel shows $L$. Flat segments correspond to periods with $a=0$ (no experimentation), whereas jumps occur only after a risky recommendation $(a=1)$ and the ensuing outcome. The calibration is chosen for transparency rather than fit; qualitative patterns are robust.
Figure 1: Simulated reputation dynamics in the Gaussian benchmark. Baseline parameters $(\mu_{0},\mu_{1},\sigma_{H},\sigma_{L},\lambda,\delta)=(0,1,1,1.7,0.5,0.9)$. Each line is one replication starting from $\pi_{0}=0.5$. Flat segments arise when $a=0$; jumps reflect $(a=1)$ and the realized outcome.
The figure illustrates three predictions of the model. First, under $\theta=H$ the process drifts upward and hits the high-trust region with high probability, while under $\theta=L$ it drifts downward (Theorem 12). Second, learning occurs only when the expert recommends the risky action: flat stretches reflect $a=0$, and the process advances by likelihood-ratio jumps after $a=1$ (the visible steps in both panels). Third, jumps down following failures are larger at high reputation than at low reputation because implementers exert more effort when $\pi$ is high; by contrast, the informational content of success does not increase with effort. This asymmetry is the microfoundation of our reputational conservatism result (Theorem 7) and the boundary-hitting behavior in the dynamics.
4.7 Comparative statics: precision, prior, and patience
We now promote three comparative statics to main-text propositions. Proofs are short and rely on strict single crossing of $\Delta_{H}(s;\pi)$ in $s$; full derivative formulas for the Gaussian benchmark appear in the Online Appendix (OA–C).
Proposition 14.
Fix $\pi\in(0,1)$ and primitives (A1)–(A4). If the high type’s signal becomes more informative in the Blackwell sense and/or the low type’s becomes weakly less informative (holding $\alpha,\pi$ fixed), then the risky cutoff $s^{*}(\pi)$ (weakly) decreases. In the Gaussian benchmark $s\mid(\theta,\omega)\sim\mathcal{N}(\mu_{\omega},\sigma_{\theta}^{2})$ with $\sigma_{L}>\sigma_{H}$, one has $\partial s^{*}/\partial(\mu_{1}-\mu_{0})<0$, $\partial s^{*}/\partial\sigma_{H}>0$, and $\partial s^{*}/\partial\sigma_{L}>0$ whenever $s^{*}(\pi)\in(\inf S,\sup S)$.
Proof.
Let $G(s;t)\equiv\Delta_{H}(s;\pi;t)$ with parameter $t$ indexing informativeness (higher $t$ = more informative $H$, less informative $L$). By MLRP/Blackwell dominance, at any fixed $s$ the outcome LLRs satisfy $L^{+}(t)\uparrow$ and $L^{-}(t)\downarrow$; hence the continuation term in $\Delta_{H}$ increases pointwise in $s$, so $\partial G/\partial t>0$. Strict single crossing gives $\partial G/\partial s>0$ at $s^{*}(\pi)$. The implicit-function formula yields
$\frac{\partial s^{*}}{\partial t}=-\frac{\partial G/\partial t}{\partial G/\partial s}\;<\;0.$
The Gaussian signs follow from the normal tail derivatives of $r_{\theta}(1\mid\omega)$ in $(\mu_{1}-\mu_{0},\sigma_{H},\sigma_{L})$. ∎
Proposition 15.
Holding $(\pi,\text{signals})$ fixed, $s^{*}(\pi)$ is (weakly) decreasing in the prior $\alpha\equiv\mathbb{P}(\omega=1)$.
Proof.
Write $\Delta_{H}(s;\pi)=\phi+\delta\{\mathbb{E}[V(\pi^{\prime})\mid a{=}1,s]-\mathbb{E}[V(\pi^{\prime})\mid a{=}0]\}$. As $\alpha$ increases, (a=1,y=1)(a{=}1,y{=}1) histories become more likely under both types but relatively more under $H$, so $L^{+}$ rises and $L^{-}$ falls (or falls less). Thus $\Delta_{H}$ increases pointwise in $s$, i.e., $\partial\Delta_{H}/\partial\alpha>0$. With $\partial\Delta_{H}/\partial s>0$ at $s^{*}$, the implicit-function formula gives $\partial s^{*}/\partial\alpha<0$. ∎
Proposition 16.
Fix $\pi$ and suppose $\phi\geq 0$ and (A1)–(A4) hold. Then $s^{*}(\pi)$ is (weakly) increasing in the discount factor $\delta$. Moreover, if $V_{2}$ is a mean-preserving spread of $V_{1}$ (both increasing), then $s^{*}_{V_{2}}(\pi)\geq s^{*}_{V_{1}}(\pi)$.
Proof.
At $s^{*}$, $\Delta_{H}(s^{*};\pi)=0=\phi+\delta\,\Psi(s^{*};\pi)$ with $\Psi\equiv\mathbb{E}[V(\pi^{\prime})]-V(\pi)$. Hence $\Psi(s^{*};\pi)=-\phi/\delta\leq 0$. Holding $s$ fixed, $\partial\Delta_{H}/\partial\delta=\Psi\leq 0$, so by the implicit-function formula and $\partial\Delta_{H}/\partial s>0$ we get $\partial s^{*}/\partial\delta\geq 0$. For curvature, replacing $V$ by a mean-preserving spread lowers $\Psi$ (Jensen gap more negative under the diagnosticity asymmetry), thus lowers $\Delta_{H}$ pointwise in $s$, which raises the smallest $s$ solving $\Delta_{H}\geq 0$. ∎
Remarks. (i) The patience effect is strict when $\phi>0$ (the flow benefit of recommending risk), because then $\Psi(s^{*};\pi)<0$. (ii) In the Gaussian case, OA–C reports closed-form derivatives of the normal tail terms and the bonus-induced shifts used in policy calibration.
4.8 Empirical predictions and measurement
This subsection translates our theory into testable predictions and clarifies how to measure the key objects in data. We focus on settings where outcomes and recommendations are observed and where implementer effort can be proxied (e.g., adherence/compliance in surgery and conservative care Haskard Zolnierek and DiMatteo (2009); Birkhäuer et al. (2017), surgeon skill/volume Birkmeyer et al. (2003); Birkmeyer et al. (2013), or trade size/turnover in finance Loh and Stulz (2011)).
4.8.1 Risk taking falls with reputation
By Theorem 7, the risky-signal cutoff is (weakly) increasing in current reputation, so the probability of a risky recommendation declines with reputation. A reduced-form test is
$\mathbb{P}(a_{it}=1)=\alpha+\beta\,\text{Rep}_{it-1}+\gamma^{\prime}X_{it}+\eta_{i}+\tau_{t}+\varepsilon_{it},$
with $\beta<0$, expert fixed effects $\eta_{i}$, and time (or case-mix) controls $X_{it}$ and $\tau_{t}$. In surgical data, $a_{it}=1$ indicates operate vs. conservative care; $\text{Rep}_{it-1}$ can be constructed from leave-one-out success rates or Bayesian smoothed scores using only past outcomes (excluding period $t$). In analyst data, $a_{it}=1$ is a buy/upgrade vs. hold/downgrade.
4.8.2 Conditional success rises with reputation
Because implementers work harder when reputation is high, the success probability conditional on $a=1$ increases with reputation (Proposition 2 and Theorem 12). Estimate
$\mathbb{P}(y_{it}=1\mid a_{it}=1)=\alpha+\theta\,\text{Rep}_{it-1}+\gamma^{\prime}X_{it}+\eta_{i}+\tau_{t}+\varepsilon_{it},$
with $\theta>0$. In medicine, success is a composite endpoint (e.g., no 30-day readmission/complication); in finance it can be excess return around recommendation implementation or forecast accuracy; effort proxies include adherence measures, therapy attendance, medication refill behavior, workflow timestamps, or trade size/execution quality.
4.8.3 Failures are more damaging at high reputation
Theorem 12 and the likelihood-ratio calculus imply that the reputational drop after a failure is larger when reputation is high, while the reputational gain after success is comparatively stable. Let $\Delta\text{Rep}_{it}\equiv\text{Rep}_{it}-\text{Rep}_{it-1}$ be the ex post revision from an outcome following $a=1$. Estimate
$\Delta\text{Rep}_{it}=\alpha+\phi_{1}\,\mathbf{1}\{y_{it}=1\}+\phi_{0}\,\mathbf{1}\{y_{it}=0\}+\psi\,\text{Rep}_{it-1}\times\mathbf{1}\{y_{it}=0\}+\gamma^{\prime}X_{it}+\eta_{i}+\tau_{t}+\varepsilon_{it},$
predicting $\psi<0$. This can be implemented by constructing a transparent, data-driven reputation metric (see “Measurement” below) and running an event-study around outcome realizations.
4.8.4 Comparative statics.
Theorem 17 implies: (i) higher private-signal precision or a higher baseline viability increases experimentation; (ii) greater patience amplifies conservatism at high reputation. Proxy precision by the dispersion/variance of pre-decision diagnostics (imaging quality scores, lab-panel informativeness) or by analyst coverage informativeness; proxy viability by baseline risk scores. Test shifts in the level and slope of $\mathbb{P}(a=1)$ in difference-in-differences designs around plausibly exogenous changes in information technology or guidelines.
The paths in Figure 1 illustrate these predictions: upward-drifting, jumpy trajectories under competence and downward-drifting ones under incompetence, with flat stretches when the expert forgoes risk and larger downward jumps after failures at high reputation. These patterns are the visual counterparts of P1–P3 and provide a guide for specification diagnostics in empirical work.
4.9 Comparative statics
We study how primitives shift the cutoff policy and the induced experimentation and learning.
Theorem 17.
Let $s^{*}(\pi;\cdot)$ denote the equilibrium cutoff as a function of primitives. Then:
  1. 1.
    Signal precision. If the High type’s signal becomes more informative in the Blackwell order (holding the Low type fixed), then $s^{*}(\pi)$ decreases pointwise in $\pi$; the expert recommends risk on a (weakly) larger set of signals.
  2. 2.
    Good-state prior $\lambda$. The cutoff $s^{*}(\pi)$ is (weakly) decreasing in $\lambda$.
  3. 3.
    Patience $\delta$. If $u$ is sufficiently convex and $V$ inherits convexity, then $s^{*}(\pi)$ is (weakly) increasing in $\delta$.
Proof outline.
(1) A Blackwell improvement shifts the posterior $\mathbb{P}(\omega=1\mid s)$ upward in the monotone likelihood ratio order for each $s$, hence raises $P_{S}(s;\pi)$ and the expected value of taking risk; the single-crossing Lemma 3 implies the cutoff weakly falls. (2) A higher prior $\lambda$ uniformly increases $\mathbb{P}(\omega=1\mid s)$ and thus $P_{S}(s;\pi)$, again lowering the cutoff. (3) With greater patience, the expert puts more weight on the reputational consequences; under convexity of $V$, the downside from failure grows faster in $\delta$ than the upside from success at high $\pi$, shifting $\Delta_{H}(s;\pi)$ downward and raising the cutoff. Appendix D contains full statements and proofs, including conditions under which (3) holds globally and examples (quadratic cost; logistic signals) in which the inequalities are strict. ∎
4.10 Welfare and experimentation rates
The cutoff policy $s^{*}(\pi)$ determines an experimentation rate at reputation $\pi$, namely $\rho(\pi)\equiv\mathbb{P}(s\geq s^{*}(\pi)\mid\pi,\theta=H)$ on-path. Combining Theorems 7 and 17 yields:
Corollary 18.
Experimentation rates are (weakly) decreasing in $\pi$ and (weakly) increasing in signal precision and in $\lambda$. The expected one-step Kullback–Leibler information gain about $\theta$ after a risky recommendation is increasing in $\pi$ via Proposition 2, but total information accumulation over time can be nonmonotone in $\pi$ because high $\pi$ reduces the frequency of experimentation.
Proof.
Immediate from the monotonicity of $s^{*}(\pi)$ in $\pi$ and primitives, and from the decomposition of expected information gain into (i) the probability of experimentation and (ii) the informativeness of outcomes conditional on experimentation (which is increasing in $\pi$ by Proposition 2). ∎
4.11 Equilibrium selection
The monotone methods used in Appendix A deliver extremal cutoffs when multiple fixed points exist. We select the smallest cutoff policy $s^{*}(\cdot)$ (the greatest experimentation equilibrium). This selection is natural for comparative statics and yields the sharpest form of reputational conservatism; it also corresponds to the limit of vanishing payoff perturbations (Appendix A.4).
The next section discusses extensions (endogenous exit, partial observability of effort, committees, continuous-time limits) and connects the results to applications. Full proofs are provided in Appendices AD.
5 Discussion and Extensions
This section develops several extensions that either sharpen the main mechanisms or broaden applicability. [^8] Persistent environments and correlated information are treated in Online Appendix OA.5. In each case we preserve the unified notation from Sections 34 and keep the public state one-dimensional (the reputation $\pi$). Formal proofs and additional details are provided in the appendices referenced below.
5.1 Flagship application: surgery vs. conservative care
We map the model to surgical decision-making. The expert is a surgeon (or surgical team lead); the risky action ($a=1$) is operating; the safe action ($a=0$) is conservative management. The period state $\omega\in\{0,1\}$ collects anatomical/physiological factors that make surgery succeed when favorable ($\omega=1$). The surgeon observes a private clinical signal $s$ (history, exam, imaging) and recommends $a\in\{0,1\}$. The outcome $y\in\{0,1\}$ is publicly observed (e.g., risk-adjusted success/complication). Implementer effort is adherence by the patient and care team (prehab, preparation, post-op protocols, physical therapy, medication adherence), chosen after observing the recommendation and the surgeon’s public reputation $\pi$. The recursive equilibrium (Theorem 5) yields a reputation-dependent cutoff $s^{*}(\pi)$; under Condition 6, the cutoff is (weakly) increasing in $\pi$ (Theorem 7).
Measurement. Reputation $\pi$ can be proxied by risk-adjusted historical performance (e.g., rolling outcome indices), surgeon/hospital quality scores, or volume-based measures; $s$ is a summary of clinical predictors (e.g., an ML risk score from EHR and imaging). Effort proxies include adherence indices (medication possession ratio, refill gaps), therapy attendance, completion of pre-op prep and post-op protocols, and peri-operative checklist compliance. The success variable $y$ is a risk-adjusted binary (e.g., no severe complication within 30/90 days).
Identification. Three sources of quasi-random variation are natural. (i) Assignment shocks: on-call/slot-availability rotations quasi-randomize surgeon–patient matches within service lines; this supports IV/event-study designs for how $\pi$ shifts $a$ and $y$ holding $s$ constant. (ii) Reputation shocks: public report releases or early idiosyncratic successes generate transitory jumps in $\pi$, allowing difference-in-differences around discrete events. (iii) Adherence instruments: congestion at therapy clinics, distance or weather shocks on rehab days, or pharmacy disruptions instrument implementer effort without directly shifting $s$ or $\omega$.
Testable predictions. (P1) Holding clinical risk $s$ fixed, higher $\pi$ reduces the propensity to recommend surgery (negative slope of $a$ on $\pi$): reputational conservatism. (P2) Conditional on recommending surgery, higher $\pi$ raises the hit rate $\mathbb{P}(y=1\mid a=1)$ through higher adherence/effort. (P3) Positive reputation shocks (e.g., early successes) temporarily reduce surgical recommendations and increase the reputational cost of subsequent failures (larger posterior drops after $y=0$). (P4) Monitoring that makes success less diagnostic (e.g., coarse pass/fail scorecards that move with $a$) depresses experimentation relative to benchmarks that preserve outcome diagnosticity; by contrast, pre-outcome transparency about effort (checklists, adherence dashboards) raises experimentation by strengthening the expected payoff of $a=1$.
Design implications. Simple success-contingent bonuses can restore target experimentation by rotating the risky–safe trade-off without requiring complex menus; the calibration maps directly to observed hit-rate shifts. Committee settings (tumor boards) fit the $k$-of-$n$ extension: higher $k$ lowers pivotality and thus raises $s^{*}$, predicting fewer risky recommendations unless countervailed by improved implementation capacity.
5.2 Endogenous exit and demand for advice
In many markets, consultation ceases when reputation is sufficiently low. We model this via an absorbing lower region: fix $\underline{\pi}\in(0,1)$ and assume that if the public reputation after any period satisfies $\pi_{t+1}<\underline{\pi}$, no further agents arrive and the game terminates with continuation value 0.
The Bellman equation (7) then becomes
$V(\pi)=\max_{a\in\{0,1\}}\left\{u(\pi)+\phi\,\mathbf{1}\{a=1\}+\delta\left[P_{S}(1,\pi)\,V\!\big{(}\pi^{+}(\pi)\big{)}+(1-P_{S}(1,\pi))\,\widehat{V}\!\big{(}\pi^{-}(\pi)\big{)}\right]\right\},$
where $\widehat{V}(x)=V(x)$ if $x\geq\underline{\pi}$ and $\widehat{V}(x)=0$ if $x<\underline{\pi}$. The advice cutoffs are defined as in Theorem 5.
Proposition 19.
Suppose Assumptions A1–A3 hold and $\underline{\pi}\in(0,1)$ is an absorbing lower boundary. Then:
  1. 1.
    A (possibly weak) cutoff policy exists for the High type at each $\pi\geq\underline{\pi}$, with the Low type mimicking on-path as before.
  2. 2.
    The cutoff $s^{*}(\pi)$ is (weakly) increasing in $\pi$ on $\underline{\pi},1)$ and has a (weak) kink at the locus where $\pi^{-}(\pi)=\underline{\pi}$.
  3. 3.
    Near the lower boundary, the sign of the change in $s^{*}(\pi)$ relative to the no-exit benchmark depends on the local comparison of the success jump $V(\pi^{+})-V(\underline{\pi})$ and the status-quo value $V(\pi)-V(\underline{\pi})$: if $V(\pi^{+})-V(\underline{\pi})$ is large relative to $V(\pi)-V(\underline{\pi})$, the expert gambles for resurrection (a lower cutoff than without exit); otherwise the policy is more conservative.
The result highlights that exit shapes policy through a discrete downside at failure. In applications where a single failure at low $\pi$ credibly triggers exit, the expert is tempted to take chances if the upside is meaningful, consistent with turnarounds by struggling advisers. Appendix E provides a full proof and examples under quadratic cost and logistic signal families.
5.2.1 Failure-driven exit risk
The next result formalizes that failures are more likely to trigger exit when current reputation is higher.
Corollary 20.
Fix an exit boundary $\underline{\pi}\in(0,1)$ and consider a period in which the expert recommends risk and the outcome is a failure $(a=1,y=0)$. For any signal realization $s$ at which both reputations $\pi^{\prime}<\pi^{\prime\prime}$ would recommend risk, the failure-updated posterior satisfies
$\pi^{-}(\pi^{\prime\prime},s)\ <\ \pi^{-}(\pi^{\prime},s).$
Consequently, for any measurable set of signals on which both reputations recommend risk, the probability of crossing the boundary,
$\mathbb{P}\!\big{\{}\pi^{-}(\pi,s)<\underline{\pi}\ \big{|}\ a=1,\ y=0\big{\}},$
is (weakly) increasing in $\pi$, and strictly increasing whenever the risky region is nondegenerate and Assumption A1 holds on a set of positive measure.
Proof.
Write odds as O(x)=x/(1x)O(x)=x/(1-x). After a risky recommendation and failure,
$O\!\big{(}\pi^{-}(\pi,s)\big{)}\;=\;O\!\big{(}\pi^{\mathrm{rec}}(1;\pi,s)\big{)}\cdot e^{J^{-}(\pi,s)}.$
By Lemma 27, the recommendation-only posterior $\pi^{\mathrm{rec}}(1;\pi,s)$ is increasing in $\pi$, hence $O(\pi^{\mathrm{rec}}(1;\pi,s))$ is increasing in $\pi$. By Lemma 31, $J^{-}(\pi,s)$ is strictly decreasing in the implementer’s effort $e^{*}(1,\pi;s)$, and $e^{*}(1,\pi;s)$ is increasing in $\pi$; thus $J^{-}(\pi,s)$ is (strictly) decreasing in $\pi$, so $e^{J^{-}(\pi,s)}$ is (strictly) decreasing in $\pi$. The product of an increasing term and a decreasing term is (strictly) decreasing here because $e^{J^{-}}$ dominates in the log-odds update: $\partial_{\pi}\log O(\pi^{-})=\partial_{\pi}\log O(\pi^{\mathrm{rec}}(1))+\partial_{\pi}J^{-}<0$ whenever Assumption A1 implies $p_{H}(\cdot)-p_{L}(\cdot)>0$ and $e^{*}$ responds to $\pi$. [^9] Equivalently, $\partial_{\pi}J^{-}(\pi,s)=-\frac{(p_{H}-p_{L})\,\partial_{\pi}e^{*}}{(1-e^{*}p_{H})(1-e^{*}p_{L})}<0$ by Lemma 31 and $\partial_{\pi}e^{*}>0$. Hence $O(\pi^{-}(\pi,s))$ and therefore $\pi^{-}(\pi,s)$ are strictly decreasing in $\pi$ for any fixed $s$ in the overlapping risky region, proving the pointwise claim.
For the probability statement, fix $\pi^{\prime}<\pi^{\prime\prime}$ and any measurable set $S$ of signals on which both reputations recommend $a=1$. The crossing event $\{\pi^{-}(\pi,s)<\underline{\pi}\}$ is an upward-closed set in $s$ (failures become more damaging at higher $s$ since $e^{*}$ increases in $s$), and by the pointwise monotonicity above its indicator is weakly larger at $\pi^{\prime\prime}$ than at $\pi^{\prime}$ on $S$. Integrating against the (absolutely continuous) conditional failure densities on $S$ yields a weakly larger probability at $\pi^{\prime\prime}$, with strict inequality whenever the risky region has positive measure and $J^{-}$ is strictly decreasing. ∎
5.3 Partial observability of effort
The main model assumes that effort $e$ is unobserved; only the binary outcome $y$ is public. In many settings, implementation intensity is partially observable (e.g., audit trails, compliance logs). Let $\tilde{e}\in[0,1]$ be a public signal of effort drawn from a kernel $g(\tilde{e}\mid e)$ satisfying the Blackwell order in an informativeness parameter $\kappa\in[0,\infty)$ (higher $\kappa$ means more precise monitoring; $\kappa=0$ recovers the baseline with no information about $e$).
[…]
Proposition 21.
Under Assumptions A1–A3 and for any $\pi\in(0,1)$, the advice cutoff $s^{*}(\pi;\kappa)$ is (weakly) decreasing in the informativeness $\kappa$ of effort monitoring. In particular, more transparent implementation makes failures less diagnostic of type (holding fixed $a$ and the signal distribution), which raises the expected continuation value of recommending risk and lowers the cutoff.
Proof.
Fix $\pi\in(0,1)$ and consider the risky branch ($a=1$). Let $\tilde{e}_{\kappa}$ denote the public effort monitor with informativeness parameter $\kappa$, ordered in the Blackwell sense (so $\kappa^{\prime}>\kappa$ means $\tilde{e}_{\kappa^{\prime}}$ is a strict refinement/less garbled version of $\tilde{e}_{\kappa}$). The implementer’s effort is chosen before the outcome and is independent of $\theta$ conditional on $(a,\pi,s)$; the signal $\tilde{e}_{\kappa}$ is therefore also independent of $\theta$ given $(a,\pi,s)$.
Let $\pi^{-,\kappa}$ be the failure-updated posterior about $\theta$ under monitor $\tilde{e}_{\kappa}$. As the monitoring $\sigma$-field is refined from $\kappa$ to $\kappa^{\prime}$, Bayes posteriors form a martingale:
$\mathbb{E}\!\big{[}\pi^{-,\kappa^{\prime}}\mid\tilde{e}_{\kappa}\big{]}\;=\;\pi^{-,\kappa}\quad\text{a.s.}$
By Assumption A3, $V$ is increasing and convex, hence by Jensen’s inequality,
$\mathbb{E}\!\big{[}V(\pi^{-,\kappa^{\prime}})\mid\tilde{e}_{\kappa}\big{]}\;\geq\;V\!\big{(}\mathbb{E}[\pi^{-,\kappa^{\prime}}\mid\tilde{e}_{\kappa}]\big{)}\;=\;V(\pi^{-,\kappa})\quad\text{a.s.}$
Taking expectations yields $\mathbb{E}[V(\pi^{-,\kappa^{\prime}})]\geq\mathbb{E}[V(\pi^{-,\kappa})]$. The success branch is unaffected by $\kappa$ because $J^{+}$ does not depend on effort. Therefore the risky–safe continuation difference $\Delta_{H}(s;\pi,\kappa)$ is (weakly) larger at $\kappa^{\prime}$ than at $\kappa$ for every $s$. By strict single crossing of $\Delta_{H}(\cdot;\pi,\kappa)$ in $s$ (Appendix A), the cutoff is (weakly) lower under the more informative monitor: $s^{*}(\pi;\kappa^{\prime})\leq s^{*}(\pi;\kappa)$, with strict inequality when the refinement is strict and the risky region has positive probability. ∎
[…]
5.4 Multiple implementers and committees
Two extensions illustrate how aggregation on the recommendation side and scale on the implementation side shape experimentation.
In a committee with $n$ experts where the risky action is taken if $\sum_{i}a^{i}\geq k$, an individual expert’s outcome-based gain from recommending risk equals the single-expert gain scaled by the pivot probability $\zeta_{k}(\pi)$—the chance her recommendation flips the committee decision. Appendix G shows
$\zeta_{k}(\pi)=\lambda\,\binom{n-1}{k-1}\rho_{1}(\pi)^{\,k-1}\big{(}1-\rho_{1}(\pi)\big{)}^{n-k}+(1-\lambda)\,\binom{n-1}{k-1}\rho_{0}(\pi)^{\,k-1}\big{(}1-\rho_{0}(\pi)\big{)}^{n-k},$
with $\rho_{\omega}(\pi)$ the risky-recommendation frequency of a representative other expert conditional on $\omega$. The cutoff structure carries over: in a symmetric equilibrium the High type uses a threshold $s^{*}(\pi;k)$ (Appendix G). Because only the outcome channel is scaled by $\zeta_{k}(\pi)$—the signaling content of one’s own recommendation is unchanged—a higher threshold $k$ lowers pivotality and thus raises the cutoff under mild conditions ($\rho_{\omega}(\pi)\leq k/n$), reducing experimentation; see Lemma 54 and Proposition 57.
Suppose that following a risky recommendation there are $m\geq 1$ implementers (e.g., a project team or multiple trial sites), each choosing effort $e_{i}\in[0,1]$ at cost $c(e_{i})$, independently across $i$. The project succeeds if at least one implementer succeeds, yielding the reduced-form success probability
$\mathbb{P}(y=1\mid a=1,\omega,\mathbf{e})=1-\prod_{i=1}^{m}\big{(}1-\omega\,e_{i}\big{)}.$
Under symmetry and common knowledge of $(a,\pi)$, each implementer solves $c^{\prime}(e_{i})=\lambda(1,\pi)$, so $e_{i}=e^{*}(1,\pi)$ and
$P_{S}^{(m)}(1,\pi)=1-\Big{(}1-\lambda(1,\pi)\,e^{*}(1,\pi)\Big{)}^{m}.$
Proposition 22.
Fix $\pi$. Then $P_{S}^{(m)}(1,\pi)$ is strictly increasing in $m$, and the High type’s cutoff $s^{*}_{m}(\pi)$ is (weakly) decreasing in $m$. Thus larger implementation scale (more parallel effort) lowers the bar for recommending risk.
The result quantifies a natural complementarity: reputational incentives are stronger when a risky recommendation triggers broader implementation capacity, because the chance of success—and hence the expected reputational gain—is higher. Appendix G provides the details and variants with success thresholds (e.g., $k$-out-of-$m$ success) and correlated efforts. [^11] General monotone aggregation rules, exchangeable/heterogeneous committees, and comparative statics beyond $k$-of-$n$ are in Online Appendix OA.6.
5.5 Policy design: success-contingent bonuses
We show that a simple success bonus implements any target experimentation rate at a given reputation. Let $\beta=(\beta_{1},-\beta_{0})$ denote transfers after $(a,y)=(1,1)$ and $(1,0)$, with limited liability $\beta_{1}\geq 0$, $\beta_{0}\geq 0$. The expert’s risky–safe advantage becomes
$\Delta_{H}^{\beta}(s;\pi)\;=\;\Delta_{H}(s;\pi)\;+\;\alpha\,\beta_{1}\;-\;(1-\alpha)\,\beta_{0},$
so the induced cutoff $s^{*}_{\beta}(\pi)$ solves $\Delta_{H}^{\beta}(s^{*}_{\beta}(\pi);\pi)=0$, and the induced risky rate is $\rho(\pi;\beta)\equiv\sum_{\omega\in\{0,1\}}\mathbb{P}(\omega)\,r_{H}(1\mid\omega,\pi;\beta)$.
Theorem 23.
Fix $\pi\in(0,1)$, assume (A1)–(A4) and strictly positive signal densities at the cutoff. Under limited liability ($\beta_{0}=0$), the map $\beta_{1}\mapsto\rho(\pi;\beta_{1})$ is continuous and strictly increasing with
$\lim_{\beta_{1}\downarrow 0}\rho(\pi;\beta_{1})=\rho(\pi;0),\qquad\lim_{\beta_{1}\uparrow\infty}\rho(\pi;\beta_{1})=1.$
Hence, for any target $\rho^{\star}\in\big{(}\rho(\pi;0),\,1\big{)}$ there exists a unique $\beta_{1}(\rho^{\star})$ such that $\rho(\pi;\beta_{1}(\rho^{\star}))=\rho^{\star}$. With affine transfers $(\beta_{1},-\beta_{0})$, uniqueness extends to any $\rho^{\star}\in(0,1)$ via the net wedge $\alpha\beta_{1}-(1-\alpha)\beta_{0}$.
Proof.
By strict single crossing, $s\mapsto\Delta_{H}(s;\pi)$ is strictly increasing; adding $\alpha\beta_{1}$ shifts $\Delta_{H}$ up and yields a unique $s^{*}_{\beta_{1}}(\pi)$, strictly decreasing and continuous in $\beta_{1}$. Since $\rho(\pi;\cdot)$ is a continuous, strictly decreasing function of the cutoff (with positive densities at $s^{*}$), it is continuous and strictly increasing in $\beta_{1}$. As $\beta_{1}\downarrow 0$, $s^{*}_{\beta_{1}}(\pi)\to s^{*}(\pi)$ and $\rho(\pi;\beta_{1})\to\rho(\pi;0)$; as $\beta_{1}\uparrow\infty$, $s^{*}_{\beta_{1}}(\pi)\to$-$\infty$ and $\rho(\pi;\beta_{1})\to 1$. The affine case follows by replacing $\alpha\beta_{1}$ with $\alpha\beta_{1}-(1-\alpha)\beta_{0}$. Full details are in Online Appendix OA.8.
Proposition 24.
Let $S(\rho)$ denote per-experiment surplus from $a=1$ (concave, differentiable). The implementer chooses $\beta_{1}\geq 0$ to maximize
$U(\beta_{1})\;=\;S(\rho(\pi;\beta_{1}))\;-\;\alpha\,\rho(\pi;\beta_{1})\,\beta_{1},$
with interior FOC
$S^{\prime}(\rho^{\star})\;=\;\alpha\,\beta_{1}^{\star}\;+\;\alpha\,\rho^{\star}\,\frac{1}{\rho^{\prime}(\beta_{1}^{\star})}\,,\qquad\rho^{\star}=\rho(\pi;\beta_{1}^{\star}),\ \ \rho^{\prime}(\beta_{1}^{\star})>0.$
If a budget $B$ imposes $\alpha\,\rho(\pi;\beta_{1})\,\beta_{1}\leq B$, the same FOC holds with a positive multiplier added to the right-hand side at the optimum.
Proof.
Envelope differentiation: $U^{\prime}(\beta_{1})=S^{\prime}(\rho)\rho^{\prime}(\beta_{1})-\alpha[\rho+\beta_{1}\rho^{\prime}(\beta_{1})]$. Setting $U^{\prime}(\beta_{1}^{\star})=0$ and dividing by $\rho^{\prime}(\beta_{1}^{\star})>0$ yields the FOC. Concavity of $S$ and monotonicity of $\rho$ ensure global optimality. See OA–D for regularity and the budgeted case. ∎
One-line Gaussian mapping. With $s\mid(\theta,\omega)\sim\mathcal{N}(\mu_{\omega},\sigma_{\theta}^{2})$ and strictly positive densities at $s^{*}_{\beta_{1}}(\pi)$,
$\rho^{\prime}(\beta_{1})\;=\;\frac{\alpha\Big{[}(1-\alpha)\,f_{H}\!\big{(}s^{*}_{\beta_{1}}(\pi)\mid 0\big{)}+\alpha\,f_{H}\!\big{(}s^{*}_{\beta_{1}}(\pi)\mid 1\big{)}\Big{]}}{\Delta^{\prime}_{H}\!\big{(}s^{*}_{\beta_{1}}(\pi);\pi\big{)}}\;>\;0,$ (9)
and $\frac{ds^{*}_{\beta_{1}}}{d\beta_{1}}=-\alpha/\Delta^{\prime}_{H}(s^{*}_{\beta_{1}}(\pi);\pi)$. These expressions give a plug-in calibration for $\beta_{1}^{\star}$ via Proposition 24. Derivations and numerical recipes are reported in OA–D.
5.6 Continuous-time approximation
Consider a vanishing-period approximation with period length $\Delta>0$ and scale primitives so that (i) signals arrive with intensity $O(\Delta^{-1})$ and the High type’s recommendation region induces a recommendation intensity $\rho(\pi)$; (ii) conditional on $a=1$, successes arrive as a Poisson process with intensity $\Lambda(\pi)\equiv\lambda(1,\pi)\,e^{*}(1,\pi)$; and (iii) belief updates follow log-odds jumps of size $J^{+}(\pi)$ at success and $J^{-}(\pi)$ at failure as in (6). Let $L_{t}\equiv\log\frac{\pi_{t}}{1-\pi_{t}}$.
Proposition 25.
Under regularity conditions on $\rho,\Lambda,J^{\pm}$ (boundedness, Lipschitz continuity), as $\Delta\to 0$ the log-odds process $L_{t}$ under the cutoff policy converges weakly to a jump-diffusion process with local drift and variance
$\begin{split}\mu(\pi)&=\rho(\pi)\,\big{[}\Lambda(\pi)\,J^{+}(\pi)+\big{(}1-\Lambda(\pi)\big{)}\,J^{-}(\pi)\big{]},\\ \sigma^{2}(\pi)&=\rho(\pi)\,\big{[}\Lambda(\pi)\,(J^{+}(\pi))^{2}+\big{(}1-\Lambda(\pi)\big{)}\,(J^{-}(\pi))^{2}\big{]},\end{split}$
where $\pi=\frac{e^{L}}{1+e^{L}}$. In particular, $\mu(\pi)$ is (weakly) increasing in $\pi$ by Proposition 2, capturing the reinforcement of reputation in continuous time.
The limit provides a tractable approximation for inference and policy experiments and connects our discrete-time structure to continuous-time reputation models. Appendix H states the functional central limit theorem used and verifies conditions in a canonical logistic-signal specification.
5.7 Heterogeneous agents and costs
If the arriving agent’s cost function $c_{i}$ varies across periods (e.g., drawn i.i.d. from a known distribution), the best response becomes $e_{i}^{*}(1,\pi)=(c_{i}^{\prime})^{-1}(\lambda(1,\pi))$. The expected effort conditional on $(a=1,\pi)$ is then $E[e_{i}^{*}(1,\pi)]$, which is still increasing in $\lambda(1,\pi)$. All results in Sections 4 carry through with $e^{*}$ replaced by its expectation; in particular, Proposition 2 and Theorem 7 continue to hold.
5.8 Policy and design: contingent fees in dynamic settings
Our companion paper Lukyanov et al. (2025) analyzes a one-shot version with contingent compensation. In the dynamic environment here, a per-period success bonus $B\geq 0$ (paid only when $a=1$ and $y=1$) augments the continuation objective by $\delta\,B$ at success.
Proposition 26.
With a per-period success bonus $B$, the High type’s cutoff $s^{*}(\pi;B)$ is (weakly) decreasing in $B$ for all $\pi$. Moreover, for small $B$ the marginal effect satisfies
$\frac{\partial s^{*}(\pi;B)}{\partial B}\Big{|}_{B=0}<0,$
and the effect is larger in magnitude at higher $\pi$ (where reputational conservatism is strongest).
Thus, modest explicit incentives can partially undo reputational conservatism and restore experimentation at high reputation. [^12] Planner’s welfare and the first-order condition linking $\partial\rho/\partial\beta$ to per-experiment surplus and the budget shadow price, as well as affine $(\beta_{1},-\beta_{0})$ contracts (including corners and implementation sets), are presented in Online Appendix OA.8. See also the minimal-implementation formulas in (31).
These extensions illustrate that the main forces—reputation–effort feedback, state-dependent conservatism, and path-dependent learning—are robust. Transparency in implementation and larger implementation scale reduce conservatism; endogenous exit creates kinks and can induce gambling near the lower boundary; and the continuous-time approximation yields a convenient representation for empirical or quantitative work.
6 Conclusion
This paper develops a dynamic theory of expert advice with reputation–effort feedback. A long-lived expert repeatedly recommends a risky action to short-lived implementers who optimally choose effort in response to the expert’s current reputation. Embedding this complementarity in a recursive formulation delivers a tractable belief-based Markov equilibrium. We prove that the competent expert’s policy is a cutoff in her private signal that depends on public reputation, that the cutoff is (weakly) increasing in reputation (reputational conservatism), and that reputation evolves as a submartingale for competent experts (supermartingale for incompetent ones), hitting boundary regions with probability one. Comparative statics clarify how signal precision, the good-state prior, and patience shift both the cutoff and the law of motion of beliefs.
Two forces drive these results. First, higher reputation elicits greater implementation effort, which raises success probabilities and—crucially—amplifies the informational content of failures without increasing the informational content of successes. Second, with an increasing and convex value of reputation, the downside from a visible failure grows faster than the upside from success at high reputation. These forces jointly tilt established experts toward caution and outsiders toward “gambling for resurrection,” organizing a range of behaviors observed in advisory markets.
The framework is intentionally parsimonious—one-dimensional public state $\pi$, i.i.d. technological state $\omega$, and MLRP signals—yet portable. It applies to financial analysts (buy vs. hold), surgeons (operate vs. conservative care), policy consultants (reform vs. status quo), and R&D leadership (greenlight vs. delay). In each case, reputational status shapes the intensity of implementation, which in turn shapes what outcomes reveal about ability. The equilibrium characterization yields transparent, testable predictions:
  1. 1.
    Holding fundamentals fixed, experts with higher current reputation recommend the risky action less frequently; when they do, success rates are higher because clients work harder.
  2. 2.
    Transitory shocks that raise perceived competence (e.g., an early success) reduce experimentation frequency but increase the informational content of each failure; the opposite holds for negative shocks.
  3. 3.
    Improved signal precision or a higher prior probability of viability increases experimentation across the reputation spectrum; greater patience amplifies conservatism at high reputation.
From a design perspective, reputation partly substitutes for explicit incentives: even without contracts, high reputation alone can sustain implementation effort. But reputation can also over-discipline, stalling experimentation when caution is excessive. Simple instruments can counterbalance this conservatism. We show that monitoring of implementation (transparent effort) and larger implementation scale reduce the cutoff by shrinking the reputational downside of failure. In dynamic environments, modest success-contingent transfers play the same role, rotating the risky–safe tradeoff most where reputational forces are strongest.
The analysis suggests empirical and quantitative agendas. On the empirical side, the model points to reduced-form tests using exogenous shifts in perceived competence (e.g., plausibly random early outcomes) and administrative measures of effort or compliance as proxies for the mechanism. On the structural side, the Gaussian–quadratic benchmark furnishes a workhorse specification for estimating $\pi$-dependent cutoffs, experimentation rates, and belief dynamics from panel advice data with outcomes.
The paper also clarifies the boundary of tractability and indicates natural extensions. Allowing persistent technological states or cross-period spillovers ($\omega_{t}$ Markov) would introduce an exploration–exploitation tradeoff atop reputational concerns, connecting our setup to dynamic bandits with dual learning about technology and ability. Richer communication (multi-message or continuous recommendations) and heterogeneous implementers (costs or observability) can be accommodated within the recursive approach at the cost of additional state variables. Finally, endogenous market frictions—entry, exit, and competition among experts—would convert the reputation–effort feedback into an industry-level selection mechanism.
We view the main contribution as conceptual and methodological: a clean recursive formulation that isolates the interaction between reputation and implementer effort, yields sharp comparative statics and dynamics, and remains close to data. We hope the framework can serve as a baseline for studying advice, disclosure, and implementation in organizations and markets where outcomes are jointly produced by ideas and execution.
Appendix A Proof of Theorem 5
This appendix proves that the competent expert ($\theta=H$) has a cutoff policy: for each public reputation $\pi\in(0,1)$ there is a threshold $s^{*}(\pi)\in S$ such that she recommends the risky action iff her private signal $s\geq s^{*}(\pi)$. The proof proceeds in three steps: (i) single crossing of the marginal value in $s$; (ii) existence and uniqueness of the zero; (iii) measurability and equilibrium consistency.
A.1 Preliminaries
Fix $\pi\in(0,1)$. Let $\Delta_{H}(s;\pi)$ denote the High type’s risky-minus-safe continuation gain when others’ strategies are fixed at the symmetric profile described in Section 3. With our timing and observables,
$\Delta_{H}(s;\pi)=\phi+\delta\left\{\underbrace{\mathbb{E}\big{[}V(\pi^{\prime})\mid a=1,\pi,s\big{]}}_{\text{recommendation \& outcome}}-\underbrace{\mathbb{E}\big{[}V(\pi^{\prime})\mid a=0,\pi,s\big{]}}_{\text{recommendation only}}\right\},$
where $V$ is the continuation value and $\pi^{\prime}$ the next-period reputation. Define the recommendation-only posterior $\pi^{\mathrm{rec}}(a;\pi,s)$ (from observing $a$) and, conditional on $a=1$, the outcome posteriors $\pi^{+}(\pi,s)$ and $\pi^{-}(\pi,s)$ after success/failure. Let $P_{S}(s;\pi)$ be the success probability under $a=1$. Then
$\begin{split}\Delta_{H}(s;\pi)=\phi&+\delta\Big{(}P_{S}(s;\pi)\,V(\pi^{+}(\pi,s))\\ &+\big{(}1-P_{S}(s;\pi)\big{)}\,V(\pi^{-}(\pi,s))-V\!\left(\pi^{\mathrm{rec}}(0;\pi,s)\right)\Big{)}.\end{split}$ (10)
A.2 Likelihood ratios and posterior monotonicity
Let $f_{\theta}(\cdot\mid\omega)$ denote the density of $s$ under type $\theta\in\{H,L\}$ and state $\omega\in\{0,1\}$. Assumption A1 (MLRP and Blackwell domination) implies the type log-likelihood ratio
$\Lambda_{\theta}(s)\equiv\log\frac{f_{H}(s\mid\omega)}{f_{L}(s\mid\omega)}$
is (strictly) increasing in $s$ for each $\omega$; likewise, the state likelihood ratio
$\Lambda_{\omega}(s)\equiv\log\frac{f_{\theta}(s\mid\omega=1)}{f_{\theta}(s\mid\omega=0)}$
is (strictly) increasing in $s$ for each $\theta$. Bayes’ rule therefore gives:
Lemma 27.
For every $\pi\in(0,1)$: (i) the recommendation-only posterior $\pi^{\mathrm{rec}}(1;\pi,s)$ is (strictly) increasing in $s$ and $\pi^{\mathrm{rec}}(0;\pi,s)$ is (strictly) decreasing in $s$; (ii) the posterior success belief $\lambda(1,\pi;s)$ after $a=1$ is (strictly) increasing in $s$; hence the agent’s best-response effort $e^{*}(1,\pi;s)$ solving $c^{\prime}(e)=\lambda(1,\pi;s)$ is (strictly) increasing in $s$.
Proof.
Part (i) follows since $\mathbb{P}(a=1\mid s,\theta)$ is (weakly) higher under $H$ than $L$ for larger $s$ in any best response (MLRP ensures the monotone likelihood ratio for the event $a=1$ in $s$), so the posterior odds about $\theta$ after observing $a$ are increasing in $s$. Part (ii) is the standard MLRP implication for the $\omega$-posterior given $s$. Strict monotonicity obtains whenever the signal has strictly increasing likelihood ratios. The mapping $e^{*}$ is increasing because $c^{\prime}$ is strictly increasing by Assumption A2. ∎
The log-likelihood jumps in $\pi$ after outcomes when $a=1$ are
$J^{+}(\pi,s)$ $=\log\frac{p_{H}(\pi,s)}{p_{L}(\pi,s)},$ (11)
$J^{-}(\pi,s)$ $=\log\frac{1-e^{*}(1,\pi;s)\,p_{H}(\pi,s)}{1-e^{*}(1,\pi;s)\,p_{L}(\pi,s)},$ (12)
where $p_{\theta}(\pi,s)\equiv\mathbb{P}(\omega=1\mid a=1,\theta,\pi,s)$. Note $J^{+}$ is independent of effort, while $J^{-}$ is strictly decreasing in $e^{*}$ and hence in $s$.
A.3 Single crossing and existence
Lemma 28.
Under Assumptions A1–A3, $s\mapsto\Delta_{H}(s;\pi)$ is continuous and strictly increasing on $S$.
Proof.
Continuity follows from continuity of the primitives and the dominated convergence theorem. To prove monotonicity, examine (8). First, $s\mapsto\pi^{\mathrm{rec}}(0;\pi,s)$ is decreasing by Lemma 27(i) while $V$ is increasing (A3), so $s\mapsto-V(\pi^{\mathrm{rec}}(0;\pi,s))$ is increasing. Second, $s\mapsto P_{S}(s;\pi)$ is increasing because both $\lambda(1,\pi;s)$ and $e^{*}(1,\pi;s)$ increase with $s$ (Lemma 27(ii) and A2). Third, conditional on $a=1$, $s\mapsto\pi^{+}(\pi,s)$ increases with $s$ and $s\mapsto\pi^{-}(\pi,s)$ decreases with $s$ because $J^{+}$ is independent of $e^{*}$ while $J^{-}$ becomes more negative as $e^{*}$ rises; since $V$ is increasing and convex (A3), the map
$s\ \longmapsto\ P_{S}(s;\pi)\,V(\pi^{+}(\pi,s))+\big{(}1-P_{S}(s;\pi)\big{)}\,V(\pi^{-}(\pi,s))$
is increasing. Adding the three components yields the claim; strictness follows from strict MLRP and strict convexity of $c$ (hence strict increase of $e^{*}$). ∎
Lemma 29.
For each $\pi$, there exist $s_{-},s_{+}\in S$ with $s_{-}<s_{+}$ such that $\Delta_{H}(s_{-};\pi)<0$ and $\Delta_{H}(s_{+};\pi)>0$.
Proof.
As $s\to\inf S$, $\lambda(1,\pi;s)$ and $e^{*}$ become small, so $P_{S}(s;\pi)$ is arbitrarily small, while the recommendation-only posterior after $a=0$ approaches its upper support; hence the Jensen gain from switching to risk is dominated by the safe continuation and $\Delta_{H}<0$ for $s$ low enough. As $s\to\sup S$, both $\lambda(1,\pi;s)$ and $e^{*}$ approach their upper supports so that $P_{S}(s;\pi)$ and $V(\pi^{+})$ make the experimentation option dominant; thus $\Delta_{H}>0$ for $s$ high enough. The argument uses that $V$ is increasing and bounded on $(0,1)$. ∎
Proof of Theorem 5.
By Lemma 28, $s\mapsto\Delta_{H}(s;\pi)$ is strictly increasing and continuous; by Lemma 29 it changes sign. Hence there exists a unique $s^{*}(\pi)\in S$ with $\Delta_{H}(s^{*}(\pi);\pi)=0$, and the High type’s best reply is the cutoff policy $a=1$ iff $s\geq s^{*}(\pi)$. Measurability and existence of a symmetric MPBE follow by standard fixed-point arguments (Appendix A.4). ∎
A.4 Existence of MPBE
Under Assumptions A1–A4 and the cutoff structure of Theorem 5, the set of symmetric strategy profiles with cutoff policies is a nonempty, compact, convex lattice under the pointwise order; the induced best–response correspondence is nonempty, convex–valued, and monotone (single–crossing of the marginal value in $s$ implies monotone best responses). By Tarski’s fixed–point (or, equivalently, by Kakutani applied to a compact convex subset of $L^{\infty}$), a symmetric MPBE exists. The continuity of posteriors and $V$ ensures the usual measurable–selection requirements.
Appendix B Proof of Theorem 7
We prove that the cutoff $s^{*}(\pi)$ is (weakly) increasing in the public reputation $\pi$. The proof uses Topkis’ monotone comparative statics: it suffices to show that $\Delta_{H}(s;\pi)$ has decreasing differences in $(s,\pi)$ (i.e., the cross-partial $\partial_{s\pi}\Delta_{H}\leq 0$ in the sense of monotone differences).
B.1 Decomposing the marginal value
Write $\Delta_{H}(s;\pi)=\phi+\delta\{\Gamma_{1}(s;\pi)-\Gamma_{0}(s;\pi)\}$ with
$\begin{split}\Gamma_{0}(s;\pi)&\equiv V\big{(}\pi^{\mathrm{rec}}(0;\pi,s)\big{)},\\ \Gamma_{1}(s;\pi)&\equiv P_{S}(s;\pi)\,V\big{(}\pi^{+}(\pi,s)\big{)}+\big{(}1-P_{S}(s;\pi)\big{)}V\big{(}\pi^{-}(\pi,s)\big{)}.\end{split}$
The dependence on $\pi$ enters through (i) the recommendation posteriors $\pi^{\mathrm{rec}}$ (because the prior odds enter Bayes’ rule), and (ii) $P_{S}$ and the outcome posteriors $\pi^{\pm}$ via the agent’s effort best reply $e^{*}(1,\pi;s)$ and the jumps $J^{\pm}$.
B.2 Decreasing differences
Lemma 30.
The map $(s,\pi)\mapsto\Gamma_{0}(s;\pi)$ has decreasing differences.
Proof.
By Bayes’ rule, $\pi^{\mathrm{rec}}(0;\pi,s)=\sigma(\operatorname{logit}(\pi)+J^{\mathrm{rec}}_{0}(s))$, where $\sigma(x)=\frac{1}{1+e^{-x}}$ and $J^{\mathrm{rec}}_{0}(s)$ is the (type) log-likelihood ratio of observing $a=0$. Lemma 27(i) implies $J^{\mathrm{rec}}_{0}(s)$ is decreasing in $s$. The map $(x,z)\mapsto\sigma(x+z)$ has decreasing differences because $\sigma$ is increasing and concave. Composing with the increasing $V$ preserves the property (Topkis). ∎
B.3 LLR asymmetry and effort
We record a basic property used repeatedly in the decreasing–differences proof.
Lemma 31.
When $a=1$, the success jump $J^{+}(\pi,s)$ is independent of effort, while the failure jump $J^{-}(\pi,s)$ is strictly decreasing in the implementer’s effort $e^{*}(1,\pi;s)$. Consequently, as $s$ or $\pi$ rise (both increase $e^{*}$), failures become (weakly) more damaging to reputation, whereas the informativeness of success is unchanged.
Proof.
By (11), $J^{+}(\pi,s)=\log\!\big{(}p_{H}(\pi,s)/p_{L}(\pi,s)\big{)}$ does not involve $e^{*}$. By the same display,
$J^{-}(\pi,s)=\log\!\frac{1-e^{*}(1,\pi;s)\,p_{H}(\pi,s)}{1-e^{*}(1,\pi;s)\,p_{L}(\pi,s)},$
whose derivative with respect to $e^{*}(1,\pi;s)$ equals $\big{(}p_{H}(\pi,s)-p_{L}(\pi,s)\big{)}\big{/}\big{(}(1-e^{*}p_{H})(1-e^{*}p_{L})\big{)}<0$ by Assumption A1. Since $e^{*}(1,\pi;s)$ increases in $s$ and $\pi$ (Lemma 27), the comparative statics follow. ∎
Lemma 32.
The map $(s,\pi)\mapsto\Gamma_{1}(s;\pi)$ has decreasing differences.
Proof.
Three effects matter.
(i) Effort amplification. By Lemma 27(ii), $s\mapsto\lambda(1,\pi;s)$ is increasing. For fixed $s$, the agent’s success belief $\lambda(1,\pi;s)$ is also increasing in $\pi$ because the recommendation posterior $\pi^{\mathrm{rec}}(1;\pi,s)$ is increasing in $\pi$ and raises the weight on the more informative type $H$; therefore the best response $e^{*}(1,\pi;s)$ is increasing in both $s$ and $\pi$. As a result, $P_{S}(s;\pi)=\lambda(1,\pi;s)\,e^{*}(1,\pi;s)$ is supermodular in $(s,\pi)$.
(ii) LLR asymmetry. The success jump $J^{+}$ is independent of effort (hence of $\pi$), while the failure jump $J^{-}$ is strictly decreasing in $e^{*}(1,\pi;s)$ (see (12)); thus $s$ and $\pi$ jointly make failures more damaging but do not change the informativeness of success. This creates decreasing differences in the pair $(s,\pi)$ for the outcome-updated value $V(\pi^{\pm}(\pi,s))$: higher $\pi$ reduces the marginal gain from raising $s$ because the downside risk of a failure (whose probability rises with $s$) becomes more severe.
(iii) Jensen’s inequality. Since $V$ is increasing and convex, the map $q\mapsto q\,V(\pi^{+})+(1-q)\,V(\pi^{-})$ has decreasing differences in $(q,\pi)$ when $\pi^{\pm}$ move apart in $\pi$ mainly through the more negative $J^{-}$. Combining (i)–(iii) delivers decreasing differences of $\Gamma_{1}$. ∎
Proof of Theorem 7.
By Lemmas 30 and 32, $\Delta_{H}(s;\pi)=\phi+\delta\{\Gamma_{1}(s;\pi)-\Gamma_{0}(s;\pi)\}$ has decreasing differences in $(s,\pi)$. Lemma 28 gives strict single crossing in $s$. Topkis’ monotone comparative statics theorem then implies the smallest (and here unique) solution $s^{*}(\pi)$ to $\Delta_{H}(s;\pi)=0$ is (weakly) increasing in $\pi$. ∎
Appendix C Proof of Theorem 12
We prove that public reputation $\{\pi_{t}\}$ is a submartingale when the true type is $H$ and a supermartingale when the true type is $L$, and that it reaches trust regions with probability one. The key step is an exact expression for the conditional drift of $\pi_{t}$ in terms of likelihood ratios.
C.1 One-step update and conditional drift
Let $L_{t+1}$ be the one-step likelihood ratio between histories under $H$ and $L$ (recommendation and, if applicable, outcome) given $\mathcal{F}_{t}$:
$L_{t+1}\;\equiv\;\frac{\mathbb{P}(\text{obs}_{t+1}\mid\theta=H,\mathcal{F}_{t})}{\mathbb{P}(\text{obs}_{t+1}\mid\theta=L,\mathcal{F}_{t})}\;\in(0,\infty).$
Bayes’ rule yields the posterior odds recursion
$\frac{\pi_{t+1}}{1-\pi_{t+1}}=\frac{\pi_{t}}{1-\pi_{t}}\,L_{t+1},\qquad\pi_{t+1}=g(\pi_{t},L_{t+1}),\quad g(\pi,L)\equiv\frac{\pi L}{\pi L+1-\pi}.$
Hence
$\pi_{t+1}-\pi_{t}\;=\;\pi_{t}(1-\pi_{t})\,\frac{L_{t+1}-1}{\pi_{t}L_{t+1}+1-\pi_{t}}.$ (13)
Lemma 33.
For every $t$, under the probability measure with $\theta=H$,
$\mathbb{E}\!\left[\pi_{t+1}-\pi_{t}\mid\mathcal{F}_{t},\theta=H\right]\ \geq\ 0,$
with equality iff $L_{t+1}=1$ almost surely (i.e., the period-$t$ observation is uninformative about type). Symmetrically, under $\theta=L$ the conditional drift is $\leq 0$, with equality iff $L_{t+1}=1$ almost surely.
Proof.
Let $a\equiv\pi_{t}\in(0,1)$ and $h(x)\equiv(x-1)/(ax+1-a)$. Then (13) gives $\pi_{t+1}-\pi_{t}=a(1-a)\,h(L_{t+1})$. Under $\theta=H$, the law of $L_{t+1}$ has density $d\mathbb{P}_{H}=L_{t+1}\,d\mathbb{P}_{L}$ with respect to the law under $L$. Thus
$\mathbb{E}_{H}[h(L_{t+1})\mid\mathcal{F}_{t}]\;=\;\mathbb{E}_{L}\!\left[\frac{L_{t+1}(L_{t+1}-1)}{aL_{t+1}+1-a}\,\Big{|}\,\mathcal{F}_{t}\right].$
The integrand is nonnegative because $x\mapsto\frac{x(x-1)}{ax+1-a}$ is nondecreasing and vanishes at $x=1$. Hence the conditional expectation is $\geq 0$, with equality only if $L_{t+1}=1$ a.s. under $L$ (and therefore under $H$). The statement under $\theta=L$ follows by symmetry (replace $L_{t+1}$ with $1/L_{t+1}$). ∎
Proof of Theorem 12.
The sub-/supermartingale property follows by Lemma 33 and boundedness of $\pi_{t}\in[0,1]$. For boundary hitting, note that if information arrives infinitely often with positive probability (i.e., $\mathbb{P}(L_{t+1}\neq 1\ \text{i.o.})>0$), then the sum of nonnegative conditional drifts $\sum_{t}\mathbb{E}_{H}[\pi_{t+1}-\pi_{t}\mid\mathcal{F}_{t}]$ diverges on that event unless $\pi_{t}$ enters an arbitrarily small neighborhood of $1$; Doob’s submartingale convergence theorem then implies $\pi_{t}\to 1$ almost surely under $H$. A symmetric argument yields $\pi_{t}\to 0$ under $L$. If informative periods cease after some (random) time, the process is eventually constant and trivially hits a boundary region. In either case, the process reaches trust neighborhoods with probability one. ∎
C.2 LLR asymmetry and effort (for reference)
When the risky action is recommended, $L_{t+1}$ decomposes as $L_{t+1}=L^{\mathrm{rec}}_{t+1}\cdot L^{\mathrm{out}}_{t+1}$, the product of the recommendation LLR and the outcome LLR. The latter has jumps $J^{+}$ and $J^{-}$ as in (11)–(12); crucially, $J^{+}$ is independent of effort while $J^{-}$ is strictly decreasing in the agent’s effort $e^{*}(1,\pi_{t};s_{t})$. This is the microfoundation for the greater reputational downside of failures at high reputation used in Theorem 7.
Appendix D Comparative Statics Proofs
This appendix provides proofs for Theorem 17. We study how the equilibrium cutoff $s^{*}(\pi)$ moves with primitives.
D.1 Signal precision (Blackwell order)
Let the High type’s signal distribution be indexed by a precision parameter $\kappa$, with $\kappa^{\prime}\succcurlyeq\kappa$ in the Blackwell order (i.e., $f_{H}(\cdot\mid\omega;\kappa^{\prime})$ is more informative about $\omega$ than $f_{H}(\cdot\mid\omega;\kappa)$ for each $\omega$). The Low type’s distribution is held fixed.
Lemma 34.
Fix $(s,\pi)$. Under $\kappa^{\prime}\succcurlyeq\kappa$, the posterior success probability $\mathbb{P}(\omega=1\mid s,a=1,\pi;\kappa^{\prime})$ is (weakly) higher than $\mathbb{P}(\omega=1\mid s,a=1,\pi;\kappa)$ for all $s$ in the MLR order. Consequently, $P_{S}(s;\pi)$ is (weakly) higher under $\kappa^{\prime}$ than under $\kappa$.
Proof.
Blackwell dominance implies that for any prior over $\omega$ and any likelihood ratio test based on $s$, the posterior places (weakly) more mass on $\omega=1$ under the more informative experiment. As $e^{*}(1,\pi)$ is increasing in the success belief and the signal of effort is unchanged, the reduced-form success probability increases. ∎
Proposition 35.
If $\kappa^{\prime}\succcurlyeq\kappa$ (High type more informative), then for all $\pi\in(0,1)$ the High type’s cutoff satisfies $s^{*}(\pi;\kappa^{\prime})\leq s^{*}(\pi;\kappa)$.
Proof.
From (10), $\Delta_{H}(s;\pi)$ is strictly increasing in $P_{S}(s;\pi)$. By Lemma 34, $P_{S}(s;\pi)$ shifts up under $\kappa^{\prime}$, so the zero set $\{s:\Delta_{H}(s;\pi)=0\}$ weakly shifts left, and the monotone selection implies the cutoff falls. ∎
D.2 Good-state prior $\lambda$
Proposition 36.
For all $\pi\in(0,1)$, the cutoff $s^{*}(\pi)$ is (weakly) decreasing in the prior probability $\lambda$ of a favorable state.
Proof.
A larger $\lambda$ raises $\mathbb{P}(\omega=1\mid s,a=1,\pi)$ for all $s$ and hence raises $P_{S}(s;\pi)$ pointwise. The argument of Proposition 35 applies verbatim. ∎
D.3 Patience $\delta$
We establish the comparative static under a sufficient curvature condition.
Proposition 37.
Suppose $V$ is increasing and convex on $(0,1)$ (Assumption A3 plus convexity). Then for all $\pi\in(0,1)$, the cutoff $s^{*}(\pi)$ is (weakly) increasing in the discount factor $\delta$.
Proof.
From (10), $\Delta_{H}(s;\pi)=\phi+\delta\big{(}P_{S}(s;\pi)\,V(\pi^{+})+(1-P_{S}(s;\pi))\,V(\pi^{-})-V(\pi)\big{)}$. The derivative with respect to $\delta$ is the bracketed term, which is negative at high $\pi$ under convexity because $V(\cdot)$ exhibits diminishing returns in $\pi$ and, by Appendix B.3, failures become more damaging (reduce $\pi$ more) as $e^{*}$ rises with $\pi$. Globally, Appendix B shows that $\Delta_{H}$ has decreasing differences in $(1,\pi)$; multiplying a function with decreasing differences by a larger scalar $\delta$ preserves the monotone comparative static: the zero-crossing in $s$ shifts (weakly) upward. A direct Topkis argument (decreasing differences in $(a,\delta)$) delivers $s^{*}(\pi)$ nondecreasing in $\delta$. ∎
D.4 A worked benchmark
This subsection fully develops the Gaussian–quadratic benchmark and derives the comparative statics in signal precision, the good-state prior, and patience. Throughout we fix $c(e)=\tfrac{1}{2}e^{2}$, so that
$e^{*}(1,\pi;s)\;=\;\lambda(1,\pi;s),$ (14)
where the dependence on the High type’s cutoff $s^{*}(\pi)$ enters through the advice likelihoods below.
D.4.1 Signal distributions, advice likelihoods, and posteriors
Conditional on the state $\omega\in\{0,1\}$, the expert’s private signal $s$ is Gaussian:
$s\mid(\theta=\!H,\omega)\sim\mathcal{N}(\mu_{\omega},\sigma_{H}^{2}),\qquad s\mid(\theta=\!L,\omega)\sim\mathcal{N}(\mu_{\omega},\sigma_{L}^{2}),$
with $\mu_{1}-\mu_{0}>0$ and $\sigma_{L}^{2}>\sigma_{H}^{2}$ (so $H$ is strictly more informative than $L$ in the Blackwell sense). Fix a belief $\pi\in(0,1)$. In equilibrium, the High type uses a cutoff $s^{*}(\pi)$ (Theorem 5), recommending risk if and only if $s\geq s^{*}(\pi)$. Define for $\omega\in\{0,1\}$:
$\begin{split}A(s)\equiv r_{H}(1\mid\omega\!=\!1,\pi)&=1-\Phi\!\Big{(}\frac{s-\mu_{1}}{\sigma_{H}}\Big{)},\\ B(s)\equiv r_{H}(1\mid\omega\!=\!0,\pi)&=1-\Phi\!\Big{(}\frac{s-\mu_{0}}{\sigma_{H}}\Big{)},\end{split}$ (15)
where $\Phi$ is the standard normal cdf. Note that $A(s)>B(s)$ and both are strictly decreasing in $s$.
For the Low type, to keep on-path beliefs Bayes-consistent while keeping the algebra transparent, we adopt a simple frequency-matching signal-independent mixed strategy: upon any $(s,\pi)$, the Low type recommends $a=1$ with probability
$p(\pi)\in(0,1),\qquad r_{L}(1\mid\omega,\pi)=p(\pi)\ \ \text{ for }\ \omega\in\{0,1\},$
so the event $\{a=1\}$ carries no information about $\omega$ under LL. [^13] Any alternative measurable construction that maintains $r_{L}(1\mid 1,\pi)\geq r_{L}(1\mid 0,\pi)$ works as well; the comparative statics below are unchanged.
Given a cutoff $s=s^{*}(\pi)$, the interim posterior that the state is good after observing $a=1$ is
$\lambda(1,\pi;s)\;=\;\mathbb{P}(\omega=1\mid a=1,\pi)\;=\;\frac{\pi\,[\lambda A(s)]+\,(1-\pi)\,[\lambda p(\pi)]}{\pi\,[\lambda A(s)+(1-\lambda)B(s)]+\,(1-\pi)\,p(\pi)}.$ (16)
Using (14), the success probability conditional on $a=1$ is
$P_{S}(1,\pi;s)\;=\;\lambda(1,\pi;s)^{2}.$ (17)
For Bayesian updating about type, define, as in Appendix B,
$\begin{split}p_{H}(\pi;s)&\equiv\mathbb{P}(\omega=1\mid\theta\!=\!H,a=1,\pi)=\frac{\lambda A(s)}{\lambda A(s)+(1-\lambda)B(s)},\\ p_{L}(\pi)&\equiv\mathbb{P}(\omega=1\mid\theta\!=\!L,a=1,\pi)=\lambda.\end{split}$
The log-likelihood jumps (Appendix C) are
$\begin{split}J^{+}(\pi;s)&=\log\frac{p_{H}(\pi;s)}{p_{L}(\pi)}=\log\frac{A(s)}{\lambda A(s)+(1-\lambda)B(s)}\ >0,\\ J^{-}(\pi;s)&=\log\frac{1-e^{*}p_{H}(\pi;s)}{1-e^{*}p_{L}(\pi)}\ <0,\end{split}$ (18)
with $e^{*}=\lambda(1,\pi;s)$.
D.4.2 Monotone effects of Gaussian precision
Let $\sigma^{2}_{H}$ be the High type’s variance. Differentiating (15) at a fixed threshold $s$ gives
$\frac{\partial A}{\partial\sigma_{H}}=\varphi\!\Big{(}\frac{s-\mu_{1}}{\sigma_{H}}\Big{)}\frac{s-\mu_{1}}{\sigma_{H}^{2}}\ <\ 0,\qquad\frac{\partial B}{\partial\sigma_{H}}=\varphi\!\Big{(}\frac{s-\mu_{0}}{\sigma_{H}}\Big{)}\frac{s-\mu_{0}}{\sigma_{H}^{2}}\ >\ 0,$ (19)
whenever $\mu_{0}<s<\mu_{1}$, where $\varphi$ is the standard normal pdf. Thus, a precision loss (higher $\sigma_{H}$) reduces the true-positive rate $A$ and increases the false-positive rate $B$ for the event $\{s\geq s^{*}(\pi)\}$.
Lemma 38.
Holding $(s,\pi)$ fixed, $\partial_{\sigma_{H}}\lambda(1,\pi;s)<0$ and $\partial_{\sigma_{H}}P_{S}(1,\pi;s)<0$.
Proof.
Write (16) as $\lambda(1,\pi;s)=N/D$ with
$N=\lambda\big{[}\pi A+(1-\pi)p(\pi)\big{]},\qquad D=\pi\big{[}\lambda A+(1-\lambda)B\big{]}+(1-\pi)p(\pi).$
Using (19), $\partial_{\sigma_{H}}N=\lambda\pi\,\partial_{\sigma_{H}}A<0$ while $\partial_{\sigma_{H}}D=\pi\big{[}\lambda\,\partial_{\sigma_{H}}A+(1-\lambda)\,\partial_{\sigma_{H}}B\big{]}$ has the sign of $(1-\lambda)\partial_{\sigma_{H}}B+\lambda\partial_{\sigma_{H}}A>0$ because $|\partial_{\sigma_{H}}A|$ and $\partial_{\sigma_{H}}B$ are both positive and $(1-\lambda)\geq\lambda$ cannot hold for all $\lambda$, but, crucially, for any $\lambda\in(0,1)$ and $\mu_{0}<s<\mu_{1}$, the Gaussian tails satisfy
$(1-\lambda)\,\varphi\!\Big{(}\frac{s-\mu_{0}}{\sigma_{H}}\Big{)}\frac{s-\mu_{0}}{\sigma_{H}^{2}}\;+\;\lambda\,\varphi\!\Big{(}\frac{s-\mu_{1}}{\sigma_{H}}\Big{)}\frac{\mu_{1}-s}{\sigma_{H}^{2}}\;>\;0,$
since each term is strictly positive. Hence $\partial_{\sigma_{H}}D>0$. By the quotient rule,
$\partial_{\sigma_{H}}\lambda(1,\pi;s)=\frac{D\,\partial_{\sigma_{H}}N-N\,\partial_{\sigma_{H}}D}{D^{2}}<0.$
Equation (17) then yields $\partial_{\sigma_{H}}P_{S}=2\lambda(1,\pi;s)\,\partial_{\sigma_{H}}\lambda(1,\pi;s)<0$. ∎
Lemma 39.
Holding $(s,\pi)$ fixed, $\partial_{\sigma_{H}}J^{+}(\pi;s)<0$ and $\partial_{\sigma_{H}}J^{-}(\pi;s)>0$.
Proof.
Differentiate (18). For $J^{+}$,
$\partial_{\sigma_{H}}J^{+}=\frac{\partial_{\sigma_{H}}A}{A}-\frac{\lambda\,\partial_{\sigma_{H}}A+(1-\lambda)\,\partial_{\sigma_{H}}B}{\lambda A+(1-\lambda)B}<0$
by (19) and the same positivity argument as in Lemma 38. For $J^{-}$, holding $p_{H}$ fixed for the moment,
$\frac{\partial J^{-}}{\partial e^{*}}=-\Big{(}\frac{p_{H}}{1-e^{*}p_{H}}-\frac{p_{L}}{1-e^{*}p_{L}}\Big{)}<0\quad\text{since }p_{H}>p_{L}=\lambda,$
so the fall in $e^{*}$ from Lemma 38 raises $J^{-}$. In addition, $\partial_{\sigma_{H}}p_{H}<0$ by the same calculus as for $J^{+}$, which further increases $J^{-}$ (makes it less negative). Thus $\partial_{\sigma_{H}}J^{-}>0$. ∎
Proposition 40.
For any $\pi\in(0,1)$ with interior experimentation, $\dfrac{\partial s^{*}(\pi)}{\partial\sigma_{H}^{2}}>0$.
Proof.
Let $\Delta_{H}(s;\pi)$ denote the risky-minus-safe continuation difference (Appendix A). The cutoff satisfies $\Delta_{H}(s^{*}(\pi);\pi)=0$ and $\partial_{s}\Delta_{H}>0$ (single crossing). By the implicit function theorem,
$\frac{\partial s^{*}(\pi)}{\partial\sigma_{H}^{2}}\;=\;-\frac{\partial_{\sigma_{H}^{2}}\Delta_{H}(s;\pi)}{\partial_{s}\Delta_{H}(s;\pi)}\Bigg{|}_{s=s^{*}(\pi)}.$
Using (17) and the update formulae for $\pi^{\pm}$ via $J^{\pm}$, one obtains
$\partial_{\sigma_{H}^{2}}\Delta_{H}=\delta\Big{[}(\partial_{\sigma_{H}^{2}}P_{S})\,\big{(}V(\pi^{+})-V(\pi^{-})\big{)}+P_{S}\,V^{\prime}(\pi^{+})\,\partial_{\sigma_{H}^{2}}\pi^{+}+(1-P_{S})\,V^{\prime}(\pi^{-})\,\partial_{\sigma_{H}^{2}}\pi^{-}\Big{]}.$
By Lemma 38, $\partial_{\sigma_{H}^{2}}P_{S}<0$. By Lemma 39, $\partial_{\sigma_{H}^{2}}\pi^{+}<0$ and $\partial_{\sigma_{H}^{2}}\pi^{-}>0$ (success becomes less, and failure more, like the prior), and with $V^{\prime}(\cdot)\geq 0$ and $V(\pi^{+})>V(\pi^{-})$, the bracket is strictly negative. Since $\partial_{s}\Delta_{H}>0$, the sign is $\partial s^{*}/\partial\sigma_{H}^{2}>0$. ∎
D.4.3 Effect of the good-state prior $\lambda$
Lemma 41.
Holding $(s,\pi)$ fixed, $\partial_{\lambda}\lambda(1,\pi;s)>0$ and $\partial_{\lambda}P_{S}(1,\pi;s)>0$.
Proof.
From (16), $N=\lambda[\pi A+(1-\pi)p]$ and $D=\pi[\lambda A+(1-\lambda)B]+(1-\pi)p$. Then
$\begin{split}\partial_{\lambda}\lambda(1,\pi;s)&=\frac{D[\pi A+(1-\pi)p]-N\cdot\pi(A-B)}{D^{2}}\\ &=\frac{(1-\pi)p\cdot\pi(A-B)+\pi A\cdot[(1-\pi)p+\pi(1-\lambda)B]}{D^{2}}>0,\end{split}$
since $A>B$ and all terms are nonnegative with at least one strictly positive. Hence $\partial_{\lambda}P_{S}=2\lambda(1,\pi;s)\,\partial_{\lambda}\lambda(1,\pi;s)>0$. ∎
Lemma 42.
Holding $(s,\pi)$ fixed, $\partial_{\lambda}J^{+}(\pi;s)<0$ and $\partial_{\lambda}J^{-}(\pi;s)>0$.
Proof.
From (18), $J^{+}=\log\frac{A}{\lambda A+(1-\lambda)B}$, so $\partial_{\lambda}J^{+}=-\frac{A-B}{\lambda A+(1-\lambda)B}<0$. For $J^{-}$, $p_{L}=\lambda$ increases with $\lambda$ while $p_{H}$ also (weakly) increases with $\lambda$ but remains $p_{L}$; holding $e^{*}$ fixed,
$\partial_{\lambda}J^{-}=\frac{\partial}{\partial\lambda}\log\frac{1-e^{*}p_{H}}{1-e^{*}\lambda}=-\frac{e^{*}\,\partial_{\lambda}p_{H}}{1-e^{*}p_{H}}+\frac{e^{*}}{1-e^{*}\lambda}\ >\ 0,$
because $\partial_{\lambda}p_{H}\in(0,1)$ and the second term dominates when $p_{H}>p_{L}=\lambda$. The rise in $e^{*}$ from Lemma 41 further increases $J^{-}$ (less negative). ∎
Proposition 43.
For any $\pi\in(0,1)$ with interior experimentation, $\dfrac{\partial s^{*}(\pi)}{\partial\lambda}<0$.
Proof.
As in the proof of Proposition 40, apply the implicit function theorem. Using Lemma 41, $\partial_{\lambda}P_{S}>0$ raises the first term in $\partial_{\lambda}\Delta_{H}$; Lemma 42 moves $\pi^{\pm}$ toward $\pi$ (success less informative, failure less damaging), which increases the expected continuation from taking risk because $V$ is increasing; thus $\partial_{\lambda}\Delta_{H}>0$ and $\partial s^{*}/\partial\lambda<0$. ∎
D.4.4 Effect of patience $\delta$
In the benchmark, $\delta$ scales the expected continuation difference from taking risk:
$\frac{\partial\Delta_{H}(s;\pi)}{\partial\delta}\;=\;P_{S}(1,\pi;s)\,V(\pi^{+}(\pi;s))+\big{(}1-P_{S}(1,\pi;s)\big{)}\,V(\pi^{-}(\pi;s))-V(\pi).$ (20)
The sign of (20) is governed by how the lottery over $\{\pi^{-},\pi^{+}\}$ compares to the sure thing $\pi$. With $V$ convex and $\pi$ the prior mean of the posterior (law of total expectation for posteriors), Jensen’s inequality yields
$P_{S}\,V(\pi^{+})+(1-P_{S})\,V(\pi^{-})\;\geq\;V\big{(}P_{S}\,\pi^{+}+(1-P_{S})\,\pi^{-}\big{)}\;=\;V(\pi),$ (21)
with strict inequality whenever $\pi^{+}\neq\pi^{-}$. However, as $\pi$ rises the downside ($\pi^{-}\!<\pi$) becomes more likely to be triggered by higher effort (Appendix B), and the informational content of success does not rise with effort—so the convexity comparison increasingly favors the status quo at high $\pi$ once we internally account for how $\pi^{\pm}$ depend on $\pi$ (Appendix B.2).
The benchmark therefore delivers the same qualitative comparative static as Theorem 17(iii), and we can make it explicit by bounding $\pi^{\pm}$.
Lemma 44.
Fix $\pi$. There exist constants $0<\underline{\eta}(\pi)<\overline{\eta}(\pi)<\infty$ (computable from $A,B$) such that
$\begin{split}\pi^{+}(\pi;s)&\in\Big{[}\pi+\underline{\eta}(\pi)\,(1-\pi)\,,\ \ \pi+\overline{\eta}(\pi)\,(1-\pi)\Big{]},\\ \pi^{-}(\pi;s)&\in\Big{[}\pi-\overline{\eta}(\pi)\,\pi\,,\ \ \pi-\underline{\eta}(\pi)\,\pi\Big{]},\end{split}$
with $\underline{\eta}(\pi)$ decreasing and $\overline{\eta}(\pi)$ increasing in $\sigma_{H}$, and $\underline{\eta}(\pi)$ decreasing in $\lambda$.
Proof.
Immediate from the odds-ratio updates with $J^{\pm}$ in (18) and the monotonicity in Lemmas 3942; the bounds follow by mapping log-odds intervals to probability space. ∎
Proposition 45.
Suppose $V$ is increasing and convex. Then for any $\pi$ at which experimentation is interior,
$\frac{\partial s^{*}(\pi)}{\partial\delta}\ \geq\ 0,$
with strict inequality whenever $\pi$ is sufficiently high. [^14] At very low $\pi$, the reputational upside can dominate under strong convexity of $V$, in which case $s^{*}(\pi)$ is locally flat in $\delta$; the inequality is weak globally and strict for $\pi$ above a model-dependent threshold.
Proof.
By (20), the sign of $\partial_{\delta}\Delta_{H}$ equals the sign of the Jensen gap in (21). Using Lemma 44, as $\pi$ rises the spread $(\pi^{+}-\pi^{-})$ around $\pi$ contracts asymmetrically toward the downside (failure becomes more anti-$H$ with higher effort), so the convexity gain from taking risk shrinks and eventually becomes negative. Therefore $\partial_{\delta}\Delta_{H}\leq 0$ for all $\pi$ (weakly) and <0<0 for $\pi$ high enough, implying $\partial s^{*}/\partial\delta\geq 0$ and >0>0 at high $\pi$ by the implicit function theorem. ∎
In the quadratic–Gaussian benchmark, for any interior $\pi$,
$\frac{\partial s^{*}(\pi)}{\partial\sigma_{H}^{2}}>0,\qquad\frac{\partial s^{*}(\pi)}{\partial\lambda}<0,\qquad\frac{\partial s^{*}(\pi)}{\partial\delta}\geq 0\ \ (\text{strict for large }\pi).$
These conclusions align with the general results in Theorem 17 and make the channels explicit: precision shapes both the frequency and informativeness of successes/failures through $(A,B)$; the prior $\lambda$ raises success odds and softens failure; and patience scales a convex-reputation tradeoff that is tilted toward conservatism at high reputation.
Appendix E Exit Options and Equilibrium Selection
This appendix proves Proposition 19. The expert may irrevocably exit at the start of a period and secure a reservation value $U_{0}$ that is independent of reputation. Let $\widehat{V}(\pi)$ denote the expert’s value function when exit is available.
E.1 Value recursion with an exit option
At belief $\pi$, after the High type observes $s$ and chooses $a\in\{0,1\}$, continuation value equals
$\widehat{V}(\pi)=\max\Big{\{}U_{0},\quad\underbrace{\max_{a\in A_{H}(s,\pi)}\big{\{}\underbrace{\phi\,a}_{\text{stage payoff}}+\delta\,\mathbb{E}\big{[}\widehat{V}(\pi^{\prime})\mid a,\pi\big{]}\big{\}}}_{\equiv W(\pi)}\Big{\}}.$
By the same arguments as in Section 3, $W(\pi)$ is well defined and bounded, and $\widehat{V}$ is the smallest bounded solution to the Bellman equation above.
Lemma 46.
$\widehat{V}(\pi)\geq V(\pi)$ for all $\pi$. If $U_{0}\leq\min_{\pi}V(\pi)$, then $\widehat{V}\equiv V$.
Proof.
Trivial from $\max\{U_{0},W(\pi)\}\geq W(\pi)$ and the definition of $V$ as the fixed point without exit. ∎
E.2 Exit thresholds and reputation reinforcement
Define the (upper hemicontinuous) experimentation correspondence $\Gamma(\pi)\subseteq\{0,1\}$ for the High type as in Section 4. Let $\Pi_{\mathrm{exp}}\equiv\{\pi:\,1\in\Gamma(\pi)\}$ and $\Pi_{\mathrm{no}}\equiv\{\pi:\,0\in\Gamma(\pi)\}$.
Proposition 47.
There exists $\pi^{\mathrm{exit}}\in[0,1)$ such that: (i) if $\pi\leq\pi^{\mathrm{exit}}$, then $\widehat{V}(\pi)=U_{0}$ and the High type exits with probability one; (ii) if $\pi>\pi^{\mathrm{exit}}$ and $U_{0}<\sup_{\pi}V(\pi)$, then exit does not occur on the equilibrium path.
Proof.
The map $\pi\mapsto W(\pi)$ is increasing and upper semicontinuous (Appendix B). The set $\{\pi:W(\pi)\leq U_{0}\}$ is therefore a (possibly empty) closed interval $0,\pi^{\mathrm{exit}}]$. On this set, $\widehat{V}(\pi)=U_{0}$; off it, $\widehat{V}(\pi)=W(\pi)>U_{0}$ and exit is strictly dominated. Monotonicity of $\Gamma$ implies that for $\pi>\pi^{\mathrm{exit}}$ experimentation occurs with (weakly) higher probability, yielding strictly more convex reputation dynamics and reinforcing the no-exit region. ∎
Corollary 48.
If multiple MPBE exist at some $\pi$, the equilibrium with (weakly) higher experimentation probability yields (weakly) lower $\pi^{\mathrm{exit}}$ and strictly dominates in welfare for any $U_{0}<\sup_{\pi}V(\pi)$.
Proof.
Follows from the sub/supermartingale arguments in Appendix C: higher experimentation raises expected log-likelihood drift under $H$ and reduces the measure of belief states at which $W(\pi)\leq U_{0}$. ∎
Appendix F Monitoring and Disclosure
We enrich the baseline with a public monitoring signal $m_{t}\in\{0,1\}$ realized after the recommendation $a_{t}\in\{0,1\}$ and observed by all players before the outcome. Conditional on the action and the expert’s type,
$\mathbb{P}(m_{t}=1\mid a_{t}=1,\theta)=q_{\theta},\qquad\mathbb{P}(m_{t}=1\mid a_{t}=0,\theta)=\bar{q}_{\theta},$ (22)
with $q_{H}>q_{L}$ and $\bar{q}_{H}\geq\bar{q}_{L}$. The realized pair $(y_{t},m_{t})$ updates the public reputation by Bayes’ rule; let $\pi_{t+1}=\Phi(\pi_{t};y_{t},m_{t})$ denote the posterior. We maintain Assumptions A1–A4.
F.1 Preliminaries and log-likelihood ratios
Fix a period with public reputation $\pi\in(0,1)$. When $a=1$, the outcome $y\in\{0,1\}$ is generated as in the baseline: conditional on $\theta$ and on the induced effort $e^{*}(1,\pi)$, the success probability equals $e^{*}(1,\pi)\,p_{\theta}(\pi)$, where $p_{H}(\pi)>p_{L}(\pi)$ by Assumption A1. The monitoring signal $m$ is conditionally independent of $y$ given $(a,\theta)$ and has distribution (22).
It is convenient to work with the log-likelihood ratio (LLR) increments for the type posterior (Appendix C). Define
$\begin{split}J^{+}(\pi)&\equiv\log\frac{\mathbb{P}(y=1\mid\theta\!=\!H,a\!=\!1,\pi)}{\mathbb{P}(y=1\mid\theta\!=\!L,a\!=\!1,\pi)}=\log\frac{p_{H}(\pi)}{p_{L}(\pi)}\;>\;0,\\ J^{-}(\pi)&\equiv\log\frac{\mathbb{P}(y=0\mid H,a\!=\!1,\pi)}{\mathbb{P}(y=0\mid L,a\!=\!1,\pi)}=\log\frac{1-e^{*}p_{H}(\pi)}{1-e^{*}p_{L}(\pi)}\;<\;0,\end{split}$
with $e^{*}=e^{*}(1,\pi)$, as in (11)–(12). The monitoring LLRs are
$J_{m}^{(1)}\equiv\log\frac{q_{H}}{q_{L}}\ \text{ if }m=1,\quad J_{m}^{(1)}=\log\frac{1-q_{H}}{1-q_{L}}\ \text{ if }m=0\qquad(a=1),$ (23)
$J_{m}^{(0)}\equiv\log\frac{\bar{q}_{H}}{\bar{q}_{L}}\ \text{ if }m=1,\quad J_{m}^{(0)}=\log\frac{1-\bar{q}_{H}}{1-\bar{q}_{L}}\ \text{ if }m=0\qquad(a=0).$ (24)
Since $m$ and $y$ are conditionally independent given $(a,\theta)$, the composite LLR for $(y,m)$ is additive. If $a=1$, the one-period log-odds innovation equals $J^{+}$ or $J^{-}$ (from $y$) plus $J_{m}^{(1)}$ (from $m$); if $a=0$, it equals $J_{m}^{(0)}$.
Lemma 49.
For each fixed action $a\in\{0,1\}$, the composite observation $(y,m)$ (with $y$ vacuous when $a=0$) is ordered by the likelihood ratio $\ell_{a}(y,m)\equiv\tfrac{\mathbb{P}(y,m\mid\theta=H,a,\pi)}{\mathbb{P}(y,m\mid\theta=L,a,\pi)}$. Then $\pi\mapsto\Phi(\pi;y,m)$ is strictly increasing in $\ell_{a}(y,m)$, and the High type’s risky-minus-safe continuation difference $\Delta_{H}(s;\pi)$ preserves single crossing in the private signal $s$.
Proof.
By Bayes’ rule, the posterior odds equal prior odds times the LLR, hence $\Phi(\pi;y,m)$ is strictly increasing in $\ell_{a}(y,m)$. Under Assumption A1, the High type’s reduced-form success probability $P_{S}(s;\pi)$ is strictly increasing in $s$, while the incremental information from $m$ does not depend on $s$. Therefore $\Delta_{H}(s;\pi)$ adds an $s$-independent term to the baseline expression (10), preserving strict single crossing in $s$ and the cutoff representation of Theorem 5. ∎
F.2 Blackwell order and the option value of experimentation
Let $\mathcal{L}_{a}(\cdot;\pi)$ denote the law of the next-period posterior $\pi^{\prime}$ conditional on current reputation $\pi$ and action $a$. (Formally, $\pi^{\prime}$ is a measurable function of $(y,m)$ and current $\pi$.) The expert’s continuation under $a$ equals $\mathbb{V}_{a}(\pi)\;\equiv\;\mathbb{E}\big{[}V(\pi^{\prime})\mid a,\pi\big{]}$, $V$ is increasing and convex by Assumption A3 and Proposition 2. The option value of experimentation is $\mathbb{V}_{1}(\pi)-\mathbb{V}_{0}(\pi)$, the bracketed term in (10) up to the factor $\delta$.
We parameterize the informativeness of the monitor by two scalars $(\kappa,\bar{\kappa})\in[0,\infty)^{2}$, writing
[…]
with $q_{\theta}(0)=\bar{q}_{\theta}(0)=\tfrac{1}{2}$. Larger $(\kappa,\bar{\kappa})$ thus represent Blackwell improvements of $m$ under $a=1$ and under $a=0$, respectively.
Lemma 50.
Fix $\pi$. If $\bar{\kappa}^{\prime}\!>\bar{\kappa}$, then $\mathcal{L}_{0}(\cdot;\pi,\bar{\kappa}^{\prime})$ Blackwell-dominates $\mathcal{L}_{0}(\cdot;\pi,\bar{\kappa})$, hence $\mathbb{V}_{0}(\pi;\bar{\kappa}^{\prime})>\mathbb{V}_{0}(\pi;\bar{\kappa})$ whenever $V$ is strictly convex. Similarly, if $\kappa^{\prime}\!>\kappa$, then $\mathcal{L}_{1}(\cdot;\pi,\kappa^{\prime})$ Blackwell-dominates $\mathcal{L}_{1}(\cdot;\pi,\kappa)$ and $\mathbb{V}_{1}(\pi;\kappa^{\prime})>\mathbb{V}_{1}(\pi;\kappa)$.
Proof.
For $a=0$, the posterior is updated only with $m$, whose binary experiment becomes strictly more informative as $\bar{\kappa}$ increases (the LLR support expands by (24)). Blackwell’s theorem implies a mean-preserving spread in the posterior, which strictly raises $\mathbb{E}[V(\pi^{\prime})]$ under strict convexity. The case $a=1$ is identical with (23). ∎
Proposition 51.
Fix $\pi$ and let $s^{*}(\pi;\kappa,\bar{\kappa})$ be the High-type cutoff under monitor parameters $(\kappa,\bar{\kappa})$. Then:
  1. (i)
    For all $\pi$ at which experimentation is interior,
    $\frac{\partial s^{*}(\pi;\kappa,\bar{\kappa})}{\partial\bar{\kappa}}\ \geq\ 0,$
    with strict inequality whenever $V$ is strictly convex and the safe monitor is not degenerate.
  2. (ii)
    For all interior $\pi$,
    $\frac{\partial s^{*}(\pi;\kappa,\bar{\kappa})}{\partial\kappa}\ \leq\ 0,$
    with strict inequality under strict convexity and a nondegenerate risky monitor.
Proof.
By Lemma 49, $s^{*}(\pi;\kappa,\bar{\kappa})$ solves the first-order condition $\Delta_{H}(s;\pi,\kappa,\bar{\kappa})=0$ with $\partial_{s}\Delta_{H}>0$. By the implicit function theorem,
$\frac{\partial s^{*}(\pi;\kappa,\bar{\kappa})}{\partial\bar{\kappa}}=-\,\frac{\partial_{\bar{\kappa}}\Delta_{H}(s;\pi,\kappa,\bar{\kappa})}{\partial_{s}\Delta_{H}(s;\pi,\kappa,\bar{\kappa})}\Bigg{|}_{s=s^{*}}.$
Using (10) and that $m$ arrives under both actions,
$\partial_{\bar{\kappa}}\Delta_{H}\;=\;\delta\left(\partial_{\bar{\kappa}}\mathbb{V}_{1}(\pi;\kappa,\bar{\kappa})\;-\;\partial_{\bar{\kappa}}\mathbb{V}_{0}(\pi;\kappa,\bar{\kappa})\right).$
The risky posterior law depends on $m$ through $q_{\theta}(\kappa)$ but not through $\bar{q}_{\theta}(\bar{\kappa})$, so $\partial_{\bar{\kappa}}\mathbb{V}_{1}=0$, whereas $\partial_{\bar{\kappa}}\mathbb{V}_{0}>0$ by Lemma 50. Hence $\partial_{\bar{\kappa}}\Delta_{H}<0$, which yields $\partial_{\bar{\kappa}}s^{*}(\pi)\geq 0$, with strict inequality under strict convexity. The proof of (ii) is analogous: $\partial_{\kappa}\mathbb{V}_{0}=0$ and $\partial_{\kappa}\mathbb{V}_{1}>0$, hence $\partial_{\kappa}\Delta_{H}>0$ and $\partial_{\kappa}s^{*}(\pi)\leq 0$, with strictness as stated. ∎
Corollary 52.
For any $\kappa\geq 0$, there exists $\bar{\kappa}_{0}>0$ such that for all $\bar{\kappa}\geq\bar{\kappa}_{0}$ the High-type cutoff satisfies $s^{*}(\pi;\kappa,\bar{\kappa})\geq s^{*}(\pi;\kappa,0)$ for all $\pi$, with strict inequality on a set of $\pi$ of positive measure. In particular, along any equilibrium path that visits this set with positive probability, the ex-ante experimentation frequency strictly falls when the safe monitor is made sufficiently informative.
Proof.
Immediate from Proposition 51(i) and the strictness conclusion when $V$ is strictly convex and the monitor is nondegenerate. ∎
F.3 Failure asymmetry with monitoring
Monitoring alters not only the option value but also the composition of diagnostic content. Under $a=1$, the composite LLR after success is $J^{+}(\pi)+J_{m}^{(1)}(m)$ while after failure it is $J^{-}(\pi)+J_{m}^{(1)}(m)$. Since $J^{+}$ is independent of effort but $J^{-}$ becomes more negative as effort rises (Lemma 31), the gap in informativeness between success and failure narrows when a part of the signal $(m)$ is available regardless of outcome or action; when the monitor under $a=0$ is precise (large $\bar{\kappa}$), the incremental informational advantage of $a=1$ comes primarily from $y$, which is exactly the channel penalized by higher effort at high reputation. The next proposition formalizes the resulting tilt toward conservatism.
Proposition 53.
Suppose $V$ is increasing and convex and let $\bar{\kappa}>0$. Then the reputational-conservatism result (Theorem 7) continues to hold with monitoring, and the slope $\frac{d}{d\pi}s^{*}(\pi;\kappa,\bar{\kappa})$ is (weakly) larger for $\bar{\kappa}$ sufficiently large. In particular, for $\pi$ in a neighborhood of $1$, the High-type cutoff under monitoring exceeds the no-monitoring cutoff.
Proof.
By Lemma 49, single crossing and Topkis monotonicity in $(a,\pi)$ are preserved, so $s^{*}(\pi)$ is (weakly) increasing in $\pi$ as in Theorem 7. To compare slopes, differentiate $\Delta_{H}(s;\pi,\kappa,\bar{\kappa})$ with respect to $\pi$ as in Appendix B.2. The effort channel (which makes failures more damaging at higher $\pi$) is unchanged, while the baseline continuation $V(\pi)$ is unaffected. The only new term is the safe-action information $\mathbb{V}_{0}(\pi;\bar{\kappa})$, which is independent of $s$ but increases with $\bar{\kappa}$. Since this term reduces the convexity gain from switching to $a=1$, the decreasing-differences argument strengthens, yielding a (weakly) larger slope of $s^{*}$ in $\pi$ when $\bar{\kappa}$ is large. The local dominance near $\pi=1$ follows because at high reputation the downside (failure) dominates the upside (success), and subtracting a convexity gain from the safe action magnifies the precautionary motive. ∎
Proposition 51 shows a sharp crowding-out: information that arrives under the safe action raises the value of not experimenting and thus increases the risky-signal cutoff. Conversely, information that arrives only under the risky action encourages experimentation. Proposition 53 documents that, beyond levels, monitoring steepens the reputation–conservatism relation at the top, where the downside of a visible failure is largest due to higher implementation effort. These forces rationalize empirical situations in which transparency of process or effort reduces appetite for risky recommendations even when monitors are accurate.
Appendix G Committees and Advice Aggregation
This appendix strengthens the committee extension by deriving pivot probabilities explicitly, showing that the High type’s problem preserves single crossing, and establishing how the cutoff depends on the aggregation threshold $k$.
G.1 Environment
There are $n\geq 2$ experts indexed by $i=1,\dots,n$. Each expert $i$ has a fixed type $\theta^{i}\in\{H,L\}$, i.i.d. across $i$ with prior $\pi\in(0,1)$ that a given expert is $H$. In period $t$, expert $i$ privately observes a signal $s^{i}\in S$ about the period state $\omega_{t}\in\{0,1\}$ (i.i.d. over $t$ with $\mathbb{P}(\omega_{t}=1)=\lambda\in(0,1)$). Signal families satisfy Assumption A1 (MLRP; $H$ Blackwell-dominates $L$). Experts simultaneously issue recommendations $a^{i}\in\{0,1\}$ (risky vs. safe). The implementer adopts the risky action iff at least $k$ of the $n$ recommendations are risky:
$a\;=\;\mathbf{1}\!\left\{\sum_{i=1}^{n}a^{i}\geq k\right\},\qquad k\in\{1,\dots,n\}.$
If $a=1$, the implementer chooses effort $e\in[0,1]$ as in the baseline; success occurs with probability $e\,\mathbf{1}\{\omega=1\}$; if $a=0$, $y=0$ with certainty. All recommendations and the outcome are publicly observed, and individual reputations about ability are updated by Bayes’ rule. The period payoffs and the expert’s continuation value are as in the baseline (Section 3); we focus on a symmetric MPBE in which all $H$-types use a cutoff $s^{*}(\pi)$ and all $L$-types use some (possibly different) measurable strategy. [^15] Nothing below requires the $L$-type to use a cutoff; we only use that for each $\omega$ and $\pi$ the induced risky-recommendation probability of a generic other expert is well defined.
Let $r_{\theta}(1\mid\omega,\pi)$ denote the probability that a type-$\theta$ expert issues $a^{j}=1$ conditional on $\omega$ when the public reputation is $\pi$. From the viewpoint of a given expert $i$, the risky-recommendation probability of any other expert $j\neq i$ conditional on $\omega$ is
$\rho_{\omega}(\pi)\;\equiv\;\pi\,r_{H}(1\mid\omega,\pi)\;+\;(1-\pi)\,r_{L}(1\mid\omega,\pi),\qquad\omega\in\{0,1\}.$ (25)
Conditional on $(\omega,\pi)$, the number of other risky recommendations
$S_{-i}\;\equiv\;\sum_{j\neq i}a^{j}$
is binomial: $S_{-i}\sim\mathrm{Bin}(n-1,\rho_{\omega}(\pi))$ and is independent of $i$’s private signal $s^{i}$.
G.2 Pivot probability
Lemma 54.
Fix $\pi$ and $k\in\{1,\dots,n\}$. The probability that expert $i$ is pivotal—i.e., that the final decision switches from $a=0$ to $a=1$ when she switches her recommendation from 0 to 11—equals
$\begin{split}\zeta_{k}(\pi)\;&\equiv\;\sum_{\omega\in\{0,1\}}\mathbb{P}(\omega)\;\mathbb{P}\big{(}S_{-i}=k-1\mid\omega,\pi\big{)}\\ \;&=\;\lambda\,b_{n-1,k-1}\!\left(\rho_{1}(\pi)\right)\;+\;(1-\lambda)\,b_{n-1,k-1}\!\left(\rho_{0}(\pi)\right),\end{split}$ (26)
where $b_{N,j}(p)\equiv\binom{N}{j}p^{j}(1-p)^{N-j}$ is the binomial pmf.
Proof.
For fixed $\omega$, $a=1$ under $a^{i}=1$ iff $S_{-i}\geq k-1$, and under $a^{i}=0$ iff $S_{-i}\geq k$. The only realizations that change the committee’s decision are those with $S_{-i}=k-1$. The independence across experts conditional on $\omega$ and $\pi$ yields $\mathbb{P}(S_{-i}=k-1\mid\omega,\pi)=b_{n-1,k-1}(\rho_{\omega}(\pi))$. Averaging over $\omega$ gives (26). ∎
G.3 Decomposing the High type’s marginal value
Let $\Delta_{H}^{\mathrm{comm}}(s;\pi,k)$ be the High type’s risky-minus-safe continuation gain at signal $s$, holding others to the symmetric profile. There are two channels:
  • (i)
    an outcome channel that operates only when the expert is pivotal (her recommendation changes the committee action and hence whether an outcome and reputational jump are realized);
  • (ii)
    a signaling channel coming from the informational content of her recommendation $a^{i}$ about $\theta^{i}$ even when the committee’s action is the same with $a^{i}=0$ or $1$.
Formally, write
$\Delta_{H}^{\mathrm{comm}}(s;\pi,k)\;=\;\underbrace{\zeta_{k}(\pi)\,\Delta_{H}(s;\pi)}_{\text{outcome channel}}\;+\;\underbrace{\Xi_{H}(s;\pi)}_{\text{signaling channel}}.$ (27)
Here $\Delta_{H}(s;\pi)$ is the baseline (single-expert) risky-minus-safe continuation difference from (10)—it values the move from $a=0$ to $a=1$ holding fixed that the recommendation determines the action—and $\Xi_{H}(s;\pi)$ collects the pure signaling effect of $a^{i}$ on the posterior about $\theta^{i}$ when the committee action is unchanged. Importantly, $\Xi_{H}$ does not depend on $k$.
Lemma 55.
For any fixed $(\pi,k)$, the map $s\mapsto\Delta_{H}^{\mathrm{comm}}(s;\pi,k)$ has the strict single-crossing property. Consequently, the High type’s best reply is a cutoff in $s$, and in a symmetric MPBE the High type uses a cutoff $s^{*}(\pi;k)$.
Proof.
By Lemma 3 in Appendix A, $s\mapsto\Delta_{H}(s;\pi)$ has strict single crossing. The pivot probability $\zeta_{k}(\pi)$ is constant in $s$. The signaling term $s\mapsto\Xi_{H}(s;\pi)$ inherits single crossing from Assumption A1 because, under MLRP, the LLR of the event $\{a^{i}=1\}$ relative to $\{a^{i}=0\}$ is increasing in $s$; hence the reputational benefit of choosing $a^{i}=1$ rather than $a^{i}=0$ is (weakly) increasing in ss. [^16] Formally, let $L_{s}\equiv\log\frac{f_{H}(s\mid\omega)}{f_{L}(s\mid\omega)}$ be the signal LLR (increasing in $s$ by MLRP). The log posterior odds after observing the recommendation $a^{i}\in\{0,1\}$ but fixing the committee action is $L_{0}+\log\frac{\mathbb{P}(a^{i}\mid H)}{\mathbb{P}(a^{i}\mid L)}$, and the difference between $a^{i}=1$ and $a^{i}=0$ is increasing in $s$. A sum of (strictly) single-crossing functions is (strictly) single crossing; thus (27) preserves the property and the cutoff characterization follows as in Theorem 5. ∎
G.4 Comparative statics in the threshold $k$
We now show how $k$ shifts the cutoff. All $k$-dependence in (27) is through the pivot probability $\zeta_{k}(\pi)$.
Lemma 56.
Fix $\pi$ and let $p_{\omega}\equiv\rho_{\omega}(\pi)\in(0,1)$ for $\omega\in\{0,1\}$. Then for any $k\in\{1,\dots,n-1\}$,
$\frac{\zeta_{k+1}(\pi)}{\zeta_{k}(\pi)}\;=\;\lambda\,R_{n-1,k}(p_{1})\,\frac{b_{n-1,k-1}(p_{1})}{\zeta_{k}(\pi)}\;+\;(1-\lambda)\,R_{n-1,k}(p_{0})\,\frac{b_{n-1,k-1}(p_{0})}{\zeta_{k}(\pi)},$
where $R_{N,k}(p)\equiv\dfrac{b_{N,k}(p)}{b_{N,k-1}(p)}=\dfrac{N-k+1}{k}\cdot\dfrac{p}{1-p}$. In particular, if $p_{0}\leq\tfrac{k}{n}$ and $p_{1}\leq\tfrac{k}{n}$, then $\zeta_{k+1}(\pi)\leq\zeta_{k}(\pi)$, with strict inequality unless $p_{0}=p_{1}=\tfrac{k}{n}$.
Proof.
The identity is the binomial pmf ratio. If $p\leq\tfrac{k}{n}$, then for N=n1N=n-1 one has $R_{N,k}(p)\leq 1$ (because $\tfrac{N-k+1}{k}\leq\tfrac{n-k}{k}$ and $\tfrac{p}{1-p}\leq\tfrac{k/n}{1-k/n}=\tfrac{k}{n-k}$). The convex combination with nonnegative weights therefore does not exceed one; strictness is immediate unless both terms equal one. ∎
Proposition 57.
Fix $\pi$ and suppose $p_{\omega}=\rho_{\omega}(\pi)\leq\tfrac{k}{n}$ for $\omega\in\{0,1\}$ (in particular, for majority rules $k\geq\lceil n/2\rceil$ this holds whenever $p_{0},p_{1}\leq 1/2$). Then the High type’s cutoff $s^{*}(\pi;k)$ is (weakly) increasing in $k$, with strict inequality whenever $\Delta_{H}(s^{*}(\pi;k);\pi)>0$ (i.e., the outcome channel has strictly positive value at the threshold).
Proof.
By Lemma 55, $s^{*}(\pi;k)$ solves $\Delta_{H}^{\mathrm{comm}}(s;\pi,k)=0$ with $\partial_{s}\Delta_{H}^{\mathrm{comm}}>0$. Differencing in $k$,
$\Delta_{H}^{\mathrm{comm}}(s;\pi,k+1)-\Delta_{H}^{\mathrm{comm}}(s;\pi,k)=\big{(}\zeta_{k+1}(\pi)-\zeta_{k}(\pi)\big{)}\,\Delta_{H}(s;\pi),$
since the signaling term $\Xi_{H}(s;\pi)$ does not depend on $k$. By Lemma 56, $\zeta_{k+1}(\pi)\leq\zeta_{k}(\pi)$ under the stated condition. At $s=s^{*}(\pi;k)$, we have $\Delta_{H}^{\mathrm{comm}}(\cdot;\pi,k)=0$, so
$\Delta_{H}^{\mathrm{comm}}(s^{*}(\pi;k);\pi,k+1)=\big{(}\zeta_{k+1}-\zeta_{k}\big{)}\,\Delta_{H}\!\big{(}s^{*}(\pi;k);\pi\big{)}\;\leq\;0,$
with strict inequality if $\Delta_{H}(s^{*}(\pi;k);\pi)>0$ and $\zeta_{k+1}<\zeta_{k}$. Because $\partial_{s}\Delta_{H}^{\mathrm{comm}}>0$, the implicit-function theorem implies $s^{*}(\pi;k+1)\geq s^{*}(\pi;k)$, strictly when the inequality above is strict. ∎
Proposition 57 shows that when other members are not “too risk-prone” (precisely, their risky probability $p_{\omega}$ does not exceed $k/n$), raising the threshold $k$ shrinks the pivot probability $\zeta_{k}$ and thereby raises the cutoff. Intuitively, stricter aggregation makes a single member’s recommendation less likely to change the committee’s action, so the outcome-based option value of recommending risk is attenuated, while the pure signaling value of the recommendation is unaffected; to restore indifference at the margin, the High type requires a stronger private signal. The result applies in particular to majority rules when $p_{0},p_{1}\leq 1/2$, a case that is empirically plausible at higher reputations where experimentation is less frequent (Theorem 7).
G.5 LLR calculus for own-recommendation signaling
For completeness, we record the likelihood-ratio contribution of the expert’s own recommendation, which underlies the signaling term $\Xi_{H}(s;\pi)$ in (27). Let $R_{\theta}(\pi)\equiv\mathbb{P}(a^{i}=1\mid\theta,\pi)$ be the (unconditional-in-$\omega$) risky frequency induced by the strategy at $\pi$. The LLR contribution of switching from $a^{i}=0$ to $a^{i}=1$ while holding the committee action fixed is
$J^{\mathrm{rec}}(\pi)\;\equiv\;\log\frac{R_{H}(\pi)}{R_{L}(\pi)}-\log\frac{1-R_{H}(\pi)}{1-R_{L}(\pi)}\;=\;\log\frac{R_{H}(\pi)\,[1-R_{L}(\pi)]}{R_{L}(\pi)\,[1-R_{H}(\pi)]},$ (28)
which is strictly positive whenever $R_{H}(\pi)>R_{L}(\pi)$. Under MLRP and cutoff strategies with (weakly) lower threshold for $H$ than for $L$, $R_{H}(\pi)\geq R_{L}(\pi)$ holds, and the mapping $s\mapsto\Xi_{H}(s;\pi)$ is increasing because the event $\{a^{i}=1\}$ is more likely for $H$ at higher signals. [^17] In the special case in which the $L$-type mixes to exactly match the $H$-type’s risky frequency ($R_{H}(\pi)=R_{L}(\pi)$), $J^{\mathrm{rec}}(\pi)=0$ and the signaling term vanishes; then $\Delta_{H}^{\mathrm{comm}}(s;\pi,k)=\zeta_{k}(\pi)\,\Delta_{H}(s;\pi)$ and the cutoff is independent of $k$. Proposition 57 therefore isolates the comparative static in $k$ to environments where recommendations carry some informational content about type.
Combining Lemma 54, Lemma 55, Proposition 57, and the recommendation-LLR (28) yields a complete characterization of the High type’s policy in committees: it remains a reputation-dependent cutoff, and it is (weakly) more conservative in larger-$k$ committees whenever others’ risk propensity is below $k/n$, with strict conservatism whenever recommending risk has positive option value at the margin.
Appendix H Continuous-Time Approximation
Let periods have length $\Delta>0$. Suppose $y_{t}\in\{0,1\}$ arrives with $\mathbb{P}(y_{t}=1\mid a_{t}=1,\theta)=e^{*}(1,\pi_{t})\,p_{\theta}\,\Delta+o(\Delta)$ and $\mathbb{P}(y_{t}=1\mid a_{t}=0,\theta)=o(\Delta)$. Define the log-likelihood increment $\Delta L_{t+\Delta}$. Standard diffusion approximations yield:
Lemma 58.
As $\Delta\to 0$, under $a_{t}=1$ the process $(L_{t})_{t\geq 0}$ converges weakly to a Lévy process with drift $\mu_{\theta}(\pi_{t})=e^{*}(1,\pi_{t})\,[p_{H}(\pi_{t})-p_{L}(\pi_{t})]$ and bounded jumps at arrival times of $y=1$. If additionally $e^{*}(1,\pi)$ and $p_{\theta}(\pi)$ are $C^{1}$, then $L_{t}$ admits an Itô decomposition with locally Lipschitz coefficients on any compact set where $a_{t}\equiv 1$.
We now give a functional central limit theorem (in the semimartingale sense) for the log-odds process.
Theorem 59.
Fix a finite horizon $T>0$ and a vanishing period length $\Delta\downarrow 0$. For each $\Delta$, let $(\mathcal{F}^{\Delta}_{k\Delta})_{k\geq 0}$ be the public filtration. Define the (càdlàg, piecewise-constant) log-odds process
$L^{\Delta}_{t}\;=\;\sum_{k<t/\Delta}\Delta L^{\Delta}_{k+1},\qquad\Delta L^{\Delta}_{k+1}\in\{0,J^{+}(\pi^{\Delta}_{k\Delta}),J^{-}(\pi^{\Delta}_{k\Delta})\},$
where $J^{\pm}(\cdot)$ are the success/failure log-likelihood jumps (Appendix C), uniformly bounded and locally Lipschitz on $(0,1)$, and $\pi^{\Delta}_{k\Delta}$ is the public reputation just before time $k\Delta$. Write
$m^{\Delta}_{k}\;\equiv\;\mathbb{E}\!\left[\Delta L^{\Delta}_{k+1}\mid\mathcal{F}^{\Delta}_{k\Delta}\right],\qquad\xi^{\Delta}_{k}\;\equiv\;\Delta L^{\Delta}_{k+1}-m^{\Delta}_{k},$
and define the martingale
$M^{\Delta}(t)\;\equiv\;\sum_{k<t/\Delta}\xi^{\Delta}_{k},\qquad t\in[0,T],$
and the drift process $B^{\Delta}(t)\equiv\sum_{k<t/\Delta}m^{\Delta}_{k}$ so that $L^{\Delta}_{t}=L^{\Delta}_{0}+B^{\Delta}(t)+M^{\Delta}(t)$.
Assume the following intensity scaling and regularity hold:
  • (H1)
    There exist bounded, locally Lipschitz maps $\rho,\Lambda:(0,1)\to[0,\infty)$ such that
    $\begin{split}\mathbb{P}(a_{k\Delta}&=1\mid\mathcal{F}^{\Delta}_{k\Delta})=\rho(\pi^{\Delta}_{k\Delta})\,\Delta+o_{p}(\Delta),\\ \mathbb{P}(y_{k\Delta}&=1\mid a_{k\Delta}=1,\mathcal{F}^{\Delta}_{k\Delta})=\Lambda(\pi^{\Delta}_{k\Delta})+o_{p}(1),\end{split}$
    uniformly on $[0,T]$ in probability. In particular, the compensators
    $A^{+,\Delta}(t)\;\equiv\;\sum_{k<t/\Delta}\rho(\pi^{\Delta}_{k\Delta})\Lambda(\pi^{\Delta}_{k\Delta})\,\Delta,\quad A^{-,\Delta}(t)\;\equiv\;\sum_{k<t/\Delta}\rho(\pi^{\Delta}_{k\Delta})\big{(}1-\Lambda(\pi^{\Delta}_{k\Delta})\big{)}\,\Delta$
    converge in probability, uniformly on $[0,T]$ (u.c.p.), to
    $A^{+}(t)=\int_{0}^{t}\rho(\pi_{s})\Lambda(\pi_{s})\,ds,\qquad A^{-}(t)=\int_{0}^{t}\rho(\pi_{s})\big{(}1-\Lambda(\pi_{s})\big{)}\,ds.$
  • (H2)
    There is $C<\infty$ such that $|J^{\pm}(\pi)|\leq C$ for all $\pi\in(0,1)$; moreover $J^{\pm}$ are locally Lipschitz in $\pi$.
  • (H3)
    The conditional means satisfy
    $B^{\Delta}(t)\;=\;\sum_{k<t/\Delta}m^{\Delta}_{k}\;=\;\int_{0}^{t}\mu(\pi^{\Delta}_{s})\,ds+o_{p}(1)\quad\text{u.c.p. on }[0,T],$
    for a bounded, locally Lipschitz $\mu:(0,1)\to\mathbb{R}$.
Then, as $\Delta\to 0$, the pair $\big{(}L^{\Delta},\,\nu^{\Delta}\big{)}$ converges weakly in $\mathbb{D}([0,T])\times\mathbb{M}$ to a special semimartingale $(L,\nu)$ with characteristics
$\text{\rm(drift)}\;\;B(t)=\int_{0}^{t}\mu(\pi_{s})\,ds,\qquad\text{\rm(continuous covariation)}\;\;C\equiv 0,$
and jump compensator
$\nu(dt,dx)\;=\;\rho(\pi_{t})\,\Lambda(\pi_{t})\,dt\,\delta_{J^{+}(\pi_{t})}(dx)\;+\;\rho(\pi_{t})\,\big{(}1-\Lambda(\pi_{t})\big{)}\,dt\,\delta_{J^{-}(\pi_{t})}(dx),$
where $\delta_{z}$ denotes the Dirac measure at $z$. In particular,
$L_{t}\;=\;L_{0}\;+\;\int_{0}^{t}\mu(\pi_{s})\,ds\;+\;J^{+}(\pi)\,\Big{(}N^{+}(t)-A^{+}(t)\Big{)}\;+\;J^{-}(\pi)\,\Big{(}N^{-}(t)-A^{-}(t)\Big{)},$
with $(N^{+},N^{-})$ independent inhomogeneous Poisson processes conditionally on $(\pi_{s})_{s\leq t}$ and compensators $A^{\pm}$ as above. The predictable quadratic variation satisfies
$\langle M\rangle(t)\;=\;\int_{0}^{t}\sigma^{2}(\pi_{s})\,ds,\qquad\sigma^{2}(\pi)\;=\;\rho(\pi)\Big{[}\Lambda(\pi)\big{(}J^{+}(\pi)\big{)}^{2}+\big{(}1-\Lambda(\pi)\big{)}\big{(}J^{-}(\pi)\big{)}^{2}\Big{]}.$
Small-jump regime. If, in addition, $\sup_{\pi\in(0,1)}\big{(}|J^{+}(\pi)|\vee|J^{-}(\pi)|\big{)}\to 0$ as $\Delta\to 0$, then $M^{\Delta}\Rightarrow W$ where $W$ is a continuous Gaussian martingale with $W](t)=\int_{0}^{t}\sigma^{2}(\pi_{s})\,ds$; hence
$L^{\Delta}\;\Rightarrow\;L_{0}\,+\,\int_{0}^{t}\mu(\pi_{s})\,ds\,+\,W(t)\quad\text{in }\mathbb{D}([0,T]).$
Proof.
Step 1. For each $\Delta$ and $k$, condition on $\mathcal{F}^{\Delta}_{k\Delta}$. Given $a_{k\Delta}\in\{0,1\}$, the next-period outcome $y_{k\Delta}\in\{0,1\}$ yields $\Delta L^{\Delta}_{k+1}\in\{0,J^{+},J^{-}\}$ with probabilities
$\mathbb{P}(\Delta L^{\Delta}_{k+1}=J^{+}|\mathcal{F}^{\Delta}_{k\Delta})\;=\;\rho(\pi^{\Delta}_{k\Delta})\,\Lambda(\pi^{\Delta}_{k\Delta})\,\Delta+o_{p}(\Delta),$
$\mathbb{P}(\Delta L^{\Delta}_{k+1}=J^{-}|\mathcal{F}^{\Delta}_{k\Delta})\;=\;\rho(\pi^{\Delta}_{k\Delta})\big{(}1-\Lambda(\pi^{\Delta}_{k\Delta})\big{)}\,\Delta+o_{p}(\Delta),$
and $\mathbb{P}(\Delta L^{\Delta}_{k+1}=0|\mathcal{F}^{\Delta}_{k\Delta})=1-O_{p}(\Delta)$, uniformly on $[0,T]$ in probability by (H1). By construction, $M^{\Delta}$ is a square-integrable martingale with respect to the càdlàg filtration $(\mathcal{F}^{\Delta}_{t})$, with jumps $\Delta M^{\Delta}_{(k+1)\Delta}=\xi^{\Delta}_{k}$ and predictable quadratic variation
$\langle M^{\Delta}\rangle(t)\;=\;\sum_{k<t/\Delta}\mathbb{E}\!\left[(\xi^{\Delta}_{k})^{2}\mid\mathcal{F}^{\Delta}_{k\Delta}\right].$
Since $|m^{\Delta}_{k}|=O_{p}(\Delta)$ and $|\Delta L^{\Delta}_{k+1}|\leq C$ by (H2), we have
$\mathbb{E}\!\left[(\xi^{\Delta}_{k})^{2}\mid\mathcal{F}^{\Delta}_{k\Delta}\right]=\mathbb{E}\!\left[(\Delta L^{\Delta}_{k+1})^{2}\mid\mathcal{F}^{\Delta}_{k\Delta}\right]+O_{p}(\Delta^{2})$
and thus, uniformly on $[0,T]$ in probability,
$\langle M^{\Delta}\rangle(t)\;=\;\sum_{k<t/\Delta}\Big{[}(J^{+})^{2}\rho\Lambda\,+\,(J^{-})^{2}\rho(1-\Lambda)\Big{]}(\pi^{\Delta}_{k\Delta})\,\Delta+o_{p}(1)\;\xrightarrow[\Delta\to 0]{\mathbb{P}}\;\int_{0}^{t}\sigma^{2}(\pi_{s})\,ds,$
where $\sigma^{2}$ is as stated and we used Riemann-sum convergence plus boundedness/Lipschitz continuity (H1)–(H2). Similarly, by (H3) $B^{\Delta}(\cdot)\to\int_{0}^{\cdot}\mu(\pi_{s})\,ds$ u.c.p. Hence the predictable characteristics of $L^{\Delta}$ (with truncation function $h(x)\equiv x$) are
$B^{\Delta}(t)\to B(t),\qquad C^{\Delta}(t)\equiv 0\to 0,\qquad\nu^{\Delta}(dt,dx)\Rightarrow\nu(dt,dx),$
where
$\begin{split}\nu^{\Delta}(dt,dx)\;=\;\sum_{k<t/\Delta}\Big{\{}&\rho(\pi^{\Delta}_{k\Delta})\Lambda(\pi^{\Delta}_{k\Delta})\,\Delta\,\delta_{J^{+}(\pi^{\Delta}_{k\Delta})}(dx)\\ +&\rho(\pi^{\Delta}_{k\Delta})\big{(}1-\Lambda(\pi^{\Delta}_{k\Delta})\big{)}\,\Delta\,\delta_{J^{-}(\pi^{\Delta}_{k\Delta})}(dx)\Big{\}}.\end{split}$
By (H1)–(H2), $\nu^{\Delta}$ converges in probability (in the sense of measures on $0,T]\times\mathbb{R}$ endowed with the vague topology) to
$\nu(dt,dx)=\rho(\pi_{t})\Lambda(\pi_{t})\,dt\,\delta_{J^{+}(\pi_{t})}(dx)+\rho(\pi_{t})\big{(}1-\Lambda(\pi_{t})\big{)}\,dt\,\delta_{J^{-}(\pi_{t})}(dx).$
Step 2. By Aldous’ criterion, it suffices to show that for any sequence of $(\mathcal{F}^{\Delta}_{t})$-stopping times $\tau_{\Delta}\leq T$ and any $\delta_{\Delta}\downarrow 0$,
$L^{\Delta}_{\tau_{\Delta}+\delta_{\Delta}}-L^{\Delta}_{\tau_{\Delta}}\;\xrightarrow{\mathbb{P}}\;0.$
Since $L^{\Delta}=B^{\Delta}+M^{\Delta}$ and $B^{\Delta}$ converges u.c.p. to a continuous limit, it suffices to control $M^{\Delta}$. For the martingale increments,
$\mathbb{E}\Big{[}\big{(}M^{\Delta}_{\tau_{\Delta}+\delta_{\Delta}}-M^{\Delta}_{\tau_{\Delta}}\big{)}^{2}\ \big{|}\ \mathcal{F}^{\Delta}_{\tau_{\Delta}}\Big{]}=\mathbb{E}\Big{[}\langle M^{\Delta}\rangle(\tau_{\Delta}+\delta_{\Delta})-\langle M^{\Delta}\rangle(\tau_{\Delta})\ \big{|}\ \mathcal{F}^{\Delta}_{\tau_{\Delta}}\Big{]}.$
By the u.c.p. convergence of $\langle M^{\Delta}\rangle$ established above and boundedness of $\sigma^{2}$, the right-hand side is bounded by $K\,\delta_{\Delta}+o_{p}(1)$ for some $K<\infty$. Hence the conditional second moment of the increment tends to 0 in probability, which implies Aldous tightness (see, e.g., Ethier and Kurtz (1986), Thm. 3.8.6). Therefore $\{L^{\Delta}\}$ is tight in $\mathbb{D}([0,T])$.
Step 3. Let $L$ be any weak limit point along a subsequence $\Delta_{n}\downarrow 0$. By Step 1, the characteristics $(B^{\Delta_{n}},C^{\Delta_{n}},\nu^{\Delta_{n}})$ converge in probability to $(B,0,\nu)$ uniformly on compacts. By the general convergence theorem for semimartingales (e.g., Jacod and Shiryaev (2003), Thm. IX.3.9), $L$ is a special semimartingale with characteristics $(B,0,\nu)$ and admits the decomposition
$L_{t}\;=\;L_{0}\;+\;\int_{0}^{t}\mu(\pi_{s})\,ds\;+\;\int_{0}^{t}\!\!\int x\,\big{(}\mu^{L}(ds,dx)-\nu(ds,dx)\big{)},$
where $\mu^{L}$ is the jump measure of $L$. Since $\nu$ is a.s. absolutely continuous w.r.t. $dt$ with atoms at $J^{\pm}(\pi_{t})$ and finite mass on bounded sets, $L$ can be represented (conditionally on the path $(\pi_{s})$) as a sum of two compensated inhomogeneous Poisson integrals with jump sizes $J^{\pm}(\pi_{t})$ and intensities $\rho(\pi_{t})\Lambda(\pi_{t})$ and $\rho(\pi_{t})\big{(}1-\Lambda(\pi_{t})\big{)}$, respectively. In particular, the martingale part has predictable quadratic variation $\langle M\rangle(t)=\int_{0}^{t}\sigma^{2}(\pi_{s})\,ds$.
Step 4. Because the limiting characteristics are unique and continuous functionals of the (prelimit) characteristics, every subsequence has the same law for any limit point; hence the entire sequence converges to the unique limit characterized above.
If $\sup_{\pi}(|J^{+}(\pi)|\vee|J^{-}(\pi)|)\to 0$, then the jump compensator $\nu^{\Delta}$ assigns vanishing mass to $\{|x|>\varepsilon\}$ for any fixed $\varepsilon>0$. The martingale functional CLT (e.g., Rebolledo (1980); Hall and Heyde (1980), Thm. 3.2) applies: $M^{\Delta}$ converges to a continuous Gaussian martingale $W$ with quadratic variation given by the limit of $\langle M^{\Delta}\rangle$, namely $\int_{0}^{t}\sigma^{2}(\pi_{s})\,ds$. Adding the drift limit from (H3) yields the stated continuous-limit representation of $L^{\Delta}$. ∎
Proposition 60.
Let $V(\pi)$ be $C^{1}$ and convex. In the diffusion regime with constant experimentation ($a\equiv 1$), the expert’s value solves
$\rho V(\pi)=\phi+\mu(\pi)\,V^{\prime}(\pi)+\sup_{s\in S}\big{\{}\delta\,\mathbb{E}[V(\pi^{\prime})]-V(\pi)\big{\}},$
where $\mu(\pi)=\pi(1-\pi)\,e^{*}(1,\pi)\,[p_{H}(\pi)-p_{L}(\pi)]$ is the expected log-odds drift mapped to belief space. The discrete-time cutoff characterization in Theorem 5 converges to the bang-bang policy of the HJB.
Appendix I Contingent Bonus on Success
We formalize the effect of a success–contingent transfer on the expert’s policy and on learning. Suppose a bonus $\beta\geq 0$ is paid to the expert whenever $(a=1,y=1)$ occurs. Period utility is $\phi\,a+\beta\,\mathbf{1}\{a=1,y=1\}$; the implementer’s problem and effort $e^{*}(1,\pi;s)$ are unchanged (the contract affects only the expert’s utility). Standing Assumptions A1–A4 continue to hold.
I.1 Cutoff shift and comparative statics in $\beta$
Given public reputation $\pi\in(0,1)$ and signal $s$, the High type’s risky–minus–safe continuation difference with bonus $\beta$ equals
$\begin{split}\Delta_{H}^{\beta}(s;\pi)\;=\;\phi\;+\;\beta\,P_{S}(s;\pi)\;&+\;\delta\!\Big{\{}P_{S}(s;\pi)\,V\!\big{(}\pi^{+}(\pi,s)\big{)}\\ &+\big{(}1-P_{S}(s;\pi)\big{)}\,V\!\big{(}\pi^{-}(\pi,s)\big{)}-V\!\big{(}\pi^{\mathrm{rec}}(0;\pi,s)\big{)}\Big{\}},\end{split}$ (29)
where $P_{S}(s;\pi)$ is the success probability under $a=1$ (Section 3), $\pi^{\mathrm{rec}}(0;\pi,s)$ is the recommendation-only posterior after $a=0$, and $\pi^{\pm}(\pi,s)$ are the outcome posteriors after $a=1$ (Appendix A). Write the equilibrium cutoff as $s^{*}_{\beta}(\pi)$, defined by $\Delta_{H}^{\beta}\!\big{(}s^{*}_{\beta}(\pi);\pi\big{)}=0$.
Proposition 61.
Fix $\pi\in(0,1)$. Then:
  1. (i)
    For each $\beta\geq 0$, the map $s\mapsto\Delta_{H}^{\beta}(s;\pi)$ is continuous and strictly increasing (single crossing), hence there exists a unique cutoff $s^{*}_{\beta}(\pi)$ with $\Delta_{H}^{\beta}\!\big{(}s^{*}_{\beta}(\pi);\pi\big{)}=0$.
  2. (ii)
    $\beta\mapsto s^{*}_{\beta}(\pi)$ is continuous and (weakly) decreasing. If $P_{S}\!\big{(}s^{*}_{\beta}(\pi);\pi\big{)}\in(0,1)$ and $\partial_{s}\Delta_{H}^{\beta}\!\big{(}s^{*}_{\beta}(\pi);\pi\big{)}>0$, then $s^{*}_{\beta}(\pi)$ is strictly decreasing and differentiable in $\beta$, with
    $\frac{\partial s^{*}_{\beta}(\pi)}{\partial\beta}\;=\;-\;\frac{P_{S}\!\left(s^{*}_{\beta}(\pi);\pi\right)}{\partial_{s}\Delta_{H}^{\beta}\!\left(s^{*}_{\beta}(\pi);\pi\right)}\;<\;0.$
  3. (iii)
    If there exists $c(\pi)>0$ with $\partial_{s}\Delta_{H}^{\beta}(s;\pi)\geq c(\pi)$ for all $s$ in a neighborhood of $s^{*}_{\beta}(\pi)$ and all $\beta$ in a compact set $B\subset\mathbb{R}_{+}$, then for any $\beta_{1},\beta_{2}\in B$,
    $\big{|}s^{*}_{\beta_{2}}(\pi)-s^{*}_{\beta_{1}}(\pi)\big{|}\;\leq\;\frac{|\beta_{2}-\beta_{1}|}{c(\pi)}.$
Proof.
(i) The only change from $\Delta_{H}$ (Appendix A) is the additive term $\beta P_{S}(s;\pi)$; by Assumptions A1–A3 and Lemma 27, $P_{S}$ is continuous and strictly increasing in $s$, while the remaining (baseline) term is continuous and strictly increasing by Lemma 28. Existence/uniqueness follows.
(ii) Monotone comparative statics: for fixed $s$, $\partial_{\beta}\Delta_{H}^{\beta}(s;\pi)=P_{S}(s;\pi)\geq 0$. Since $s\mapsto\Delta_{H}^{\beta}(s;\pi)$ has strict single crossing, Topkis implies that the smallest root $s^{*}_{\beta}(\pi)$ is (weakly) decreasing in $\beta$. When interior, the implicit function theorem applies and yields the derivative displayed.
(iii) Integrate the IFT formula along a path from $\beta_{1}$ to $\beta_{2}$ and bound the denominator below by $c(\pi)$; use that $0\leq P_{S}\leq 1$. ∎
By the IFT formula, the local sensitivity satisfies $\big{|}\partial s^{*}_{\beta}(\pi)/\partial\beta\big{|}=P_{S}\!\big{(}s^{*}_{\beta}(\pi);\pi\big{)}/\partial_{s}\Delta_{H}^{\beta}\!\big{(}s^{*}_{\beta}(\pi);\pi\big{)}$. Since $P_{S}$ rises with reputation via effort (Lemma 27), the absolute effect is largest at reputations where $P_{S}$ is high (typically at higher $\pi$). Corners: if $P_{S}=0$ at the threshold, the bonus has no first-order effect at that $\pi$; if $P_{S}=1$ and $\partial_{s}\Delta_{H}^{\beta}>0$, the formula still applies.
I.2 Targeting a desired threshold
A minimal bonus that implements any target cutoff $\tilde{s}(\pi)$ (with $\Delta_{H}(\tilde{s}(\pi);\pi)<0$ at $\beta=0$) is
$\beta^{\min}(\pi,\tilde{s})\;=\;-\frac{\Delta_{H}\!\big{(}\tilde{s}(\pi);\pi\big{)}}{P_{S}\!\big{(}\tilde{s}(\pi);\pi\big{)}}\,$, (30)
with the convention that if $P_{S}(\tilde{s}(\pi);\pi)=0$ the local instrument is ineffective (no success can occur at the target). Any $\beta>\beta^{\min}$ strictly lowers the cutoff relative to $\tilde{s}(\pi)$.
I.3 Effect on the speed of learning (KL drift)
Let $L_{t+1}$ be the one-step likelihood ratio between period-(t+1)(t{+}1) observations under $H$ and $L$ (Appendix C). Define the per-period expected Kullback–Leibler drift at reputation $\pi$ under contract $\beta$ by
$\begin{split}\kappa(\pi;\beta)\;&\equiv\;\mathbb{E}\!\left[\log L_{t+1}\,\middle|\,\pi_{t}=\pi;\beta\right]\\ \;&=\;\mathbb{P}(a_{t}=1\mid\pi;\beta)\cdot D_{\mathrm{KL}}\!\big{(}\mathcal{L}_{H}(y\mid a{=}1,\pi)\,\|\,\mathcal{L}_{L}(y\mid a{=}1,\pi)\big{)}\;+\;\text{(rec. term)},\end{split}$
where the “rec. term” collects the (nonnegative) KL from observing the recommendation itself when it is informative about type. The outcome KL is strictly positive whenever $p_{H}(\pi,s)\neq p_{L}(\pi,s)$ on a set of positive probability (Assumption A1).
Proposition 62.
At $\beta=0$, a marginal increase in the bonus strictly raises the expected KL drift:
$\begin{split}\left.\frac{d}{d\beta}\,\mathbb{E}\big{[}\kappa(\pi_{t};\beta)\big{]}\right|_{\beta=0}\;&=\;\mathbb{E}\!\left[\frac{\partial\,\mathbb{P}(a_{t}=1\mid\pi_{t};\beta)}{\partial\beta}\Big{|}_{\beta=0}\right]\\ &\cdot\underbrace{D_{\mathrm{KL}}\!\big{(}\mathcal{L}_{H}(y\mid a{=}1,\pi_{t})\,\|\,\mathcal{L}_{L}(y\mid a{=}1,\pi_{t})\big{)}}_{>0}\;+\;\text{\emph{(rec. term)}}\;>\;0,\end{split}$
provided recommendations are not exactly uninformative at $\beta=0$ (in which case the “rec. term” vanishes).
Proof.
By Proposition 61, $\partial_{\beta}s^{*}_{\beta}(\pi)\big{|}_{\beta=0}<0$ wherever $P_{S}\big{(}s^{*}(\pi);\pi\big{)}>0$, so the risky region in $s$ (hence $\mathbb{P}(a=1\mid\pi;\beta)$) expands at $\beta=0$. Each experiment contributes strictly positive outcome KL by Assumption A1 (the Bernoulli success parameters under $H$ and $L$ differ on a set of positive probability). Linearity of expectation gives the display. The recommendation KL term is weakly positive and independent of $\beta$ to first order unless the bonus also shifts recommendation informativeness directly; here it does not, since the agent’s problem is unchanged. ∎
I.4 Affine success/failure contracts
More generally, consider an affine contract paying $\beta_{1}$ upon (a=1,y=1)(a{=}1,y{=}1) and $\beta_{0}$ upon (a=1,y=0)(a{=}1,y{=}0), with $\beta_{1},\beta_{0}\geq 0$. Then
$\Delta_{H}^{\beta_{1},\beta_{0}}(s;\pi)\;=\;\phi\;+\;\big{[}\beta_{1}P_{S}(s;\pi)-\beta_{0}(1-P_{S}(s;\pi))\big{]}\;+\;\delta\Big{\{}\cdots\Big{\}},$ (31)
so all arguments above go through with the additive rotation $\beta_{1}P_{S}-\beta_{0}(1-P_{S})]$. In particular, increasing $\beta_{1}$ (holding $\beta_{0}$ fixed) lowers the cutoff, and increasing $\beta_{0}$ raises it; the IFT derivative becomes
$\frac{\partial s^{*}_{\beta_{1},\beta_{0}}(\pi)}{\partial\beta_{1}}=-\,\frac{P_{S}\!\left(s^{*}_{\beta_{1},\beta_{0}}(\pi);\pi\right)}{\partial_{s}\Delta_{H}^{\beta_{1},\beta_{0}}\!\left(s^{*}_{\beta_{1},\beta_{0}}(\pi);\pi\right)},\qquad\frac{\partial s^{*}_{\beta_{1},\beta_{0}}(\pi)}{\partial\beta_{0}}=\;\;\frac{1-P_{S}\!\left(s^{*}_{\beta_{1},\beta_{0}}(\pi);\pi\right)}{\partial_{s}\Delta_{H}^{\beta_{1},\beta_{0}}\!\left(s^{*}_{\beta_{1},\beta_{0}}(\pi);\pi\right)}.$
The bonus rotates the risky–safe margin by an amount proportional to $P_{S}$, so it is most potent where effort (hence $P_{S}$) is high. Because experiments carry positive KL, any policy that increases their frequency speeds up learning; the appendix’s formulas make these effects transparent and easy to calibrate.
Online Appendix
Appendix OA.1 Value monotonicity and convexity: full proof
This appendix provides a complete proof of Theorem 4. We first show that for any fixed Markov cutoff policy $s(\cdot)$, the associated Bellman operator preserves increasingness and convexity, and is a contraction in the sup norm. We then consider the equilibrium policy $s^{*}(\cdot)$ and conclude that the equilibrium value $V=V^{\,s^{*}}$ inherits these properties.
Lemma 63 (Monotonicity and contraction).
Fix a measurable cutoff policy $s(\cdot)$. Define $(T^{\,s}V)(\pi)\equiv\phi\,\rho(\pi;s)+\delta\,\mathbb{E}\big{[}V(\pi^{\prime})\mid\pi,s\big{]}$, where $\rho(\pi;s)$ is the risky recommendation frequency induced by $s(\cdot)$ at public belief $\pi$ and $\pi^{\prime}$ is the posterior after the publicly observed history. Then $T^{\,s}$ is monotone: if $V_{1}\leq V_{2}$, then $T^{\,s}V_{1}\leq T^{\,s}V_{2}$. Moreover, $\|T^{\,s}V_{1}-T^{\,s}V_{2}\|_{\infty}\leq\delta\,\|V_{1}-V_{2}\|_{\infty}$.
Proof.
Monotonicity is immediate since the flow term $\phi\,\rho(\pi;s)$ does not depend on $V$ and the expectation is taken with nonnegative weights. The contraction inequality follows from the discount factor: $\big{|}(T^{\,s}V_{1})(\pi)-(T^{\,s}V_{2})(\pi)\big{|}=\delta\,\big{|}\mathbb{E}[V_{1}(\pi^{\prime})-V_{2}(\pi^{\prime})\mid\pi,s]\big{|}\leq\delta\,\|V_{1}-V_{2}\|_{\infty}$. ∎
Lemma 64 (Convexity preservation).
Suppose (A1)–(A3) hold and Condition (6) is satisfied at the policy $s(\cdot)$. If $V$ is convex, then $\pi\mapsto(T^{\,s}V)(\pi)$ is convex.
Proof.
Write $(T^{\,s}V)(\pi)=\phi\,\rho(\pi;s)+\delta\,\Xi(\pi)$ with $\Xi(\pi)\equiv\mathbb{E}[V(\pi^{\prime})\mid\pi,s]$. The flow term is affine in $\rho(\pi;s)$ and does not affect convexity in $V$. It suffices to show $\Xi$ is convex. Under a fixed cutoff policy, the public posterior after $a=1$ takes two values $\{\pi^{+}(\pi),\pi^{-}(\pi)\}$ corresponding to $y\in\{1,0\}$; after $a=0$ it is the recommendation-only posterior $\tilde{\pi}(\pi)$. Both $\{\pi^{+}(\pi),\pi^{-}(\pi)\}$ and $\tilde{\pi}(\pi)$ are increasing in prior odds (Bayes). Condition (6) ensures that, as $\pi$ increases, the two-point distribution on $\{\pi^{+}(\pi),\pi^{-}(\pi)\}$ becomes a mean-preserving spread in the convex order (the failure branch’s log-likelihood jump weakly dominates the success branch). Therefore, for any convex $V$, the map $\pi\mapsto\mathbb{E}[V(\pi^{\prime})\mid a=1,\pi,s]$ is convex. The recommendation-only posterior yields a single-valued update; composing a convex $V$ with an increasing fractional-linear map preserves convexity locally and, combined with the dominant two-branch term under $a=1$, yields global convexity of $\Xi$. Formalizing this, for any $\pi_{1},\pi_{2}$ and $\lambda\in[0,1]$, construct the standard coupling of the $(a,y)$ draws and use the convex-order inequality to obtain
$\Xi(\lambda\pi_{1}+(1-\lambda)\pi_{2})\;\leq\;\lambda\,\Xi(\pi_{1})+(1-\lambda)\,\Xi(\pi_{2}).$
Hence $\Xi$ is convex and so is $T^{\,s}V$. ∎
Proof of Theorem 4.
Fix any cutoff policy $s(\cdot)$. By Lemma 63, $T^{\,s}$ is a monotone contraction; by Lemma 64, it preserves convexity. Starting from a bounded convex $V_{0}$ (e.g., $V_{0}\equiv 0$), the Picard iterates $V_{n+1}=T^{\,s}V_{n}$ are convex and converge uniformly to the unique fixed point $V^{\,s}$, which is thus convex and increasing. In equilibrium, $s^{*}(\cdot)$ maximizes the Bellman expression pointwise; since the supremum of convex functions is convex, $V=\sup_{s}V^{\,s}$ evaluated at $s^{*}$ is increasing and convex. ∎
Appendix OA.2 Empirical Measurement of Reputation and Effort
This appendix gives a compact recipe to build (i) a panel measure of reputation for each expert over time and (ii) empirical proxies for effort on the implementer side. The goal is to operationalize the testable predictions in the main text while avoiding mechanical look-ahead or reflection problems.
OA.2.1 Reputation from binary outcomes: Beta–Bernoulli with leave-one-out
[…]
$S_{i,t}^{(-t)}=\sum_{\tau<t}y_{i\tau},\qquad N_{i,t}^{(-t)}=\sum_{\tau<t}1,$
so the information set at $t$ excludes the current case. The posterior mean success rate for $i$ at $t$ is then
$\text{Rep}_{it}^{\text{BB}}\;=\;\frac{\alpha_{0}+S_{i,t}^{(-t)}}{\alpha_{0}+\beta_{0}+N_{i,t}^{(-t)}}\in(0,1).$ (OA.1)
This is a shrinkage estimator: with few past cases, $\text{Rep}_{it}^{\text{BB}}$ remains close to the prior mean; as $N_{i,t}^{(-t)}$ grows, it approaches the raw frequency $S_{i,t}^{(-t)}/N_{i,t}^{(-t)}$. When case difficulty varies observably (risk adjustment), use Section OA.2.2.
Practical notes
(i) Compute (OA.1) in rolling calendar time (or event time) with exact leave-one-out to avoid using $y_{it}$ in its own predictor. (ii) If experts enter/exit, carry the prior forward; do not backfill. (iii) Cluster subsequent inference at the expert level.
OA.2.2 Risk-adjusted reputation: hierarchical logit with case-mix
When outcomes depend on observables $x_{i\tau}$ (case mix), estimate a hierarchical logit
$\mathbb{P}\big{(}y_{i\tau}=1\mid x_{i\tau}\big{)}=\frac{1}{1+\exp\!\big{(}-(\alpha_{i}+x_{i\tau}^{\prime}\beta)\big{)}},\qquad\alpha_{i}\sim\mathcal{N}(\mu_{\alpha},\sigma_{\alpha}^{2}).$
Fit the model by GLMM (Laplace or adaptive quadrature) on a training sample that excludes the current fold of observations for each $i$ (cross-fold or rolling-split). Let $\widehat{\alpha}_{i}^{\,\text{post}}$ be the Empirical Bayes posterior mean of $\alpha_{i}$. Define the risk-adjusted reputation at time $t$ by transforming expert $i$’s intercept:
$\text{Rep}_{it}^{\text{HL}}\;=\;\frac{1}{1+\exp\!\big{(}-(\widehat{\mu}_{\alpha}+\widehat{\alpha}_{i}^{\,\text{post}})\big{)}}\in(0,1).$ (OA.2)
For small-panel experts, $\widehat{\alpha}_{i}^{\,\text{post}}$ shrinks toward $\widehat{\mu}_{\alpha}$; with abundant data it approaches the individual MLE.
OA.2.2.1 Cross-fold construction
Partition time (or cases) into $K$ folds. For each fold $k$, fit the model on the complement and produce $\widehat{\alpha}_{i}^{\,\text{post}(-k)}$ for cases in fold $k$. Concatenate across $k$. This avoids look-ahead and ensures $\text{Rep}_{it}$ uses only information available before $t$.
OA.2.3 Effort proxies and monotone mapping
The model requires only that observed effort proxies are monotone in the effort $e^{*}$. Suggested domain-specific proxies:
  • Medicine. Adherence indices (proportion of days covered; medication possession ratio), therapy completion indicators, pre-/post-operative compliance scores, scheduled vs. attended rehabilitation sessions.
  • Finance. Trade size (or turnover) conditional on recommendation strength; fraction of the book adjusted; execution slippage (smaller slippage indicates higher attention/effort).
  • R&D/Policy. Staffed-hours or milestone on-time completion rates following a “greenlight”; number of design iterations completed before deadline.
Map raw proxies $E_{it}$ into $[0,1]$ by a monotone transformation, e.g. $e_{it}=\frac{E_{it}-\min}{\max-\min}$ or rank-normalization. Comparative-statics tests (e.g. higher reputation $\Rightarrow$ higher effort when $a=1$) are invariant to any strictly monotone remapping.
OA.2.4 Design choices to avoid mechanical feedback
  1. 1.
    Leave-one-out / sample splitting. Construct $\text{Rep}_{it}$ using only $\{\tau<t\}$ or out-of-fold data; never include contemporaneous $y_{it}$.
  2. 2.
    Event-time windows. For shock analyses (early successes/failures), use symmetric windows around the event; exclude the event itself from the predictor on the left-hand side.
  3. 3.
    Controls. Include case mix $x_{i\tau}$ and calendar/time fixed effects to separate secular trends.
  4. 4.
    Clustering. Cluster standard errors at the expert level (and two-way cluster by expert $\times$ time block if serial correlation is a concern).
OA.2.5 From data to model objects
OA.2.5.1 Estimating $\rho(\pi)$
Bin observations by pre-period $\text{Rep}_{it}$ (e.g., deciles). Within bin $b$, estimate the risky-recommendation frequency $\widehat{\rho}_{b}=\mathbb{P}(a_{it}=1\mid\text{Rep}_{it}\in b)$ and report $\widehat{\rho}(\pi_{b})$ at the bin midpoints $\pi_{b}$.
OA.2.5.2 Estimating $P_{S}(\cdot;\pi)$
Within each bin $b$, estimate the success rate conditional on $a=1$: $\widehat{P}_{S,b}=\mathbb{P}(y_{it}=1\mid a_{it}=1,\text{Rep}_{it}\in b)$. If effort proxies are available, also report $\mathbb{E}[e_{it}\mid a_{it}=1,\text{Rep}_{it}\in b]$ to document the mechanism.
OA.2.5.3 Minimal bonus calibration
Given a target threshold $\tilde{s}(\pi)$ (e.g., a policy target for experimentation), the minimal success bonus that implements it is
$\beta^{\min}(\pi,\tilde{s})\;=-\;\frac{\,\Delta_{H}\!\big{(}\tilde{s}(\pi);\pi\big{)}}{P_{S}\!\big{(}\tilde{s}(\pi);\pi\big{)}}\,$,
Empirically, approximate the numerator by the model-implied continuation gap at the target (or by a calibrated proxy), and plug in $\widehat{P}_{S,b}$ for the denominator at the corresponding reputation bin.
OA.2.6 Testing the model’s predictions
OA.2.6.1 Reputational conservatism
Test that risky calls are less frequent at higher reputation:
$a_{it}=\gamma_{0}+\gamma_{1}\,\text{Rep}_{it}+\gamma_{2}^{\prime}x_{it}+\eta_{i}+\tau_{t}+\varepsilon_{it},\qquad\gamma_{1}<0,$
where $\eta_{i}$ and $\tau_{t}$ are expert and time fixed effects. Use a logit/probit or linear probability model; inference clustered at $i$.
OA.2.6.2 Higher conditional success at high reputation
Among $a_{it}=1$,
$y_{it}=\delta_{0}+\delta_{1}\,\text{Rep}_{it}+\delta_{2}^{\prime}x_{it}+\eta_{i}+\tau_{t}+u_{it},\qquad\delta_{1}>0.$
If effort proxies $e_{it}$ are available, show $\partial\mathbb{E}[e_{it}\mid a_{it}=1]/\partial\text{Rep}_{it}>0$ and that controlling for $e_{it}$ attenuates $\delta_{1}$ (mediation via effort).
OA.2.6.3 Asymmetric learning from success vs failure
Estimate the absolute reputation change after outcomes at different reputation levels:
$\big{|}\text{Rep}_{i,t+1}-\text{Rep}_{it}\big{|}=\zeta_{0}+\zeta_{1}\,\mathbf{1}\{y_{it}=0\}+\zeta_{2}\,\text{Rep}_{it}+\zeta_{3}\,\mathbf{1}\{y_{it}=0\}\cdot\text{Rep}_{it}+\cdots$
and test $\zeta_{3}>0$ (failures hurt more at high reputation), consistent with the model’s $J^{-}$ asymmetry.
OA.2.7 Summary checklist
  • Build $\text{Rep}_{it}$ by (OA.1) or (OA.2) with strict leave-one-out / cross-fold.
  • Normalize effort proxies to $[0,1]$ and use only monotone mappings for comparative statics.
  • Bin by reputation to visualize $\rho(\pi)$ and $P_{S}(\cdot;\pi)$; then estimate regression counterparts with fixed effects and clustered SEs.
  • For policy calibration, use $\widehat{P}_{S}$ in the minimal-bonus formula (30).
Appendix OA.3 Algorithms and Replication
This appendix records the numerical procedures used in Section 4 for computing the equilibrium value function $V(\pi)$, the High type’s threshold $s^{*}(\pi)$, the experimentation rate $\rho(\pi)$, and the reputation-path simulations for the Gaussian benchmark. All objects and notation match the main text.
OA.3.1 Bellman equation, grids, and convergence
We solve the recursive problem on a uniform reputation grid $\Pi=\{\pi_{j}\}_{j=1}^{J}\subset(0,1)$ with $J\in\{201,321\}$ points (baseline $J=321$), covering $[0.05,0.95]$ to avoid numerical singularities at the boundaries. The per-period flow is $u(\pi)$; in the benchmark we set $u(\pi)=\pi^{2}$ (any increasing convex $u$ is admissible). The discounted value satisfies
$V(\pi)\;=\;u(\pi)\;+\;\delta\,\mathbb{E}\!\left[V(\pi^{\prime})\,\middle|\,\pi\right],$
where $\pi^{\prime}$ is the next-period reputation induced by the equilibrium recommendation rule and observed outcomes (Appendix A). We iterate on $V$ using damped value iteration: starting from $V^{(0)}(\pi)=u(\pi)/(1-\delta)$, construct $V^{\text{raw}}$ via the one-step Bellman update (detailed below), then set
$V^{(t+1)}\;=\;(1-\eta)\,V^{(t)}\;+\;\eta\,V^{\text{raw}},\qquad\eta\in[0.3,0.5],$
until $\|V^{(t+1)}-V^{(t)}\|_{\infty}<10^{-6}$. Linear interpolation on $\Pi$ is used whenever $V$ is evaluated at off-grid beliefs.
OA.3.2 Cutoff computation at a given $\pi$
Fix $\pi\in\Pi$. Let $s_{H}$ denote the candidate High-type threshold and $s_{L}=s_{H}+\Delta_{L}$ the Low-type threshold (in the Gaussian benchmark we take $\Delta_{L}=0.30$ so recommendations are informative). For any $(s_{H},s_{L})$:
  1. 1.
    Recommendation frequencies. With Gaussian signals $s\sim\mathcal{N}(\mu_{\omega},\sigma_{\theta}^{2})$, the risky recommendation tails are
    $A_{\theta}(s_{\theta})=1-\Phi\!\left(\frac{s_{\theta}-\mu_{1}}{\sigma_{\theta}}\right),\qquad B_{\theta}(s_{\theta})=1-\Phi\!\left(\frac{s_{\theta}-\mu_{0}}{\sigma_{\theta}}\right).$
    The unconditional risky frequency by type is $R_{\theta}=\lambda A_{\theta}+(1-\lambda)B_{\theta}$.
  2. 2.
    Post-recommendation reputation. Let $\pi^{\mathrm{rec}}(a)$ be the posterior about type after observing the recommendation $a\in\{0,1\}$:
    $\begin{split}\pi^{\mathrm{rec}}(1)&=\frac{\pi R_{H}}{\pi R_{H}+(1-\pi)R_{L}},\\ \pi^{\mathrm{rec}}(0)&=\frac{\pi(1-R_{H})}{\pi(1-R_{H})+(1-\pi)(1-R_{L})}.\end{split}$
  3. 3.
    Outcome branch under $a=1$. Conditional on $a=1$, the posterior success belief is
    $\begin{split}p_{\theta}&=\mathbb{P}(\omega=1\mid a=1,\theta)=\frac{\lambda A_{\theta}}{\lambda A_{\theta}+(1-\lambda)B_{\theta}}\,,\\ \lambda^{\mathrm{post}}&=\pi^{\mathrm{rec}}(1)\,p_{H}+(1-\pi^{\mathrm{rec}}(1))\,p_{L}.\end{split}$
    With quadratic cost $c(e)=\tfrac{1}{2}e^{2}$, the implementer’s best reply is $e^{*}=\lambda^{\mathrm{post}}$ and the success probability is $P_{S}=\lambda^{\mathrm{post}}\,e^{*}$. The log-likelihood jumps for reputation after observing $y\in\{0,1\}$ are
    $J^{+}=\log\frac{p_{H}}{p_{L}},\qquad J^{-}=\log\frac{1-e^{*}p_{H}}{1-e^{*}p_{L}}.$
    Hence the outcome posteriors are
    $\pi^{+}=\sigma\!\big{(}\operatorname{logit}(\pi^{\mathrm{rec}}(1))+J^{+}\big{)},\qquad\pi^{-}=\sigma\!\big{(}\operatorname{logit}(\pi^{\mathrm{rec}}(1))+J^{-}\big{)},$
    with $\sigma(x)=1/(1+e^{-x})$.
  4. 4.
    Risky–safe margin. The High type’s risky-minus-safe continuation difference is
    $\Delta_{H}(s_{H};\pi)=\phi+\delta\Big{(}P_{S}\,V(\pi^{+})+(1-P_{S})\,V(\pi^{-})-V(\pi^{\mathrm{rec}}(0))\Big{)}.$
  5. 5.
    Bisection on the cutoff. Because $\Delta_{H}(\cdot;\pi)$ is strictly increasing in $s_{H}$ (Appendix A), we find the unique $s^{*}(\pi)$ with $\Delta_{H}(s^{*}(\pi);\pi)=0$ by bisection on a wide bracket (e.g., $\mu_{0}-6\sigma_{H},\mu_{1}+6\sigma_{H}]$). If the function does not change sign within the bracket, we clamp to the nearest boundary.
The value-iteration update at this $\pi$ averages next-period values across the recommendation branch: with probability $R_{H}$ the High type recommends $a=1$ and realizes $P_{S}V(\pi^{+})+(1-P_{S})V(\pi^{-})$; with probability $1-R_{H}$ she recommends $a=0$ and realizes $V(\pi^{\mathrm{rec}}(0))$.
OA.3.3 Experimentation rate and figures
Given the equilibrium cutoffs $\{s^{*}(\pi_{j})\}_{j=1}^{J}$, the High type’s experimentation rate is
$\rho(\pi)\;=\;R_{H}\!\big{(}s^{*}(\pi)\big{)}\;=\;\lambda\,A_{H}\!\big{(}s^{*}(\pi)\big{)}+(1-\lambda)\,B_{H}\!\big{(}s^{*}(\pi)\big{)}.$
Figure panels in Section 4 were produced as follows:
  • Panel $s^{*}(\pi)$. Plot $\pi\mapsto s^{*}(\pi)$ on $\Pi$.
  • Panel $\rho(\pi)$. Plot $\pi\mapsto\rho(\pi)$ on $\Pi$.
Baseline parameters for the Gaussian benchmark:
$(\mu_{0},\mu_{1},\sigma_{H},\sigma_{L},\lambda,\delta)=(0,1,0.8,1.6,0.5,0.95),\quad\phi=0,\quad\text{and }\Delta_{L}=0.30.$
OA.3.4 Reputation-path simulations
To generate Figure 1 (reputation trajectories under $\theta=H$ and $\theta=L$):
  1. 1.
    Initialize $\pi_{0}=0.5$; fix a type $\theta\in\{H,L\}$; choose the number of replications $R$ (e.g., $R=250$) and horizon $T$ (e.g., $T=150$).
  2. 2.
    For each replication $r=1,\dots,R$ and period $t=0,\dots,T-1$:
    1. (a)
      Draw the period state $\omega_{t}\sim\text{Bernoulli}(\lambda)$ and the private signal $s_{t}\sim\mathcal{N}(\mu_{\omega_{t}},\sigma_{\theta}^{2})$.
    2. (b)
      Compute $a_{t}=\mathbbm{1}\{s_{t}\geq s^{*}(\pi_{t})\}$. If $a_{t}=0$, set $\pi_{t+1}=\pi^{\mathrm{rec}}(0)$.
    3. (c)
      If $a_{t}=1$, draw $y_{t}\sim\text{Bernoulli}(P_{S}(\pi_{t},s_{t}))$ and update
      $\pi_{t+1}=\begin{cases}\pi^{+}(\pi_{t},s_{t}),&\text{if }y_{t}=1,\\[2.5pt] \pi^{-}(\pi_{t},s_{t}),&\text{if }y_{t}=0.\end{cases}$
  3. 3.
    Plot the $R$ sample paths $\{\pi_{t}^{(r)}\}_{t=0}^{T}$ for each $\theta$.
Flat segments arise whenever $a_{t}=0$; jumps coincide with $a_{t}=1$ and realized outcomes.
OA.3.5 Implementation details
  • Interpolation. Linear interpolation on $\Pi$ for $V(\cdot)$ suffices; cubic interpolation can be used but is unnecessary.
  • Damping and tolerance. We use $\eta\in[0.3,0.5]$ and stop when $\|V^{(t+1)}-V^{(t)}\|_{\infty}<10^{-6}$ (results are unchanged at $10^{-7}$).
  • Bracketing and bisection. $80$ bisection iterations per $\pi$ are ample given the tolerance above; widening the bracket to $\mu_{0}-8\sigma_{H},\mu_{1}+8\sigma_{H}]$ guarantees a sign change in extreme calibrations.
  • Randomness. Fix random seeds for reproducibility. In the code used for figures, seeds are set once per panel (not per path) to ensure cross-panel comparability.
Code and data. A minimal replication script (value iteration, cutoff solver, and figure generation) is available upon request and mirrors the steps in this appendix exactly. No external data are required for the benchmark simulations.
Appendix OA.4 Alternative Primitives and Robustness
This appendix records robustness of our results to (i) alternative effort-cost primitives and (ii) outcome/observability frictions. Throughout we retain Assumptions A1 (MLRP/Blackwell dominance of $H$ over $L$), A3 (value $V$ increasing and convex), and the information structure of Section 3, except where explicitly modified.
OA.4.1 Effort costs beyond the quadratic
Let the implementer’s cost be any $C^{2}$ function $c:[0,\bar{e}]\to\mathbb{R}_{+}$, strictly convex with $c^{\prime}(0)\geq 0$, $c^{\prime\prime}(e)>0$ for $e\in[0,\bar{e})$, and a possibly finite effort cap $\bar{e}\in(0,1]$. The implementer’s best reply under $a=1$ is
$e^{*}(1,\pi;s)\;=\;\min\!\left\{\bar{e},\;\,(c^{\prime})^{-1}\!\big{(}\lambda(1,\pi;s)\big{)}\right\},$ (OA.2)
where $\lambda(1,\pi;s)$ is the success belief under risk before choosing effort.
Lemma 65.
Under A1 and the cost assumptions above, $e^{*}(1,\pi;s)$ is (weakly) increasing in $s$ and in $\pi$; hence $P_{S}(s;\pi)=\lambda(1,\pi;s)\,e^{*}(1,\pi;s)$ is (weakly) increasing in $s$ and in $\pi$.
Proof.
By A1, $s\mapsto\lambda(1,\pi;s)$ is increasing and, by the recommendation posterior, so is $\pi\mapsto\lambda(1,\cdot\,;s)$. The map $(c^{\prime})^{-1}$ is increasing by strict convexity of $c$, and the projection to $0,\bar{e}]$ preserves monotonicity. Multiplying by $\lambda$ preserves monotonicity of $P_{S}$. ∎
Proposition 66.
Fix $\pi$. With $c$ as above, $s\mapsto\Delta_{H}(s;\pi)$ remains continuous and (strictly) increasing except possibly at a kink where $e^{*}$ hits the cap $\bar{e}$. Consequently, a (unique) cutoff $s^{*}(\pi)$ exists. Moreover, $s^{*}(\pi)$ is (weakly) increasing in $\pi$ (reputational conservatism).
Proof.
The proof of Appendix A carries over using Lemma 65: $V(\pi^{\mathrm{rec}}(0;\pi,s))$ is increasing in $s$ and the outcome mixture $P_{S}(\cdot)\,V(\pi^{+})+(1-P_{S}(\cdot))V(\pi^{-})$ is increasing in $s$. Strictness holds wherever $e^{*}$ is interior; at a binding cap the function can be kinked but remains nondecreasing, so the smallest zero is unique. For conservatism, Appendix B uses only decreasing differences of the risky and safe branches in $(s,\pi)$, which are preserved because $e^{*}(1,\pi;s)$ is increasing in both arguments (Lemma 65) and the $J^{-}$ asymmetry in (12) continues to make failures more damaging when effort is higher. ∎
OA.4.1.1 What if $c^{\prime}$ is flat near 0?
If $c^{\prime}(e)=0$ on $0,\underline{e}]$ for some $\underline{e}>0$, then for any $\lambda(1,\pi;s)>0$ the best reply jumps to $e^{*}\in[\underline{e},\bar{e}]$ (highest payoff in the flat region), creating a kink in $s\mapsto e^{*}(1,\pi;s)$. Monotonicity survives and so do the main results (weak single crossing and conservatism), but $\Delta_{H}$ may be nondifferentiable at the corresponding $s$.
OA.4.1.2 Effort cap
If $\bar{e}<1$, then $e^{*}(1,\pi;s)$ saturates once $\lambda(1,\pi;s)\geq c^{\prime}(\bar{e})$. Threshold existence and conservatism are unaffected; at high reputation the additional marginal push from $\pi$ on $e^{*}$ vanishes, making reputational conservatism weaker in very high-reputation regions.
OA.4.2 Outcome technology and observability frictions
We now relax the outcome/observability primitives. Let $y\in\{0,1\}$ denote the latent technological success and $z\in\{0,1\}$ the observed success flag (possibly noisy). Unless stated, recommendations remain publicly observed.
OA.4.2.1 Baseline success under the safe action
Allow a baseline success risk $b\in[0,1)$ under $a=0$:
$\mathbb{P}(y=1\mid a=0)=b,\qquad\text{independent of }\theta,\ \omega.$
Because $b$ is type-independent, outcomes under $a=0$ carry no information about ability. The recommendation-only posterior $\pi^{\mathrm{rec}}(0;\pi,s)$ is unchanged, and the safe-branch continuation remains $V(\pi^{\mathrm{rec}}(0;\pi,s))$. Thus all results—threshold, conservatism, and reputation dynamics—go through unchanged. (Welfare and policy experiments that reward success would of course be affected by $b$.)
OA.4.2.2 Noisy success/failure classification
Suppose $z$ is a noisy indicator of $y$ with
$\alpha\equiv\mathbb{P}(z=1\mid y=1)\in(0,1],\qquad\beta\equiv\mathbb{P}(z=1\mid y=0)\in[0,1),\qquad\alpha>\beta.$
Under $a=1$, type–specific true success probabilities are $p_{\theta}(\pi,s)\,e^{*}(1,\pi;s)$, where $p_{\theta}(\pi,s)=\mathbb{P}(\omega=1\mid a=1,\theta,\pi,s)$. The observed success probability is
$q_{\theta}(\pi,s)\;=\;\alpha\,e^{*}p_{\theta}\;+\;\beta\,(1-e^{*}p_{\theta}).$
The reputation jumps upon observing $z$ are
$J^{\text{obs},+}(\pi,s)=\log\frac{q_{H}(\pi,s)}{q_{L}(\pi,s)},\qquad J^{\text{obs},-}(\pi,s)=\log\frac{1-q_{H}(\pi,s)}{1-q_{L}(\pi,s)}.$ (OA.3)
A direct calculation yields
$\frac{\partial J^{\text{obs},-}}{\partial e^{*}}\;=\;-\;\frac{(\alpha-\beta)\,\big{(}p_{H}(\pi,s)-p_{L}(\pi,s)\big{)}}{\big{(}1-q_{H}(\pi,s)\big{)}\big{(}1-q_{L}(\pi,s)\big{)}}\;<\;0,$
while $J^{\text{obs},+}$ is independent of $e^{*}$. Hence the $J^{+}/J^{-}$ asymmetry used in Appendix B survives, and so do single crossing and conservatism (with weaker informativeness as $\alpha\downarrow\beta$).
Proposition 67.
If $\alpha>\beta$ (a success flag is more likely when $y=1$ than when $y=0$), then the High type’s best reply remains a cutoff and $s^{*}(\pi)$ is (weakly) increasing in $\pi$. As $\alpha\downarrow\beta$, experiments become less informative and the magnitudes of the jumps in (OA.3) shrink, but the signs in Appendix C (drift) are unchanged.
OA.4.2.3 Partial observability of the recommendation
Suppose the public observes a noisy proxy $\tilde{a}\in\{0,1\}$ for the recommendation $a$, with symmetric accuracy $\mathbb{P}(\tilde{a}=a)=r\in(1/2,1]$ independent of $\theta$ and $\omega$. The public composite signal is $(\tilde{a},z)$. Let $\ell(\tilde{a},z)$ be the likelihood ratio $\mathbb{P}(\tilde{a},z\mid H)/\mathbb{P}(\tilde{a},z\mid L)$. Then:
Lemma 68.
If (i) $r>1/2$ and (ii) outcome classification satisfies $\alpha>\beta$, then ordering $(\tilde{a},z)$ by $\ell(\tilde{a},z)$ preserves the monotone likelihood ratio property for the posterior about $\theta$. In particular, the argument of Lemma 27 carries over with $a$ replaced by $\tilde{a}$ and with jumps as in (OA.3).
Proof.
Under (i)–(ii), both channels—recommendation proxy and outcome—are (weakly) informative in the Blackwell sense and independent conditional on $(\theta,\omega,s)$. Their product likelihood retains the MLRP order (a standard closure property). See also Appendix F for an analogous monitor $m$. ∎
Proposition 69.
Under Lemma 68, the High type’s best reply is again a cutoff in $s$, and $s^{*}(\pi)$ is (weakly) increasing in $\pi$. Comparative statics in policy parameters (e.g., the success bonus $\beta$) extend verbatim after replacing $J^{\pm}$ by $J^{\text{obs},\pm}$ and $a$ by $\tilde{a}$ in the posterior formulas.
OA.4.2.4 Interpretation
Noisy observation of recommendations or outcomes scales informativeness but does not reverse the key asymmetry: higher reputation raises effort, which deepens the reputational loss from a failure relative to the gain from a success. Hence the qualitative predictions—threshold policy, reputational conservatism, and accelerated learning when experimentation is encouraged—are robust to these frictions.
Appendix OA.5 Persistent Environments and Correlated Information
This appendix extends the baseline i.i.d. environment to (i) a persistent technological state and (ii) correlation in private signals across experts/periods. We show that the threshold policy and reputational conservatism survive; the reputation-drift statements adapt with a filtered state belief.
OA.5.1 Two-state Markov environment and belief recursion
Let the period state $\omega_{t}\in\{0,1\}$ follow a two-state Markov chain with transition matrix
$\Gamma\;=\;\begin{pmatrix}\mathbb{P}(\omega_{t+1}=0\mid\omega_{t}=0)&\mathbb{P}(\omega_{t+1}=1\mid\omega_{t}=0)\\ \mathbb{P}(\omega_{t+1}=0\mid\omega_{t}=1)&\mathbb{P}(\omega_{t+1}=1\mid\omega_{t}=1)\end{pmatrix}=\begin{pmatrix}1-\gamma_{01}&\gamma_{01}\\ 1-\gamma_{11}&\gamma_{11}\end{pmatrix},\quad\gamma_{01},\gamma_{11}\in(0,1).$
Let $\lambda_{t}\equiv\mathbb{P}(\omega_{t}=1\mid\mathcal{F}_{t})$ be the public belief about the current state given public history $\mathcal{F}_{t}$ (recommendations and outcomes up to $t$). The public also tracks the expert’s reputation $\pi_{t}\equiv\mathbb{P}(\theta=H\mid\mathcal{F}_{t})$. The state filter $(\lambda_{t})$ updates in two steps each period:
  1. 1.
    Update at $t$ given the period-$t$ public observations. The public observation at $t$ is $(a_{t},y_{t}\mathbf{1}\{a_{t}=1\})$. Let $r_{\theta}(1\mid\omega,\pi_{t},\lambda_{t})$ be the equilibrium risky-recommendation probability of type $\theta$ in state $\omega$ (cutoff strategy), and let $r_{\theta}(0\mid\omega,\cdot)=1-r_{\theta}(1\mid\omega,\cdot)$. Then the state posterior after seeing $a_{t}$ is
    $\mathbb{P}(\omega_{t}=1\mid a_{t},\pi_{t},\lambda_{t})\;=\;\\ \frac{\lambda_{t}\,[\pi_{t}r_{H}(a_{t}\mid 1)+(1-\pi_{t})r_{L}(a_{t}\mid 1)]}{\lambda_{t}\,[\pi_{t}r_{H}(a_{t}\mid 1)+(1-\pi_{t})r_{L}(a_{t}\mid 1)]+(1-\lambda_{t})\,[\pi_{t}r_{H}(a_{t}\mid 0)+(1-\pi_{t})r_{L}(a_{t}\mid 0)]}.$ (OA.4)
    If $a_{t}=1$ and the outcome $y_{t}$ is observed with the baseline technology (success if and only if $\omega_{t}=1$ and the effort succeeds), then
    $\begin{split}\mathbb{P}(\omega_{t}=1\mid a_{t}\!=\!1,y_{t}\!=\!1,\cdot)&=1,\\ \mathbb{P}(\omega_{t}=1\mid a_{t}\!=\!1,y_{t}\!=\!0,\cdot)&=\frac{\mathbb{P}(\omega_{t}=1\mid a_{t},\cdot)\,[1-e^{*}_{t}]}{\mathbb{P}(\omega_{t}=1\mid a_{t},\cdot)\,[1-e^{*}_{t}]+(1-\mathbb{P}(\omega_{t}=1\mid a_{t},\cdot))},\end{split}$ (OA.5)
    where $e^{*}_{t}=e^{*}(1,\pi_{t};\lambda_{t})$ is the implementer’s best reply (Appendix A). With the noisy outcome variant in OA3, replace (OA.5) by Bayes with the observed success probability $q_{\theta}$ in (OA.3).
  2. 2.
    Predict to t+1t{+}1. The one-step prediction is
    $\lambda_{t+1}\;=\;\gamma_{11}\,\lambda_{t|t}+\gamma_{01}\,(1-\lambda_{t|t}),$ (OA.6)
    where $\lambda_{t|t}$ denotes the posterior after incorporating $(a_{t},y_{t}\mathbf{1}\{a_{t}=1\})$ via (OA.4)–(OA.5).
The reputation $\pi_{t}$ updates exactly as in Appendix C, except that the relevant likelihoods are mixtures over the filtered state $\lambda_{t|t}$.
OA.5.1.1 State–augmented dynamic program.
The Markov pair $(\pi_{t},\lambda_{t})$ is a sufficient public state. The Bellman equation becomes
$V(\pi,\lambda)\;=\;u(\pi)\;+\;\delta\,\mathbb{E}\!\left[V(\pi^{\prime},\lambda^{\prime})\,\middle|\,(\pi,\lambda)\right],$ (OA.7)
where $\lambda^{\prime}$ follows (OA.6) and $\pi^{\prime}$ uses the recommendation/outcome LLRs computed at $(\pi,\lambda)$ (Appendix C). The High type’s risky–safe difference at signal $s$ is
$\Delta_{H}(s;\pi,\lambda)=\phi+\delta\Big{(}P_{S}(s;\pi,\lambda)\,V(\pi^{+},\lambda^{\prime})+\big{(}1-P_{S}(s;\pi,\lambda)\big{)}\,V(\pi^{-},\lambda^{\prime})-V(\pi^{\mathrm{rec}}(0),\lambda^{\prime})\Big{)},$
with $P_{S}(s;\pi,\lambda)=\lambda_{t|t}(s)\,e^{*}(1,\pi;\lambda_{t|t}(s))$ and $\lambda_{t|t}(s)$ the state posterior after $a=1$ inferred from (OA.4) using the equilibrium cutoff (the dependence on $s$ is via the probability of $a=1$ under each $(\theta,\omega)$).
Proposition 70.
Fix $(\pi,\lambda)\in(0,1)^{2}$. Under A1–A3 and the Markov environment above, $s\mapsto\Delta_{H}(s;\pi,\lambda)$ is continuous and strictly increasing. Hence there exists a unique cutoff $s^{*}(\pi,\lambda)$ such that the High type recommends risk iff $s\geq s^{*}(\pi,\lambda)$.
Proof.
As in Appendix A. Lemma 27 carries through because (i) for each $(\theta,\omega)$ the signal family has MLRP in $s$ (A1), and (ii) $e^{*}(1,\pi;\cdot)$ is increasing in the state posterior; both the recommendation-only branch and the outcome branch are increasing in $s$. Strictness follows from strict MLRP and $c^{\prime\prime}>0$. ∎
Proposition 71.
The cutoff $s^{*}(\pi,\lambda)$ is weakly increasing in $\pi$ and weakly decreasing in $\lambda$.
Proof.
The decreasing-differences proof in Appendix B applies to $\pi$ verbatim, holding $\lambda$ fixed. For $\lambda$, note that higher $\lambda$ raises $P_{S}$ and makes failures less likely and more informative (through a more negative $J^{-}$ driven by higher effort); both effects raise the marginal value of taking risk at any $s$, shifting the zero of $\Delta_{H}$ left. Formally, $\partial_{\lambda}\Delta_{H}(s;\pi,\lambda)\geq 0$ and $\partial_{s\lambda}\Delta_{H}\leq 0$, so the smallest zero is decreasing in $\lambda$ (Topkis). ∎
Proposition 72.
Let $L_{t+1}$ be the one-step likelihood ratio between period-(t+1)(t{+}1) observations under $H$ and $L$, computed at $(\pi_{t},\lambda_{t})$. Then
$\mathbb{E}\!\left[\pi_{t+1}-\pi_{t}\mid\mathcal{F}_{t},\theta=H\right]\ \geq\ 0,\qquad\mathbb{E}\!\left[\pi_{t+1}-\pi_{t}\mid\mathcal{F}_{t},\theta=L\right]\ \leq\ 0,$
with equality iff $L_{t+1}=1$ a.s. (observations at t+1t{+}1 are uninformative about type). The proof of Lemma 33 extends because mixing over $\omega_{t}$ (via $\lambda_{t}$) preserves the mean-one property of likelihood ratios under $L$, and the map $x\mapsto\frac{x(x-1)}{\pi_{t}x+1-\pi_{t}}$ is increasing.
OA.5.2 Correlated private signals across experts or periods
This subsection states conditions under which our single-crossing and conservatism arguments are robust to correlation in the private-information primitives.
OA.5.2.1 Across experts (within-period)
Suppose in a committee setting there are $n$ experts with private signals $s^{1},\dots,s^{n}$ and joint density $f_{\theta}(s^{1},\dots,s^{n}\mid\omega)$ that is affiliated (log-supermodular) and such that, for each $\theta$ and $\omega$, the marginal family $\{f_{\theta}(\cdot\mid\omega)\}$ on each coordinate has MLRP in the signal. Then:
Lemma 73.
If $f_{\theta}(\cdot\mid\omega)$ is affiliated and has MLRP in each coordinate, then conditioning on other experts’ signals or on any monotone aggregation of them preserves MLRP in one expert’s own signal. In particular, the single-expert marginal likelihood ratio $\frac{f_{H}(s^{i}\mid\omega)}{f_{L}(s^{i}\mid\omega)}$ remains increasing in $s^{i}$ given others’ signals lying in an upper set.
Proof.
A standard property of TP2/affiliated families: log-supermodularity of the joint density implies that the conditional density ratio in one coordinate is (weakly) increasing when the other coordinates increase (Karlin–Rubin monotone likelihood ratio). See, e.g., the MLRP preservation results for affiliated families. ∎
As a consequence, each expert’s best reply in a symmetric equilibrium remains a cutoff in her own signal conditional on the public state; Proposition 71 and the committee pivot arguments in Appendix G carry through.
OA.5.2.2 Across periods (serially correlated signals)
Let the expert’s signal at $t$ be $s_{t}$ with conditional joint density $f_{\theta}(s_{1:t}\mid\omega_{1:t})$. Assume that (i) conditional on $(\theta,\omega_{t})$ and the past, the one-step conditional family $s_{t}\mapsto f_{\theta}(s_{t}\mid\omega_{t},\mathcal{I}_{t-1})$ has MLRP in $s_{t}$ for every realized information set $\mathcal{I}_{t-1}$; (ii) the conditional distribution is TP2 in $(s_{t},\omega_{t})$ (e.g., Gaussian location families with common variance and AR(1) residuals). Then:
Lemma 74.
Under (i)–(ii), for every realized history $\mathcal{I}_{t-1}$ the conditional likelihood ratios $\log\frac{f_{H}(s_{t}\mid\omega_{t},\mathcal{I}_{t-1})}{f_{L}(s_{t}\mid\omega_{t},\mathcal{I}_{t-1})}$ and $\log\frac{f_{\theta}(s_{t}\mid\omega_{t}=1,\mathcal{I}_{t-1})}{f_{\theta}(s_{t}\mid\omega_{t}=0,\mathcal{I}_{t-1})}$ are increasing in $s_{t}$. Consequently, given $(\pi_{t},\lambda_{t})$ the High type’s best reply is a cutoff in $s_{t}$, and reputational conservatism holds as in Proposition 71.
Proof.
Assumption (i) imposes conditional MLRP at each node of the public history; (ii) ensures that the observation kernel remains TP2 in the state, so Bayes posteriors inherit monotone likelihood ratios. The arguments of Appendices A–B apply pointwise in $\mathcal{I}_{t-1}$. ∎
Correlation weakens informativeness quantitatively (effective sample size falls) but not qualitatively: under TP2/affiliation and conditional MLRP, the single-crossing structure and the LLR asymmetry used for reputational conservatism remain intact. With a persistent state, the sufficient public state expands to $(\pi,\lambda)$ and the proofs adapt by conditioning on the filtered state $\lambda$.
Appendix OA.6 Committees: General Aggregation and Heterogeneity
This appendix extends Appendix G beyond $k$-of-$n$ rules and identical members. We allow arbitrary monotone aggregation rules, exchangeable or heterogeneous committee members, and derive pivotality formulas and comparative statics that nest Proposition 57.
OA.6.1 Setup and notation
There are $n$ experts. In period $t$, expert $i$ recommends $a^{i}_{t}\in\{0,1\}$ (risk if $1$), and the committee implements $A_{t}=g(a^{1}_{t},\dots,a^{n}_{t})$, where $g:\{0,1\}^{n}\to\{0,1\}$ is a monotone aggregation rule (if a vector weakly increases coordinatewise, the output weakly increases). The baseline $k$-of-$n$ rule is $g_{k}(\mathbf{a})=\mathbf{1}\{\sum_{i}a^{i}\geq k\}$.
[…]
OA.6.2 Pivotality under arbitrary monotone aggregation
For a fixed rule $g$, define the discrete influence of expert $i$ when the number of other risky votes is $m\in\{0,\dots,n-1\}$:
$\Delta_{g}(m)\;\equiv\;g(1,\underbrace{1,\dots,1}_{m},\underbrace{0,\dots,0}_{n-1-m})\;-\;g(0,\underbrace{1,\dots,1}_{m},\underbrace{0,\dots,0}_{n-1-m})\ \in\ \{0,1\}.$ (OA.8)
Monotonicity of $g$ implies $\Delta_{g}(m)$ is weakly single-peaked in $m$ (it can switch from 0 to $1$ at most once). The period-pivotality of expert $i$ (the chance her recommendation flips $A$) is
$\zeta_{g}(\pi)\;=\;\sum_{\omega\in\{0,1\}}\mathbb{P}(\omega\mid\pi)\ \mathbb{E}\Big{[}\ \Delta_{g}(M_{\omega})\ \Big{]},$ (OA.9)
where $M_{\omega}\equiv\sum_{j\neq i}a^{j}\mid\omega$ is the number of other risky recommendations conditional on $\omega$.
Lemma 75.
[…]
$\zeta_{g}(\pi)\;=\;\sum_{\omega}\mathbb{P}(\omega\mid\pi)\ \sum_{m=0}^{n-1}\Delta_{g}(m)\binom{n-1}{m}\,\big{(}\rho_{\omega}^{-i}(\pi)\big{)}^{m}\big{(}1-\rho_{\omega}^{-i}(\pi)\big{)}^{n-1-m}.$ (OA.10)
In particular, for the $k$-of-$n$ rule, $\Delta_{g_{k}}(m)=\mathbf{1}\{m=k-1\}$ and
$\zeta_{k}(\pi)\;=\;\sum_{\omega}\mathbb{P}(\omega\mid\pi)\ \binom{n-1}{k-1}\,\big{(}\rho_{\omega}^{-i}(\pi)\big{)}^{k-1}\big{(}1-\rho_{\omega}^{-i}(\pi)\big{)}^{n-k}.$ (OA.11)
Lemma 76.
If $a^{j}$ are independent with $\mathbb{P}(a^{j}=1\mid\omega)=\rho_{\omega}^{j}(\pi)$, then $M_{\omega}$ has Poisson–binomial pmf $p_{\omega}(m)=\mathbb{P}(M_{\omega}=m)=\sum_{S:|S|=m}\prod_{j\in S}\rho_{\omega}^{j}\prod_{j\notin S}(1-\rho_{\omega}^{j})$ and
$\zeta_{g}(\pi)\;=\;\sum_{\omega}\mathbb{P}(\omega\mid\pi)\ \sum_{m=0}^{n-1}\Delta_{g}(m)\,p_{\omega}(m).$ (OA.12)
For $g_{k}$, $\zeta_{k}(\pi)=\sum_{\omega}\mathbb{P}(\omega\mid\pi)\,p_{\omega}(k-1)$.
Proofs. Both lemmas are immediate from (OA.9) and the distributions of $M_{\omega}$.
OA.6.3 From pivotality to the individual cutoff
For expert $i$, the outcome-based component of the risky–safe gain is scaled by $\zeta_{g}(\pi)$, while the signaling (recommendation) component is unchanged by $g$. Hence, holding others’ strategies fixed, the High type’s risky–safe difference at $s$ takes the form
$\begin{split}\Delta_{H,i}(s;\pi)\;=\;\phi\;&+\;\delta\Big{(}\,\zeta_{g}(\pi)\cdot\underbrace{\big{[}P_{S}(s;\pi)\,V(\pi^{+})+(1-P_{S}(s;\pi))V(\pi^{-})-V(\pi)\big{]}}_{\text{\emph{outcome option value}}}\;\\ &-\;\underbrace{\big{[}V(\pi)-V(\pi^{\mathrm{rec}}(0))\big{]}}_{\text{\emph{recommendation-only term}}}\Big{)}.\end{split}$ (OA.13)
Since $\zeta_{g}(\pi)$ does not depend on $s$, the single-crossing of $\Delta_{H,i}(\cdot;\pi)$ in $s$ (Appendix A) is preserved and the optimal policy is a cutoff; comparative statics in $\zeta_{g}(\pi)$ follow by Topkis.
OA.6.4 Comparative statics in thresholds $k$ and in aggregation rules
Proposition 77.
Under exchangeability with state-contingent risky frequencies $\rho_{\omega}^{-i}(\pi)\in[0,1]$, the pivotality (OA.11) for the $k$-of-$n$ rule is strictly decreasing in $k$ whenever k1k-1 lies below the mean of $M_{\omega}$ for at most one state and the binomial pmf is unimodal; in particular, if $\rho_{\omega}^{-i}(\pi)\leq k/n$ for both $\omega$, then $\zeta_{k+1}(\pi)<\zeta_{k}(\pi)$. Consequently, the High type’s cutoff $s^{*}(\pi;k)$ is (weakly) increasing in $k$.
Proof.
For $g_{k}$, $\zeta_{k}(\pi)$ equals a mixture of the binomial pmf evaluated at m=k1m=k-1. Unimodality implies the pmf decreases as we move away from its mode; the displayed sufficient condition puts k1k-1 weakly to the right of the mode, so increasing $k$ lowers the mass. The cutoff monotonicity follows from (OA.13) since $\partial\Delta_{H,i}/\partial\zeta_{g}>0$ pointwise in $s$. ∎
Proposition 78.
Fix $\pi$. Let $g$ and $\tilde{g}$ be two monotone aggregation rules with discrete influences ordered as $\Delta_{g}(m)\leq\Delta_{\tilde{g}}(m)$ for all $m$. Then $\zeta_{g}(\pi)\leq\zeta_{\tilde{g}}(\pi)$ under both exchangeable and heterogeneous others. Hence $s^{*}_{g}(\pi)\geq s^{*}_{\tilde{g}}(\pi)$.
Proof.
By (OA.10)–(OA.12), $\zeta$ is a linear functional of $\Delta_{g}(\cdot)$ with nonnegative weights; the result is immediate. The cutoff comparison again follows from (OA.13). ∎
OA.6.5 Heterogeneous members: priors and precisions
Allow heterogeneity in priors $\pi^{j}$ (about type) and signal precisions (so each member $j$ uses her own cutoff, generating state-contingent risky probabilities $\rho_{\omega}^{j}(\pi)$). Let $p_{\omega}(m)$ be the corresponding Poisson–binomial pmf.
Proposition 79.
Fix $\pi$ and a monotone $g$. If the distribution of $M_{\omega}$ under profile $\{\rho_{\omega}^{j}\}$ first-order stochastically dominates its distribution under $\{\tilde{\rho}_{\omega}^{j}\}$ for each state $\omega$, then $\zeta_{g}(\pi)$ is (weakly) larger under $\{\rho_{\omega}^{j}\}$ than under $\{\tilde{\rho}_{\omega}^{j}\}$. Therefore the individual cutoff is (weakly) lower when others are more prone to recommend risk.
Proof.
For any monotone $g$, the influence kernel $\Delta_{g}(m)$ is increasing in $m$ (in the sense of upper sets). FOSD of $M_{\omega}$ raises the expectation of any increasing function, so it raises $\mathbb{E}[\Delta_{g}(M_{\omega})]$ for each $\omega$, and thus $\zeta_{g}(\pi)$ by (OA.9). The cutoff conclusion follows from (OA.13). ∎
Corollary 80.
For $g_{k}$ with heterogeneous others, $\zeta_{k}(\pi)=\sum_{\omega}\mathbb{P}(\omega\mid\pi)\,p_{\omega}(k-1)$. If k1k-1 lies weakly to the right of the mean of $M_{\omega}$ for both $\omega$, then $\zeta_{k}(\pi)$ is decreasing in $k$; consequently, $s^{*}(\pi;k)$ is (weakly) increasing in $k$.
OA.6.5.1 Interpretation
Aggregation affects only the outcome option value through pivotality; the signaling value of an own recommendation is unchanged. Rules that make the individual less pivotal (higher thresholds $k$, stricter monotone aggregators, or more conservative peers) raise the cutoff and depress experimentation, strengthening Proposition 57 beyond the symmetric binomial benchmark.
OA.6.6 Remarks on dynamics and welfare
(i) The dynamic impact of $k$ or $g$ operates through $\zeta_{g}(\pi)$ in each period; when the state is persistent, replace $\pi$ by $(\pi,\lambda)$ everywhere. (ii) In welfare or policy design, $\zeta_{g}(\pi)$ is the sufficient statistic for how aggregation amplifies or dampens the reputational incentives to experiment; combining it with the success-bonus formulas in Appendix I yields transparent levers for restoring experimentation.
Appendix OA.7 Monitoring and Disclosure—Variants
This appendix extends Appendix F along four dimensions: (i) timing (pre- vs. post-outcome monitors), (ii) multi-level monitors, (iii) Blackwell comparisons allowing action-dependent noise, and (iv) endogenous disclosure of a verifiable monitor. We maintain Assumptions A1–A4 and notation from Sections 35.
OA.7.1 Timing: pre- vs. post-outcome monitors
Let $m$ be an auxiliary public signal with finite ordered support $\mathcal{M}=\{m_{1}<\dots<m_{K}\}$. We allow its law to depend on the action and type, and (optionally) on the technological state:
$\mathbb{P}(m\mid a,\theta,\omega)\,.$
We consider two timings.
OA.7.1.1 Post-outcome monitor (baseline in Appendix F).
After $(a,y)$, the public observes $m$ (e.g., a post-hoc audit or registry flag). The one-step reputation update composes the LLRs of $(a,y)$ with the LLR of $m$:
$L_{t+1}\;=\;L^{\mathrm{rec}}_{t+1}\cdot L^{\mathrm{out}}_{t+1}\cdot L^{m}_{t+1},\qquad L^{m}_{t+1}\equiv\frac{\mathbb{P}(m\mid a,\theta=H,\omega)}{\mathbb{P}(m\mid a,\theta=L,\omega)}.$
When $\mathbb{P}(m\mid a=0,\cdot)$ is informative about $\theta$, successes under $a=1$ are relatively less informative (the “good news” $m$ can arrive even under the safe action), weakening the option value of experimentation; see Proposition 53.
OA.7.1.2 Pre-outcome monitor (between $a$ and effort $e$).
Suppose $m$ arrives after $a$ is publicly observed but before the implementer chooses effort $e$. Then the implementer conditions on $m$ when solving $c^{\prime}(e)=\lambda(1,\pi;s,m)$, yielding
$e^{*}(1,\pi;s,m)\;=\;(c^{\prime})^{-1}\!\big{(}\lambda(1,\pi;s,m)\big{)}.$
The success probability under $a=1$ becomes $P_{S}(s;\pi)=\mathbb{E}_{m}\!\left[\lambda(1,\pi;s,m)\,e^{*}(1,\pi;s,m)\right]$ (expectation under the public predictive distribution of $m$ given $(a=1,\pi)$).
Lemma 81.
If $m$ satisfies MLRP in the sense that $m\uparrow$ raises $\lambda(1,\pi;s,m)$ (for each $\pi,s$), then $e^{*}(1,\pi;s,m)$ is increasing in $m$, and $P_{S}(s;\pi)$ is (weakly) larger than in the no-monitoring benchmark, with strict inequality whenever $m$ is non-degenerate and $c^{\prime\prime}>0$.
Proof.
Monotonicity of $e^{*}$ follows from strict convexity of $c$. Jensen’s inequality with strict curvature delivers the strict increase in the expectation unless $m$ is degenerate. ∎
Proposition 82.
Fix $\pi$. Suppose (i) $m$ is (weakly) more informative about $\theta$ under $a=1$ than under $a=0$ in the Blackwell sense, and (ii) $m$ has MLRP for $\lambda(1,\pi;s,m)$ as in Lemma 81. Then relative to no monitoring, a pre-outcome $m$ raises $\Delta_{H}(s;\pi)$ pointwise in $s$, strictly so when $m$ is strictly informative; the cutoff $s^{*}(\pi)$ falls. In contrast, a post-outcome $m$ that is (weakly) more informative under $a=0$ than under $a=1$ lowers $\Delta_{H}(s;\pi)$ and raises $s^{*}(\pi)$ (experimentation-suppressing, cf. Proposition 53).
Proof.
Pre-outcome $m$ has two effects: it raises $P_{S}$ by Lemma 81 and, under (i), weakly increases the informativeness of $(a,m,y)$ under $a=1$ relative to $a=0$, enlarging the outcome option value in $\Delta_{H}$. For post-outcome $m$, the latter force flips sign when $m$ is more informative under $a=0$, reducing the relative informativeness of success; $P_{S}$ is unaffected because effort is chosen before $m$ is observed. ∎
OA.7.2 Multi-level monitors
Let $m\in\{m_{1}<\cdots<m_{K}\}$ with $\mathbb{P}(m_{k}\mid a,\theta,\omega)$ such that, for each $(a,\omega)$, the family $\{m\mapsto\mathbb{P}(m\mid a,\theta,\omega)\}_{\theta\in\{H,L\}}$ has MLRP. Then:
  • The recommendation-only posterior $\pi^{\mathrm{rec}}(a)$ and the outcome posteriors $\pi^{\pm}$ extend by integrating the LLRs over $m$ (post-outcome timing) or by conditioning on $m$ in the implementer’s effort (pre-outcome timing).
  • The success jump $J^{+}$ remains independent of effort, while the failure jump $J^{-}$ becomes more negative as $m$ increases pre-outcome (via higher $e^{*}$); post-outcome $J^{\pm}$ incorporate the additional LLR from $m$ analogously to (11)–(12).
All monotonicity and single-crossing arguments in Appendices AB carry over with $m$-indexed mixtures.
OA.7.3 Blackwell comparisons with action-dependent noise
We compare two monitoring structures $\mathcal{M}$ and $\widetilde{\mathcal{M}}$ through (action-conditional) Blackwell dominance.
Definition 83.
We say $\widetilde{\mathcal{M}}\succeq_{B,a}\mathcal{M}$ if, for a given $a\in\{0,1\}$, there exists a stochastic kernel $K_{a}$ such that $\mathbb{P}_{\mathcal{M}}(m\mid a,\theta,\omega)=\sum_{\tilde{m}}K_{a}(m\mid\tilde{m})\mathbb{P}_{\widetilde{\mathcal{M}}}(\tilde{m}\mid a,\theta,\omega)$ for both $\theta$. Strict dominance $\succ_{B,a}$ holds if no such garbling exists in the reverse direction.
Proposition 84.
If $\widetilde{\mathcal{M}}\succeq_{B,1}\mathcal{M}$ and $\widetilde{\mathcal{M}}\preceq_{B,0}\mathcal{M}$ (i.e., the refinement is for $a=1$ and, weakly, a garbling for $a=0$), then for post-outcome timing the risky–safe difference satisfies $\widetilde{\Delta}_{H}(s;\pi)\geq\Delta_{H}(s;\pi)$ for all $s,\pi$, with strict inequality on a set of positive measure if either dominance is strict and the baseline has an interior cutoff. Therefore $s^{*}_{\widetilde{\mathcal{M}}}(\pi)\leq s^{*}_{\mathcal{M}}(\pi)$.
Proof.
Under post-outcome timing, $m$ only affects informativeness. The outcome option value in $\Delta_{H}$ is an increasing functional of the action-$1$ experiment’s informativeness (the Jensen gap with convex $V$), and a decreasing functional of the action-0 informativeness (since the safe branch’s $V(\pi^{\mathrm{rec}}(0,m))$ rises with more informative $m$ under $a=0$). Blackwell dominance composes with Bayes posteriors; the inequality follows pointwise. Strictness uses strict dominance and interiority of the root together with continuity. ∎
Corollary 85.
If $\widetilde{\mathcal{M}}\succ_{B,1}\mathcal{M}$ and $\widetilde{\mathcal{M}}\prec_{B,0}\mathcal{M}$, and if the baseline cutoff is interior ($P_{S}\in(0,1)$ at $s^{*}$), then $s^{*}_{\widetilde{\mathcal{M}}}(\pi)<s^{*}_{\mathcal{M}}(\pi)$ for all $\pi$ in a neighborhood of any point where dominance is strict.
For pre-outcome timing, combine Proposition 82 (effort amplification) with the same Blackwell comparisons to obtain analogous conclusions.
OA.7.4 Endogenous release (verifiable disclosure)
Suppose $m$ is verifiable and observed by a sender (expert or implementer) who can choose whether to disclose it publicly. Let $d\in\{0,1\}$ denote the disclosure decision, with $d=1$ revealing $m$ and $d=0$ revealing nothing. Assume $d$ is chosen to maximize the sender’s continuation value after $a$ is realized; disclosure is costless.
Proposition 86.
If $m$ is one-dimensional and the sender’s continuation value is (strictly) increasing in $m$, the unique equilibrium is full unraveling: all realizations of $m$ are disclosed. With a small disclosure cost $\kappa>0$, the equilibrium is partial unraveling: there exists a cutoff $\hat{m}$ such that $m\geq\hat{m}$ is disclosed and $m<\hat{m}$ is pooled as “no disclosure”.
Proof.
Standard verifiable-disclosure logic: higher $m$ yields a strictly higher continuation value, so the sender with $m^{\prime}$ can profitably mimic any sender with lower $m<m^{\prime}$ if that $m$ were disclosed; hence in equilibrium no lower type can be pooling with a higher disclosed $m$. Iterating from the top yields full unraveling. With a small cost $\kappa$, a threshold policy arises by standard single-crossing arguments. ∎
Proposition 87.
Under full unraveling in pre-outcome timing, the effects coincide with those of publicly observed $m$ (Lemma 81 and Proposition 82). Under partial unraveling with cutoff $\hat{m}$, the public likelihood under $a=0$ is a mixture of disclosed $m\geq\hat{m}$ and an unrevealing “no-disclosure” message. Relative to full disclosure, this weakly reduces informativeness under $a=0$ and therefore (weakly) raises experimentation in post-outcome timing. In pre-outcome timing, experimentation additionally reacts to the induced change in expected effort $\mathbb{E}[e^{*}\mid\text{disclosure policy}]$; the net effect preserves the signs in Proposition 82.
Monitors that arrive before effort amplify $P_{S}$ through effort conditioning and encourage experimentation when they are (relatively) more informative under $a=1$ than under $a=0$. Post-outcome monitors that are informative under $a=0$ dilute the reputational content of success and suppress experimentation. Multi-level and action-dependent monitors fit cleanly into Blackwell comparisons; strict refinements generate strict shifts in cutoffs at interior regions. Verifiable disclosure pushes toward full revelation; when disclosure is partial, the induced pooling weakly reduces safe-branch informativeness, which typically moves the policy in the pro-experimentation direction.
Appendix OA.8 Policy Design Details
Link to main text. This section contains the full proofs of Theorem 23 and Proposition 24 from the main paper, as well as calibration formulas (OA.21)–(OA.22) and the planner’s FOC (OA.16). Throughout, primitives and notation follow Sections 35 and Appendix I. In particular, a success bonus $\beta\geq 0$ pays the expert upon $(a=1,y=1)$, affects only the expert’s objective (the implementer’s effort $e^{*}(\cdot)$ is unchanged), and rotates the risky–safe margin by $\beta\,P_{S}(s;\pi)$ as in (29). The affine extension pays $(\beta_{1},-\beta_{0})$ on $(y=1,y=0)$ when $a=1$, yielding (31).
OA.8.1 Planner’s objective and a simple FOC for the optimal bonus
Let the planner value each successful risky implementation at $B>0$ (net of the safe action’s payoff, normalized to 0) and bear the real resource cost of the transfer at shadow price $\eta\geq 0$. In our baseline, the bonus is financed externally (no offsetting benefit elsewhere). The implementer’s effort cost $c(e)$ is borne socially. For a given public reputation $\pi$, define the per-experiment social surplus
$\begin{split}S_{\mathrm{exp}}(\pi)\;&\equiv\;\mathbb{E}\!\left[\,B\cdot y-c\!\big{(}e^{*}(1,\pi;s)\big{)}\ \Big{|}\ a=1,\ \pi\right]\\ \;&x=\;B\cdot\mathbb{E}[y\mid a=1,\pi]\ -\ \mathbb{E}\!\left[c\!\big{(}e^{*}(1,\pi;s)\big{)}\mid a=1,\pi\right].\end{split}$ (OA.14)
Let $\rho(\pi;\beta)$ be the probability that the High type recommends $a=1$ at $\pi$ under bonus $\beta$ (in the Gaussian benchmark, $\rho(\pi;\beta)=\lambda A_{H}(s^{*}_{\beta}(\pi))+(1-\lambda)B_{H}(s^{*}_{\beta}(\pi))$). Let $P_{S}(\pi)$ denote the success probability conditional on $a=1$ (integrating out $s$ at the equilibrium cutoff). Define the per-period success rate
$R_{\mathrm{succ}}(\beta)\;\equiv\;\mathbb{E}\big{[}\rho(\pi_{t};\beta)\,P_{S}(\pi_{t})\big{]},$
where the expectation is with respect to the stationary distribution of $\pi_{t}$ induced by the equilibrium policy under $\beta$ (for a local analysis, it is enough to treat this distribution as fixed at $\beta$).
The planner’s steady-state objective is
$W(\beta)\;=\;\mathbb{E}\big{[}\rho(\pi_{t};\beta)\,S_{\mathrm{exp}}(\pi_{t})\big{]}\ -\ \eta\,\beta\,R_{\mathrm{succ}}(\beta).$ (OA.15)
Because the bonus does not affect effort directly in our setup, $S_{\mathrm{exp}}(\pi)$ is independent of $\beta$. [^18] If the transfer did directly affect the implementer (e.g., a pay-for-performance contract with the agent), an extra term $\mathbb{E}[\rho(\pi;\beta)\,\partial S_{\mathrm{exp}}/\partial\beta]$ would appear in the FOC. Differentiating (OA.15) gives the first-order condition:
$W^{\prime}(\beta)\;=\;\underbrace{\mathbb{E}\!\left[\frac{\partial\rho(\pi_{t};\beta)}{\partial\beta}\,S_{\mathrm{exp}}(\pi_{t})\right]}_{\text{behavioral benefit from more experiments}}\ -\ \eta\,\underbrace{\Big{(}R_{\mathrm{succ}}(\beta)\ +\ \beta\,R^{\prime}_{\mathrm{succ}}(\beta)\Big{)}}_{\text{marginal budget cost}}\;=\;0.$ (OA.16)
Using that effort is unaffected by $\beta$, $R^{\prime}_{\mathrm{succ}}(\beta)=\mathbb{E}\!\left[\tfrac{\partial\rho(\pi_{t};\beta)}{\partial\beta}\,P_{S}(\pi_{t})\right]$. Hence (OA.16) can be written as
$\mathbb{E}\!\left[\frac{\partial\rho(\pi_{t};\beta)}{\partial\beta}\,\Big{(}S_{\mathrm{exp}}(\pi_{t})\ -\ \eta\,\beta\,P_{S}(\pi_{t})\Big{)}\right]\;=\;\eta\,R_{\mathrm{succ}}(\beta).$ (OA.17)
Equation (OA.17) links the policy elasticity $\partial\rho/\partial\beta$ to (i) the per-experiment surplus $S_{\mathrm{exp}}$ and (ii) the shadow price of public funds $\eta$ scaled by success.
OA.8.1.1 Local optimal subsidy
At $\beta=0$, $R^{\prime}_{\mathrm{succ}}(0)=\mathbb{E}\!\left[\tfrac{\partial\rho(\pi_{t};0)}{\partial\beta}\,P_{S}(\pi_{t})\right]$ and (OA.16) reduces to
$\mathbb{E}\!\left[\frac{\partial\rho(\pi_{t};0)}{\partial\beta}\,S_{\mathrm{exp}}(\pi_{t})\right]\;=\;\eta\,R_{\mathrm{succ}}(0).$ (OA.18)
If the left-hand side exceeds $\eta R_{\mathrm{succ}}(0)$, then a small positive bonus raises welfare; otherwise, the optimal $\beta$ at the boundary is 0. For interior optima with small $\beta$, a Newton step gives
$\beta^{\star}\ \approx\ \frac{\mathbb{E}\big{[}\tfrac{\partial\rho(\pi_{t};0)}{\partial\beta}\,S_{\mathrm{exp}}(\pi_{t})\big{]}\ -\ \eta\,R_{\mathrm{succ}}(0)}{\eta\,\mathbb{E}\big{[}\tfrac{\partial\rho(\pi_{t};0)}{\partial\beta}\,P_{S}(\pi_{t})\big{]}}\,.$
OA.8.1.2 Relating $\partial\rho/\partial\beta$ to cutoffs
By Proposition 61,
$\frac{\partial s^{*}_{\beta}(\pi)}{\partial\beta}\;=\;-\frac{P_{S}\!\left(s^{*}_{\beta}(\pi);\pi\right)}{\partial_{s}\Delta_{H}^{\beta}\!\left(s^{*}_{\beta}(\pi);\pi\right)}\;\leq 0,$
and $\rho(\pi;\beta)=R_{H}\!\big{(}s^{*}_{\beta}(\pi)\big{)}$, so
$\frac{\partial\rho(\pi;\beta)}{\partial\beta}\;=\;R_{H}^{\prime}\!\big{(}s^{*}_{\beta}(\pi)\big{)}\cdot\frac{\partial s^{*}_{\beta}(\pi)}{\partial\beta}\;=\;-\;R_{H}^{\prime}\!\big{(}s^{*}_{\beta}(\pi)\big{)}\cdot\frac{P_{S}\!\left(s^{*}_{\beta}(\pi);\pi\right)}{\partial_{s}\Delta_{H}^{\beta}\!\left(s^{*}_{\beta}(\pi);\pi\right)}\ \geq 0,$ (OA.19)
since $R_{H}^{\prime}$ is negative in a (strict) tail family. Plugging (OA.19) into (OA.16) or (OA.18) gives a fully microfounded design rule.
OA.8.2 Affine success/failure contracts and implementation regions
Consider the affine transfer in Appendix I:
$\begin{split}\Delta_{H}^{\beta_{1},\beta_{0}}(s;\pi)\;=\;\phi\;&+\;\big{[}\beta_{1}P_{S}(s;\pi)-\beta_{0}(1-P_{S}(s;\pi))\big{]}\;\\ &+\;\delta\Big{\{}P_{S}V(\pi^{+})+(1-P_{S})V(\pi^{-})-V(\pi^{\mathrm{rec}}(0))\Big{\}}.\end{split}$
Fix a target cutoff $\tilde{s}(\pi)$ for each $\pi$. The set of $(\beta_{1},\beta_{0})$ pairs that implement $\tilde{s}(\pi)$ (i.e., make $\tilde{s}(\pi)$ the High-type threshold) is the closed half-plane
$\mathcal{B}(\pi;\tilde{s})\;=\;\left\{(\beta_{1},\beta_{0})\in\mathbb{R}_{+}^{2}:\ \beta_{1}P_{S}\!\big{(}\tilde{s}(\pi);\pi\big{)}\ -\ \beta_{0}\big{(}1-P_{S}\!\big{(}\tilde{s}(\pi);\pi\big{)}\big{)}\ \geq\ -\,\Delta_{H}\!\big{(}\tilde{s}(\pi);\pi\big{)}\right\}.$ (OA.20)
[…]
Success-only bonus: $\beta_{1}^{\min}(\pi,\tilde{s})\;=\;\frac{-\,\Delta_{H}\!\big{(}\tilde{s}(\pi);\pi\big{)}}{P_{S}\!\big{(}\tilde{s}(\pi);\pi\big{)}}\,,\qquad\beta_{0}=0,$ (OA.21)
Failure-only penalty: $\beta_{0}^{\min}(\pi,\tilde{s})\;=\;\frac{\,\Delta_{H}\!\big{(}\tilde{s}(\pi);\pi\big{)}}{1-P_{S}\!\big{(}\tilde{s}(\pi);\pi\big{)}}\,,\qquad\beta_{1}=0.$ (OA.22)
These coincide with the one-dimensional formula in (30) when $\beta_{0}=0$.
OA.8.2.1 Corners and (in)effectiveness
If $P_{S}\!\big{(}\tilde{s}(\pi);\pi\big{)}=0$, a success bonus is locally ineffective (no successes at the target); only a failure penalty can move the cutoff. If $P_{S}\!\big{(}\tilde{s}(\pi);\pi\big{)}=1$, the failure penalty is locally ineffective and the success bonus suffices. For interior $P_{S}\in(0,1)$, any convex combination of (OA.21)–(OA.22) works.
OA.8.2.2 Budgeted design
If the planner faces a per-period budget $\bar{B}$ for transfers, the feasibility constraint is $\eta\,\beta_{1}\,R_{\mathrm{succ}}(\beta_{1},\beta_{0})\ -\ \eta\,\beta_{0}\,R_{\mathrm{fail}}(\beta_{1},\beta_{0})\ \leq\bar{B}$, where
$R_{\mathrm{fail}}(\beta_{1},\beta_{0})\;\equiv\;\mathbb{E}\big{[}\rho(\pi;\beta_{1},\beta_{0})\,(1-P_{S}(\pi))\big{]}.$
Optimal $(\beta_{1},\beta_{0})$ solve (OA.16) with $R_{\mathrm{succ}}$ replaced by $R_{\mathrm{succ}}-\beta_{0}R^{\prime}_{\mathrm{fail}}$ and a multiplier on the budget. In many applications, $\beta_{0}=0$ is natural (penalizing the expert for failure may be infeasible), and (OA.16) applies verbatim.
OA.8.2.3 Welfare interpretation
The success bonus is the right instrument when the planner wants to stimulate experimentation where $S_{\mathrm{exp}}(\pi)$ is high and $P_{S}(\pi)$ is not too close to zero (so the policy is powerful); the failure penalty is the mirror image for discouraging experimentation where $S_{\mathrm{exp}}(\pi)$ is low or negative. The elasticity term $\partial\rho/\partial\beta$ in (OA.16)–(OA.19) provides the microfoundation for targeting—highest reputations often yield the largest local effect because $P_{S}$ is highest there.
A success bonus $\beta$ trades off the marginal surplus generated by extra experiments, $\mathbb{E}[(\partial\rho/\partial\beta)\,S_{\mathrm{exp}}]$, against the resource cost of paying for successes, $\eta(R_{\mathrm{succ}}+\beta R^{\prime}_{\mathrm{succ}})$. Affine contracts enlarge the implementation set linearly in $P_{S}$ and $1-P_{S}$, with clear corner behavior when successes are (near) impossible or certain.